### v1

# Boosting Robustness by Clipping Gradients in Distributed Learning

Abstract
Robust distributed learning consists in achieving good learning performance despite the presence of misbehaving workers. State-of-the-art (SOTA) robust distributed gradient descent (Robust-DGD) methods, relying on robust aggregation, have been proven to be optimal: Their learning error matches the lower bound established under the standard heterogeneity model of $(G,B)$-gradient dissimilarity. The learning guarantee of SOTA Robust-DGD cannot be further improved when model initialization is done arbitrarily. However, we show that it is possible to circumvent the lower bound, and improve the learning performance, when the workers’ gradients at model initialization are assumed to be bounded. We prove this by proposing pre-aggregation clipping of workers’ gradients, using a novel scheme called adaptive robust clipping (ARC). Incorporating ARC in Robust-DGD provably improves the learning, under the aforementioned assumption on model initialization. The factor of improvement is prominent when the tolerable fraction of misbehaving workers approaches the breakdown point. ARC induces this improvement by constricting the search space, while preserving the robustness property of the original aggregation scheme at the same time. We validate this theoretical finding through exhaustive experiments on benchmark image classification tasks.
1 Introduction
Distributed machine learning, a.k.a. federated learning, has emerged as a dominant paradigm to cope with the increasing computational cost of learning tasks, mainly due to growing model sizes and datasets [25]. Worker machines, holding each a fraction of the training dataset, collaborate over a network to learn an optimal common model over the collection of their datasets. Workers typically collaborate with the help of a central coordinator, that we call server [36]. Besides scalability, distributed learning is also helpful in preserving data privacy, since the workers do not have to share their local datasets during the learning.
Conventional distributed learning algorithms are known to be vulnerable to misbehaving workers that could behave unpredictably [7, 25, 18]. Misbehavior may result from software and hardware bugs, data poisoning, or malicious players controlling part of the network. In the parlance of distributed computing, misbehaving workers are referred to as Byzantine [32]. Due to the growing influence of distributed learning in critical public-domain applications, such as healthcare [38] and finance [34], the problem of robustness to misbehaving workers, a.k.a. robust distributed learning, has received significant attention in the past [7, 44, 12, 14, 27, 17, 3, 13].
Robust distributed learning algorithms primarily rely on robust aggregation, such as coordinate-wise trimmed mean (CWTM) [44], geometric median (GM) [10] and multi-Krum (MK) [7]. Specifically, in a robust distributed gradient descent (Robust-DGD) method, the server aggregates the workers’ local gradients using a robust aggregation rule, instead of simply averaging them. This protects the learning procedure from erroneous gradients sent by misbehaving workers. Recent work has made significant improvements over these aggregation rules: by incorporating a pre-aggregation step such as bucketing [27, 17] and nearest-neighbor mixing (NNM) [3] to tackle gradient dissimilarity that results from data heterogeneity. The learning guarantee of a resulting Robust-DGD method has been proven to be optimal [4], i.e., it cannot be improved unless additional assumptions are made, under the standard heterogeneity model of $(G,B)$-gradient dissimilarity [28].
Despite the theoretical tightness of state-of-the-art (SOTA) Robust-DGD, we demonstrate that its performance can be improved in pragmatic cases wherein model initialization incurs bounded errors on workers’ local data. We do so by proposing clipping of workers’ gradients prior to their aggregation (at the server) using a novel clipping scheme, named adaptive robust clipping (ARC). We show that incorporating ARC provably improves the robustness of Robust-DGD when the local gradients at the initial model estimate are bounded. The factor of improvement is prominent when the tolerable fraction of misbehaving workers approaches the breakdown point, [^1] The breakdown point is the minimum fraction of misbehaving workers that can break the system, i.e., makes it impossible to guarantee any bound on the learning error [18]. established in [4] under $(G,B)$-gradient dissimilarity. In general, when the local gradients can be arbitrarily large at initialization, ARC recovers the original learning guarantee of Robust-DGD.
Main results & contributions.
We consider a system comprising $n$ workers and a server. The goal is to tolerate up to $f$ misbehaving (a.k.a. adversarial) workers. The impact of data heterogeneity is characterized by the standard definition of $(G,B)$-gradient dissimilarity [28], recalled in Section 2.
Figure 1: Worst-case maximal accuracies across several types of misbehaviour (see Section 6) of CWTM $\circ$ NNM and GM $\circ$ NNM, with ARC and without (i.e., no clipping), for a heterogeneously distributed MNIST dataset with $10$ honest workers. Heterogeneity is simulated using Dirichlet [23] distribution of parameter $\alpha=0.5$.
(1) Adaptive robust clipping (ARC). We propose ARC, wherein prior to aggregating the gradients, the server clips the largest $k\coloneqq\left\lfloor{2(f/n)(n-f)}\right\rfloor$ gradients using a clipping parameter given by the (Euclidean) norm of the $(k+1)$-th largest gradient. We prove 2 crucial properties: ARC (i) preserves the robustness guarantee of original aggregation and (ii) constrains an adversarial gradient’s norm by that of an honest (non-adversarial) worker’s gradient.
(2) Improved learning error. Under $(G,B)$-gradient dissimilarity, [4] proves a lower bound on the learning error in the presence of adversarial workers. We show, however, that this lower bound can be circumvented under the assumption that the honest workers’ gradients are bounded at model initialization. Specifically, under this assumption, we prove that incorporating ARC in Robust-DGD improves the learning error beyond the lower bound when $f/n$ approaches the breakdown point. Thanks to its two aforementioned properties, ARC induces this improvement by effectively reducing the heterogeneity from $(G,B)$ to $(G+B\rho,0)$-gradient dissimilarity, where $\rho$ depends linearly on the norm of workers’ gradients at model initialization. We remark that the above assumption on model initialization is often satisfiable in practice [15].
(3) Empirical validation. We conduct experiments on MNIST [11], Fashion-MNIST [42], and CIFAR-10 [31], under several data heterogeneity and adversarial regimes. Our results show that ARC boosts the performance of SOTA robust aggregation methods, especially when considering a large number of adversarial workers. A preview of our results on MNIST is shown in Figure 1. We make similar observations on the effectiveness of ARC in scenarios of high heterogeneity (see Figure 3).
Related work.
Gradient clipping is a well-known technique for tackling exploding gradients in deep learning [16]. It has been extensively analyzed in the centralized setting [46, 45, 30], with applications to differential privacy [39]. There is also some work analyzing the impact of clipping in non-robust distributed learning [47, 29]. It has been shown that the clipping of local gradients causes distributed gradient descent to diverge. This motivates the need for an analysis of the clipping technique in our adversarial setting, and also hints at the inadequacy of clipping with a fixed threshold.
In the context of robust distributed learning, prior work has proposed iterative clipping for robust aggregation [26]. The analysis, however, cannot be applied to study the impact of pre-aggregation clipping that we consider. An adaptive gradient clipping scheme was also proposed for robust aggregation in [20, 33]. However, the robustness guarantees are only provided for strongly convex loss functions and under a specific redundancy condition that is inadequate for our non-convex machine learning setting. Other work [37, 14, 3] has only used static pre-aggregation clipping to enhance the empirical robustness of the proposed algorithms, but fail to provide any theoretical guarantees. Recent work [35] has proposed clipping of the gradient differences, in conjunction with variance reduction, to tackle partial worker participation when adversarial workers can form a majority. Our focus is however different, as we consider the impact of pre-aggregation clipping on the robustness obtained by a general class of aggregation rules.
Lastly, prior work [19, 48, 5] has also considered the challenge of privacy preservation alongside robustness in distributed learning. In this line of work, the gradients are assumed to have bounded norms in order to control the sensitivity of the algorithm, as is common in differential privacy (DP) [1]. This is often ensured in practice by clipping the gradients using a static clipping parameter. Our findings suggest that it is better to use adaptive clipping instead of static clipping. Nevertheless, thoroughly analysing ARC in the context of DP remains an interesting research direction.
Paper outline.
Section 2 presents the problem statement. Section 3 presents the background on robust distributed learning, while Section 4 introduces ARC and its useful properties. Section 5 presents the convergence analysis of Robust-DGD with ARC, and characterizes the improvement in its robustness. Section 6 presents our empirical results. Lastly, Section 7 presents concluding remarks. Appendix A presents the computational complexity of ARC. Appendices B and C contain proofs for results presented in Sections 4 and 5, respectively. Appendix D describes the experimental setup, while Appendices E and F include additional experimental results.
2 Problem Statement
We consider the problem of distributed learning in the server-based architecture. The system comprises $n$ workers represented by $w_{1},\dots,w_{n}$, that collaborate with the help of a trusted server. The workers hold local datasets $\mathcal{D}_{1},\dots,\mathcal{D}_{n}$ respectively, each composed of $m$ data points from an input space $\mathcal{Z}$. Specifically, for any $i\in[n]$, $\mathcal{D}_{i}\coloneqq\{z_{1}^{(i)},\dots,z_{m}^{(i)}\}\subset\mathcal{Z}^{m}$. For a given model parameterized by vector $\theta\in\mathbb{R}^{d}$, $d$ being the number of trainable parameters in the model, each worker $w_{i}$ incurs a loss given by the loss function $\mathcal{L}_{i}{(\theta)}\coloneqq\frac{1}{m}\sum_{k=1}^{m}\ell(\theta,z_{k}^{(i)})\enspace$, where $\ell:\mathbb{R}^{d}\times\mathcal{Z}\rightarrow\mathbb{R}$ is the point-wise loss. We make the following standard assumptions [8]: (i) the point-wise loss function $\ell$ is differentiable with respect to $\theta$. (ii) For all $i\in[n]$, $\mathcal{L}_{i}$ is $L$-Lipschitz smooth, i.e., there exists $L\in\mathbb{R}^{+}$ such that all $\theta,\theta^{\prime}\in\mathbb{R}^{d}$, $\left\lVert{\nabla\mathcal{L}_{i}(\theta{})-\nabla\mathcal{L}_{i}(\theta{}^{\prime})}\right\rVert\leq L\left\lVert{\theta{}-\theta{}^{\prime}}\right\rVert$. In the ideal setting where all workers are assumed to be honest, i.e., follow the prescribed algorithm correctly, the server aims to minimize the global average loss function given by $\mathcal{L}(\theta{})\coloneqq\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}_{i}(\theta{})$. However, this learning objective is rendered vacuous when some workers could be adversarial, described in the following.
Robust distributed learning. A robust distributed learning algorithm aims to output a good model despite the presence of some adversarial (a.k.a., Byzantine [32]) workers in the system [40, 44, 3]. Specifically, the goal is to design a distributed learning algorithm that can tolerate up to $f$ adversarial workers, of a priori unknown identity, out of $n$ workers. Adversarial workers need not follow a prescribed algorithm, and can send arbitrary information to the server. In the context of distributed gradient descent, adversarial workers can send incorrect gradients to the server [6, 2, 43, 27]. We let $\mathcal{H}\subseteq[n]$, with $\left\lvert{\mathcal{H}}\right\rvert=n-f$, denote the set of honest workers that do not deviate from an algorithm. The objective of the server is to minimize the honest workers’ average loss function given by
$\mathcal{L}_{\mathcal{H}}(\theta)\coloneqq\frac{1}{\left\lvert{\mathcal{H}}\right\rvert}\sum_{i\in\mathcal{H}}\mathcal{L}_{i}(\theta),~{}\forall\theta\in\mathbb{R}^{d}.$
While we can find a minimum of $\mathcal{L}_{\mathcal{H}}(\theta)$ when $\mathcal{L}_{\mathcal{H}}$ is convex, in general however the loss function is non-convex, and the optimization problem is NP-hard [9]. Therefore, we aim to find a stationary point of $\mathcal{L}_{\mathcal{H}}$ instead, i.e., $\theta^{*}$ such that $\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}(\theta^{\ast})}\right\rVert=0$. Formally, we define robustness to adversarial workers by $(f,\epsilon)$-resilience [3].
Definition 2.1.
A distributed learning algorithm $\mathcal{A}$ is said to be $(f,\varepsilon)$-resilient if, despite the presence of $f$ adversarial workers, $\mathcal{A}$ enables the server to output a model $\hat{\theta}$ such that $\mathop{{}\mathbb{E}}\left[{\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}\left(\hat{\theta}\right)}\right\rVert^{2}}\right]\leq\varepsilon$, where the expectation $\mathop{{}\mathbb{E}}\left[{\cdot}\right]$ is over the randomness of the algorithm.
If a distributed learning algorithm $\mathcal{A}$ is $(f,\varepsilon)$-resilient then it can tolerate up to $f$ adversarial workers. Thus, $f/n$ is referred to as the maximum tolerable fraction of adversarial workers for $\mathcal{A}$. It is standard in robust distributed learning to design an algorithm using $f$ as a parameter [40, 7, 44, 27, 17, 3].
Data heterogeneity. In the context of distributed learning, data heterogeneity is characterized by the following standard notion of $(G,B)$-gradient dissimilarity [41, 28, 27, 17, 4].
Definition 2.2.
Loss functions $\mathcal{L}_{i},\;i\in\mathcal{H},$ are said to satisfy $(G,B)$-gradient dissimilarity if,
$\frac{1}{\left\lvert{\mathcal{H}}\right\rvert}\sum_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}(\theta)-\nabla\mathcal{L}_{\mathcal{H}}(\theta)}\right\rVert^{2}\leq G^{2}+B^{2}\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}(\theta)}\right\rVert^{2},\quad\forall\theta\in\mathbb{R}^{d}.$
3 Relevant Background
We present in this section useful precursors on robust distributed learning.
General limitations on robustness. Note that $(f,\varepsilon)$-resilience is impossible (for any $\varepsilon$) when $f/n\geq 1/2$ [20]. Moreover, even for $f/n<1/2$, it is generally impossible to achieve $(f,\varepsilon)$-resilience for arbitrarily small $\varepsilon$ due to the disparity amongst the local datasets $\mathcal{D}_{i},\;i\in\mathcal{H}$ (a.k.a., data heterogeneity) [33]. Henceforth, we assume that $n>2f$ by default. Under $(G,B)$-gradient dissimilarity, we have the following lower bound on the achievable resilience.
Lemma 3.1 (Non-convex extension of Theorem 1 in [4]).
Under $(G,B)$-gradient dissimilarity, a distributed learning algorithm $\mathcal{A}$ satisfies $(f,\varepsilon)$-resilience only if $\frac{f}{n}<\frac{1}{2+B^{2}}$ and $\varepsilon\geq\frac{1}{4}\cdot\frac{f}{n-(2+B^{2})f}G^{2}$.
For a given distributed learning problem, the minimum fraction of adversarial workers that renders any distributed learning algorithm ineffective is referred the breakdown point [18, 4]. Thus, according to Lemma 3.1, there exists a distributed learning problem satisfying $(G,B)$-gradient dissimilarity whose breakdown point is given by $1/(2+B^{2})$.
In the ideal setting, i.e., when all the workers are honest, the server can execute the classic distributed gradient descent (DGD) method in collaboration with the workers to find a stationary point of the global average loss function. In DGD, the server maintains the model, which is updated periodically using the average of the gradients computed by the workers on their local datasets. The averaging of the gradients is rendered ineffective when some workers are adversarial, i.e., they can send malicious vectors posing as their gradients to the server [18]. To remedy this, averaging is replaced with robust aggregation, e.g., coordinate-wise trimmed mean (CWTM) and median (CWMed) [44], geometric median (GM) [10], and multi-Krum (MK) [7]. Useful properties of a resulting Robust-DGD method, which is summarized in Algorithm 1, are recalled below.
Robust distributed gradient descent (Robust-DGD). In Robust-DGD, robust aggregation protects the learning procedure from incorrect gradients sent by the adversarial workers. The robustness of an aggregation can be quantified by the following property of $(f,\kappa)$-robustness, proposed in [3].
Algorithm 1 Robust Distributed Gradient Descent (Robust-DGD)
  Initialization: Server chooses a model ${{\theta}_{1}}\in\mathbb{R}^{d}$, a learning rate $\gamma\in\mathbb{R}^{+}$ and a robust aggregation rule $\mathbf{F}:\mathbb{R}^{n\times d}\to\mathbb{R}^{d}$.
  for $t=1$ to $T$ do
     Server broadcasts ${{\theta}_{t}}$ to all workers.
     for each honest worker ${{w}_{i}}$ in parallel do
        Compute local gradient ${{g}^{\left({i}\right)}_{t}}\coloneqq\nabla{\mathcal{L}_{i}({{\theta}_{t}})}$, and send ${{g}^{\left({i}\right)}_{t}}$ to Server.
        // An adversarial worker ${{w}_{j}}$ can send an arbitrary vector in $\mathbb{R}^{d}$ for ${{g}^{\left({j}\right)}_{t}}$
     end for
     Server computes $R_{t}\coloneqq\mathbf{F}\left({{g}^{\left({1}\right)}_{t}},\dots,\,{{g}^{\left({n}\right)}_{t}}\right)$.
     Server computes the updated model ${{\theta}_{t+1}}\coloneqq{{\theta}_{t}}-\gamma\,R_{t}$
  end for
  Output: Server outputs $\hat{{{\theta}}}$ chosen uniformly at random from $\{{{\theta}_{1}},\ldots,\,{{\theta}_{T}}\}$.
Definition 3.2.
Aggregation rule ${\mathbf{F}:\mathbb{R}^{n\times d}\to\mathbb{R}^{d}}$ is said to be $(f,\kappa)$-robust if there exists $\kappa\in\mathbb{R}$ such that, for all $x_{1},\ldots,x_{n}\in\mathbb{R}^{d}$ and any set $S\subseteq[n]$, $\left\lvert{S}\right\rvert=n-f$, the following holds true:
$\left\lVert{\mathbf{F}(x_{1},\ldots,x_{n})-\overline{x}_{S}}\right\rVert^{2}\leq\frac{\kappa}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}\enspace,\quad\text{where }~{}\overline{x}_{S}=\frac{1}{\left\lvert{S}\right\rvert}\sum\limits_{i\in S}x_{i}.$
We refer to $\kappa$ as the robustness coefficient.
The property of $(f,\kappa)$-robustness encompasses several robust aggregations, including the ones mentioned above and more (e.g., see [3]).222$(f,\kappa)$-robustness unifies other robustness definitions: $(\delta_{\text{max}},c)$-robustness [26] and $(f,\lambda)$-resilience [14]. It has been shown that an aggregation rule is $(f,\kappa)$-robust only if $\kappa\geq\frac{f}{n-2f}$ [3]. Importantly, the asymptotic error of Robust-DGD with $(f,\kappa)$-robust aggregation, where $\kappa\in\mathcal{O}(f/(n-2f))$, matches the lower bound (recalled in Lemma 3.1[4]. Note that all the aforementioned aggregation rules, i.e., CWTM, CWMed, GM and MK, attain optimal robustness, i.e., are $(f,\mathcal{O}\left(f/(n-2f)\right))$-robust, when composed with the pre-aggregation step: nearest neighbor mixing (NNM) [3]. [^3] CWTM has been shown to be optimal even without NNM [3]. Specifically, we recall the following result.
Lemma 3.3 (Lemma 1 in [3]).
For $\mathbf{F}\in\left\{\mathbf{CWTM},\mathbf{CWMed},\mathbf{GM},\mathbf{MK}\right\}$, the composition $\mathbf{F}\circ\mathbf{NNM}$ is $(f,\kappa)$-robust with
$\kappa\leq\frac{8f}{n-f}\left(1+\left(1+\frac{f}{n-2f}\right)^{2}\right).$
Thus, if $n\geq(2+\delta)f$ for $\delta>0$, then $\mathbf{F}\circ\mathbf{NNM}$ is $(f,\kappa)$-robust with $\kappa\leq\frac{16f}{n-f}\left(\frac{\delta+1}{\delta}\right)^{2}\in\mathcal{O}\left(\frac{f}{n-2f}\right)$.
Convergence of Robust-DGD. Lastly, we recall the convergence result for Robust-DGD with an $(f,\kappa)$-robust aggregation $\mathbf{F}$. We let $\mathcal{L}_{\mathcal{H}}^{*}$ denote the minimum value of $\mathcal{L}_{\mathcal{H}}({{\theta}})$.
Lemma 3.4 (Theorem 2 in [4]).
Consider Algorithm 1 with $\gamma\leq 1/L$. Let $\Delta_{o}\in\mathbb{R}^{+}$ such that $\mathcal{L}_{\mathcal{H}}({{\theta}_{1}})-\mathcal{L}_{\mathcal{H}}^{*}\leq\Delta_{o}$. If the aggregation rule $\mathbf{F}$ is $(f,\kappa)$-robust with $\kappa B^{2}<1$, then
$\frac{1}{T}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}({{\theta}_{t}})}\right\rVert^{2}\leq\frac{2\Delta_{o}}{(1-\kappa B^{2})\gamma T}+\frac{\kappa G^{2}}{1-\kappa B^{2}}.$
Thus, when $\kappa\in\mathcal{O}\left(f/(n-2f)\right)$ (and smaller than $1/B^{2}$), Robust-DGD is optimal, i.e., its error matches the lower bound recalled in Lemma 3.1, as soon as $T\geq\Delta_{o}/\gamma\kappa G^{2}$. [^4] For $f=0$, Lemma 3.1 recovers the convergence of DGD, provided that $\kappa$ is tight, i.e., $\kappa\in\mathcal{O}(f/(n-2f))$.
4 Adaptive Robust Clipping (ARC) and its Properties
In this section, upon presenting a preliminary observation on the fragility of static clipping, we introduce adaptive robust clipping (i.e., ARC). Lastly, we present some useful properties of ARC.
For a clipping parameter $C\in\mathbb{R}^{+}$ and a vector $x\in\mathbb{R}^{d}$, we denote $\mathrm{clip}_{C}(x)\coloneqq\min\left(1,\frac{C}{\left\lVert{x}\right\rVert}\right)x$. For a set of $n$ vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$, we denote $\mathbf{Clip}_{C}(x_{1},\dots,x_{n})\coloneqq(\mathrm{clip}_{C}(x_{1}),\dots,\mathrm{clip}_{C}(x_{n}))\enspace.$
Let $\mathbf{F}$ be an $(f,\kappa)$-robust aggregation. Given a set of $n$ vectors $x_{1},\ldots,x_{n}$, let $\mathbf{F}\circ\mathbf{Clip}_{C}(x_{1},\dots,x_{n})\coloneqq\mathbf{F}\left(\mathbf{Clip}_{C}(x_{1},\dots,x_{n})\right)$. We make the following observation.
Lemma 4.1.
For any fixed $C\in\mathbb{R}^{+}$ and $\kappa^{\prime}\geq 0$, ${\mathbf{F}\circ\mathbf{Clip}_{C}}$ is not $(f,\kappa^{\prime})$-robust.
This means that if the clipping threshold is a priori fixed, i.e., independent from the input vectors, pre-aggregation clipping does not preserve the robustness of the original aggregation. This fragility of such static clipping is also apparent in practice, as shown by our experimental study in Appendix F.
Description of ARC. In ARC, the clipping threshold is determined according to the input vectors. More precisely, ARC clips the largest $k=\left\lfloor{2(f/n)(n-f)}\right\rfloor$ vectors using a clipping parameter given by the norm of the $(k+1)$-th largest input vector. The overall scheme is formally presented in Algorithm 2, and its computational complexity is $\mathcal{O}(nd+n\log(n))$ (see Appendix A). ARC is adaptive in the sense that the clipping threshold $C$ is not fixed but depends on the input vectors.
Algorithm 2 Adaptive Robust Clipping (ARC)
  Input: $f$ and $x_{1},\ldots,x_{n}\in\mathbb{R}^{d}$.
  Find a permutation $\pi:[n]\to[n]$ such that $\left\lVert{x_{\pi(1)}}\right\rVert\geq\left\lVert{x_{\pi(2)}}\right\rVert\ldots\geq\left\lVert{x_{\pi(n)}}\right\rVert$.
  Set $k=\left\lfloor{2\frac{f}{n}(n-f)}\right\rfloor$ and $C=\left\lVert{x_{\pi(k+1)}}\right\rVert$.
  Output: $\mathbf{Clip}_{C}(x_{1},\ldots,x_{n})$
Properties of ARC. We present below important properties of ARC. Let $\mathbf{F}$ be an $(f,\kappa)$-robust aggregation rule and $\mathbf{F}\circ\mathbf{ARC}(x_{1},\dots,x_{n})\coloneqq\mathbf{F}(\mathbf{ARC}(x_{1},\dots,x_{n}))$
Theorem 4.2.
If $\mathbf{F}$ is $(f,\kappa)$-robust, then $\mathbf{F}\circ\mathbf{ARC}$ is $\left(f,\kappa+\frac{2f}{n-2f}\right)$-robust.
Since $\kappa\geq\frac{f}{n-2f}$ (recalled in Section 3), Theorem 4.2 implies that $\mathbf{F}\circ\mathbf{ARC}$ is $\left(f,3\kappa\right)$-robust. In other words, ARC preserves the robustness of the original aggregation scheme.
Besides preserving robustness in the context of Robust-DGD, ARC also limits the norms of adversarial workers’ gradients by the norm of an honest worker’s gradient. Specifically, when $f/n<1/2$, the clipping threshold $C$ in ARC is bounded by $\max_{i\in\mathcal{H}}\left\lVert{x_{i}}\right\rVert$ (see Appendix C.2). This property is useful in proving our convergence result, stated below in Theorem 5.1. Proofs for Lemma 4.1 and Theorem 4.2 are deferred to Appendix B.
5 Learning Guarantee of Robust-DGD with ARC
In this section, we first present our convergence result for Robust-DGD with ARC (or Robust-DGD $\circ$ ARC), and then characterize its improvement over the lower bound recalled in Lemma 3.1. Specifically, throughout this section, we consider Algorithm 1 with aggregation $\mathbf{F}\circ\mathbf{ARC}$, i.e., in each learning step $t$, $R_{t}\coloneqq\mathbf{F}\circ\mathbf{ARC}\left({{g}^{\left({1}\right)}_{t}},\dots,\,{{g}^{\left({n}\right)}_{t}}\right)$ Note that, since we have already shown in Theorem 4.2 that $\mathbf{F}\circ\mathbf{ARC}$ is $(f,3\kappa)$-robust, a general convergence result for Robust-DGD $\circ$ ARC follows verbatim from Lemma 3.4, replacing $\kappa$ with $3\kappa$. However, we obtain below an alternative convergence result that enables us to show the improvement in learning error due to ARC.
5.1 Convergence of Robust-DGD with ARC
We assume that, for any set of vectors $x_{1},\ldots,x_{n}\in\mathbb{R}^{d}$, $\left\lVert{\mathbf{F}\left(x_{1},\ldots,x_{n}\right)}\right\rVert\leq\max_{i\in[n]}\left\lVert{x_{i}}\right\rVert$. This assumption can be made without loss of generality, the reason for which is deferred to Appendix C.1.
Theorem 5.1.
Let $\mathbf{F}$ be $(f,\kappa)$-robust. Let $\Delta_{o}$ and $\rho$ be real values such that $\mathcal{L}_{\mathcal{H}}({{\theta}_{1}})-\mathcal{L}_{\mathcal{H}}^{*}\leq\Delta_{o}$ and $\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{1}})}\right\rVert\leq\exp\left(-\frac{\Delta_{o}}{\kappa G^{2}}L\right)\rho$. If $\gamma=\min\left\{\left(\frac{\Delta_{o}}{\kappa G^{2}}\right)\frac{1}{T},~{}\frac{1}{L}\right\}$, then
$\frac{1}{T}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}\leq\max\left\{5\kappa\left(G^{2}+B^{2}\rho^{2}\right),~{}\frac{2\Delta_{o}L}{T}+3\kappa\left(G^{2}+B^{2}\rho^{2}\right)\right\}.$
We have the following corollary of Theorem 5.1 and Lemma 3.4.
Corollary 5.2.
For parameters given in Theorem 5.1, if $T\geq\frac{\Delta_{o}L}{\kappa G^{2}}$, then Robust-DGD $\circ$ ARC achieves
$\mathop{{}\mathbb{E}}\left[{\left\lVert{\nabla\mathcal{L}\left(\hat{{{\theta}}}\right)}\right\rVert^{2}}\right]\leq 5\kappa~{}\min\left\{G^{2}+B^{2}\rho^{2},~{}\frac{G^{2}}{\max\{(1-\kappa B^{2}),0\}}\right\},$
where $\mathop{{}\mathbb{E}}\left[{\cdot}\right]$ denotes the expectation over the choice of $\hat{{{\theta}}}$.
5.2 Improvement in Robustness
In the following, we show that if the local gradients at the initial model ${{\theta}_{1}}$ are bounded, then Robust-DGD $\circ$ ARC overcomes the lower bound recalled in 3.1. Specifically, we make this assumption:
Assumption 5.3.
There exists $\zeta\in\mathbb{R}^{+}$ such that, at the initial model ${{\theta}_{1}}$, $\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}({{\theta}_{1}})}\right\rVert\leq\zeta$.
For ease of presentation, we denote $\mathsf{BP}\coloneqq\frac{1}{2+B^{2}},$
$\quad\varepsilon_{o}\left(f/n\right)\coloneqq\frac{1}{4}\cdot\frac{G^{2}(f/n)}{1-(2+B^{2})(f/n)},\quad\text{ and }$ $\Psi(G,B,\rho)\coloneqq 640\left(1+\frac{1}{B^{2}}\right)^{2}\left(1+\frac{B^{2}\rho^{2}}{G^{2}}\right).$
Recall from Lemma 3.1 that $\mathsf{BP}$ and $\varepsilon_{o}(f/n)$ are the breakdown point and the achievable stationarity error when $f/n<\mathsf{BP}$, respectively, for the worst-case scenario under $(G,B)$-gradient dissimilarity. We obtain the following theorem for Robust-DGD $\circ$ ARC.
Theorem 5.4.
Suppose Assumption 5.3 and $B>0$. Let $\mathbf{F}\in\left\{\mathbf{CWTM},\mathbf{CWMed},\mathbf{GM},\mathbf{MK}\right\}$, and let the remaining parameters for Robust-DGD $\circ$ ARC be as defined in Corollary 5.2. Let $\xi_{o}<1$ be a (strictly) positive real value, and let $\rho\coloneqq\exp\left(\frac{(2+B^{2})\Delta_{o}}{(1-\xi_{o})G^{2}}L\right)\zeta$. Then, the following holds true:
  1. 1.
    If $\frac{f}{n}\coloneqq(1-\xi)\mathsf{BP}$, where $\xi\in(0,\xi_{o}]$, then, $\mathop{{}\mathbb{E}}\left[{\left\lVert{\nabla\mathcal{L}\left(\hat{{{\theta}}}\right)}\right\rVert^{2}}\right]\leq\xi\,\Psi(G,B,\rho)\cdot\varepsilon_{o}\left(f/n\right)$
  2. 2.
    If $\mathsf{BP}\leq f/n<1/2$, then $\mathop{{}\mathbb{E}}\left[{\left\lVert{\nabla\mathcal{L}\left(\hat{{{\theta}}}\right)}\right\rVert^{2}}\right]\leq\Psi(G,B,\rho)\cdot\frac{G^{2}}{8}$
According to Theorem 5.4, if the local gradients at model initialization are bounded, then Robust-DGD $\circ$ ARC can tolerate a larger fraction of adversarial workers than the breakdown point $\mathsf{BP}$ under $(G,B)$-gradient dissimilarity. Moreover, even if $f/n<\mathsf{BP}$, Robust-DGD $\circ$ ARC can improve over the worst-case stationarity error under $(G,B)$-gradient dissimilarity, formally shown in the following corollary of Theorem 5.4. For the corollary, we consider the same parameters as in Theorem 5.4.
Corollary 5.5.
Suppose Assumption 5.3 and that $f/n\coloneqq(1-\xi)~{}\mathsf{BP}$, where $\xi\in(0,\xi_{o}]$. For $\upsilon\in(0,1)$, if $\xi\leq\frac{1}{\Psi(G,B,\rho)}\,\upsilon$, then $\mathop{{}\mathbb{E}}\left[{\left\lVert{\nabla\mathcal{L}\left(\hat{{{\theta}}}\right)}\right\rVert^{2}}\right]\leq\upsilon\cdot\varepsilon_{o}\left(f/n\right)$.
Corollary 5.5 implies that when local gradients at model initialization are bounded, Robust-DGD $\circ$ ARC improves over the lower bound $\varepsilon_{o}\left(f/n\right)$ for sufficiently small $\xi$ (or $f/n$ sufficiently close to $\mathsf{BP}$). [^5] An improvement over the lower bound automatically implies improvement over the original learning guarantee of Robust-DGD (recalled in Lemma 3.4), i.e., when model initialization is arbitrary. The closer $f/n$ is to $\mathsf{BP}$, the smaller is the factor of reduction $\upsilon$, and thus higher is the improvement. The threshold on $\xi$ approaches 00 when $B\rightarrow 0$, i.e., the space of improvement is quite narrow. This is expected, since the lower bound approaches $G^{2}f/(4(n-2f))$ in this case, which is already smaller than $5\kappa G^{2}$ (the error for Robust-DGD $\circ$ ARC in this case) for any $(f,\kappa)$-robust $\mathbf{F}$.
Reconciliation with the lower bound. When the workers’ gradients at model initialization can be arbitrarily large, the convergence guarantee of Robust-DGD $\circ$ ARC reduces to that of Robust-DGD. On the contrary, if the model initialization is good, i.e., $\zeta$ is small, then ARC improves the learning error, as soon as the tolerable fraction of adversarial workers is in the proximity of $\mathsf{BP}$. We remark that Assumption 5.3, pertaining to model initialization, is not too strong and can be satisfied in practice [15], also validated by our experimental study shown in Figure 17 of Appendix F. Effectively, under Assumption 5.3, ARC reduces the heterogeneity model to $(G+B\rho,0)$-gradient dissimilarity, with $\rho$ defined in Theorem 5.4, since $\max_{t\in[T]}\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}({{\theta}_{t}})}\right\rVert\leq\rho$ in Robust-DGD $\circ$ ARC for the parameters defined in Theorem 5.1, shown implicitly in the proof of Theorem 5.1 in Appendix C. Proofs for all results presented in this section are deferred to Appendix C.
6 Empirical Evaluation
In this section, we delve into the practical performance of ARC when incorporated in Robust-DSGD (see Algorithm 3 in Appendix D), an order-optimal stochastic variant of Robust-DGD [3]. We conduct experiments on standard image classification tasks, covering different adversarial scenarios. We empirically test four SOTA aggregation rules when pre-composed with ARC. We also contrast these outcomes against the performance of Robust-DSGD when no gradient clipping is used. Our findings underscore that clipping workers’ gradients using ARC prior to the aggregation is crucial to ensure the robustness of existing algorithms, especially in extreme scenarios, i.e., when either the data heterogeneity is high or there is a large fraction of adversarial workers to be tolerated.
We plot in Table 1, and Figures 1 and 3, the metric of worst-case maximal accuracy. For each of the five Byzantine attacks executed (see Appendix D.3), we record the maximal accuracy achieved by Robust-DSGD during the learning procedure under that attack. The worst-case maximal accuracy is thus the smallest maximal accuracy reached across the five attacks. As the adversarial attack cannot be known in advance, this metric is critical to accurately evaluate the robustness of aggregation methods, since it provides an estimate of the potential worst-case performance of the algorithm. Furthermore, we use the Dirichlet [23] distribution of parameter $\alpha$ to simulate data heterogeneity in the workers’ datasets. The comprehensive experimental setup can be found in Appendix D. In this section, we present results on MNIST [11] and CIFAR-10 [31], and defer our findings on Fashion-MNIST [42] to Appendix E.2. Finally, for the sake of clarity, we mainly consider here the aggregations CWTM and GM, and show the remaining results on CWMed and MK in Appendix E.
(a) $\alpha=0.5$
(b) $\alpha=0.1$
(c) extreme heterogeneity
Figure 2: Performance of Robust-DSGD when using ARC and without clipping on MNIST. Among $n=11$, we consider $f=1$ adversarial worker executing FOE [43], and vary the heterogeneity.
ARC boosts robustness in high heterogeneity. Figure 2 shows the performance of Robust-DSGD against the FOE [43] attack when using ARC opposed to no clipping with $f=1$ adversarial worker among $n=11$.When the heterogeneity is low (i.e., $\alpha=0.5$, see Figure 5), the performance of ARC and no clipping are comparable for both aggregations. When the heterogeneity increases ($\alpha=0.1$), the benefits of using ARC are more visible. CWTM $\circ$ NNM and GM $\circ$ NNM significantly struggle to learn (with a very large variance), while the same aggregations composed with ARC almost match the accuracy of DSGD towards the end of the training. In extreme heterogeneity, the improvement induced by our method is the most pronounced, as ARC enables both aggregations to reach a final accuracy close to 90%. Contrastingly, the same aggregations without clipping stagnate at an accuracy close to 10% throughout the training. Interestingly, only $f=1$ adversarial worker among $n=11$ (i.e., less than 10%) suffices to completely deteriorate the learning when no clipping is applied.
(a) $f=1$
(b) extreme heterogeneity
Figure 3: Worst-case maximal accuracies achieved by Robust-DSGD, with and without ARC, on heterogeneously-distributed MNIST with $10$ honest workers. Left: $f=1$ adversarial worker among $n=11$ for varying levels of heterogeneity. Right: extreme heterogeneity for varying $f$.
Varying heterogeneity for fixed $f$. Figure 3a compares the worst-case maximal accuracies achieved by Robust-DSGD when $f=1$, with ARC and without, for varying levels of heterogeneity. While CWTM $\circ$ NNM and GM $\circ$ NNM yield comparable accuracies with ARC in low heterogeneity ($\alpha\geq 0.5$), their performance significantly degrades when $\alpha$ drops below that threshold. Indeed, when $\alpha=0.1$, their accuracies drop below 65%, while ARC enables the same aggregations to maintain their accuracy at just below 98%. In extreme heterogeneity, the performance of the aggregations without clipping completely deteriorates with accuracies close to 15%. Contrastingly, ARC efficiently mitigates the Byzantine attacks, resulting in accuracies above 85% in the worst case for both aggregations. Similar plots for $f=3,5,7$ convey similar observations (see Appendix E.1). Essentially, in settings where gradient clipping is not needed (e.g., low heterogeneity or small $f$), ARC performs at least as well as no clipping. Nonetheless, ARC induces a significant improvement near the breakdown point of SOTA robust aggregation schemes (e.g., high heterogeneity or large $f$).
$\alpha=0.2$ $\alpha=0.075$
Aggregation No Clipping ARC No Clipping ARC
CWTM $\circ$ NNM 51.6 $\pm$ 5.1 67.8 $\pm$ 0.9 40.7 $\pm$ 0.5 60.5 $\pm$ 1.2
GM $\circ$ NNM 41.2 $\pm$ 3.5 67.0 $\pm$ 1.0 16.0 $\pm$ 2.3 60.0 $\pm$ 2.0
CWMed $\circ$ NNM 43.4 $\pm$ 4.2 69.4 $\pm$ 0.8 13.7 $\pm$ 12.1 62.7 $\pm$ 1.2
MK $\circ$ NNM 50.9 $\pm$ 5.1 68.7 $\pm$ 0.5 40.5 $\pm$ 0.7 59.9 $\pm$ 1.9
Table 1: Worst-case maximal accuracies (%) achieved by Robust-DSGD on heterogeneously-distributed CIFAR-10 with ARC and without. There is $f=1$ adversarial worker among $n=17$. As a baseline, DSGD ($f=0$) reaches 76.5% and 70% when $\alpha=0.2$ and 0.075, respectively.
SOTA robust aggregations fail in extreme heterogeneity. Figure 3b shows that for all values of $f\in\{1,...,5\}$, Robust-DSGD with CWTM $\circ$ NNM and GM $\circ$ NNM completely fails to learn, consistently yielding worst-case maximal accuracies close to 15%. This suggests that $f=1$ constitutes the breakdown point for these aggregations in extreme heterogeneity. However, composing them with ARC increases the breakdown point of these aggregations to $f=3$. Indeed, for $f=1$ and 2, ARC enables CWTM $\circ$ NNM and GM $\circ$ NNM to achieve accuracies greater than 85% in the worst case. However, when $f\geq 3$, their performance degrades, although CWTM $\circ$ NNM $\circ$ ARC reaches a satisfactory accuracy close to 70%. Moreover, even when the heterogeneity is not large, ARC still produces a significant improvement when the fraction of adversarial workers increases in the system. In Figure 1 of Section 1 where $\alpha=0.5$, the performances of ARC and no clipping are comparable for $f\leq 3$ for both aggregations.
Figure 4: FOE [43] on heterogeneous CIFAR-10 with $\alpha=0.075$. $f=1$ among $n=17$.
However, when $f\geq 5$, the improvement induced by ARC is much more visible. Particularly when $f=7$, ARC enables CWTM $\circ$ NNM and GM $\circ$ NNM to reach accuracies greater than 80% in the worst case, whereas the same aggregations without clipping yield accuracies below 40%. This observation confirms the raise of the breakdown point caused by ARC to robust aggregations in practice. Plots for $\alpha=0.1$ and 1 convey the same observations (see Appendix E.1).
Improved robustness on CIFAR-10. We also conduct experiments on the more challenging CIFAR-10 task with $n=17$ and $f=1$. Table 2 shows the worst-case maximal accuracies achieved by ARC and no clipping for four robust aggregations. For $\alpha=0.2$, ARC consistently outputs accuracies greater than 67% for all aggregations, while the same aggregations without clipping yield lower accuracies (with a large variance across seeds). For instance, GM $\circ$ NNM achieves 41.2% on average, i.e., 26% less than its counterpart with ARC. In the more heterogeneous setting $\alpha=0.075$, ARC enables all aggregations to reach worst-case maximal accuracies close to 60% (with a small variance). On the contrary, executing the same methods without clipping deteriorates the performance of Robust-DSGD, with GM $\circ$ NNM and CWMed $\circ$ NNM reaching 16% and 13.7% in worst-case accuracy, respectively. This can also be seen in Figure 4, where FOE completely degrades the learning when ARC is not used, despite CWTM $\circ$ NNM reaching a peak of 40% at the start of the learning.
7 Conclusion
We have introduced a pre-aggregation clipping scheme, named adaptive robust clipping (ARC), that helps to circumvent the general limitations on the robustness to adversarial workers in distributed learning, under the standard heterogeneity model of $(G,B)$-gradient dissimilarity. We have obtained formal guarantees on the improvement in the learning error of state-of-the-art robust distributed gradient descent (Robust-DGD) when incorporating ARC. We have validated our theoretical findings through an exhaustive set of experiments on benchmark image classification tasks.
Appendix A Computational Complexity of ARC.
The algorithm starts by computing the norms of all input vectors, which takes $\mathcal{O}(nd)$ time. Note that this step is also required for static clipping. The sorting operation of the norms takes $\mathcal{O}(n\log(n))$ time. Note that if $n$ is large compared to ${k=\left\lfloor{2(f/n)(n-f)}\right\rfloor}$, one can use more efficient algorithms for finding the ${(k+1)}$-th largest element, such as quick-select [22] which has an average-case complexity of $\mathcal{O}(n)$. Finally, clipping the vectors requires $\mathcal{O}(nd)$ time. Therefore, the total complexity of ARC is $\mathcal{O}(nd+n\log(n))$. Overall, the complexity is comparable to static clipping, given that the sorting (or search) is crucially dimension independent.
Appendix B Proofs for Results in Section 4
Proofs for Lemma 4.1 and Theorem 4.2 are presented in B.1 and B.2, respectively.
Notation
Let $n\in\mathbb{N}^{*}$. Given vectors $x_{1},\ldots,x_{n}\in\mathbb{R}^{d}$ and a set $S\subseteq[n]$, we denote by $\overline{x}_{S}$ the mean of the vectors in $S$ $\overline{x}_{S}=\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}x_{i}$. Given a clipping parameter $C\geq 0$, let
$y_{i}\coloneq\mathrm{clip}_{C}(x_{i})=\min\left(1,\frac{C}{\left\lVert{x_{i}}\right\rVert}\right)x_{i},$
and
$\overline{y}_{S}\coloneqq\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}y_{i}.$
We denote by $S_{c}$ the set of clipped vectors in $S$,
$S_{c}\coloneq\{i\in S,\left\lVert{x_{i}}\right\rVert>C\}.$ (1)
Recall that we denote by $\mathbf{Clip}_{C}$ be the operator such that, for $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$
$\mathbf{Clip}_{C}(x_{1},\dots,x_{n}):=(\mathrm{clip}_{C}(x_{1}),\dots,\mathrm{clip}_{C}(x_{n}))\enspace.$
Further, let $\mathbf{F}:\mathbb{R}^{n\times d}\to\mathbb{R}^{d}$. We denote by $\mathbf{F}\circ\mathbf{Clip}_{C}$ the aggregation rule that first clips the input vectors using parameter $C$, and then aggregates them using $\mathbf{F}$
$\mathbf{F}\circ\mathbf{Clip}_{C}(x_{1},\dots,x_{n}):=\mathbf{F}(\mathbf{Clip}_{C}(x_{1},\dots,x_{n}))\enspace.$
B.1 Proof of Lemma 4.1
See 4.1
Proof.
We use reasoning by contradiction to prove the lemma.
We first consider the case when $C>0$. Let $S\coloneqq\{1,\ldots,n-f\}$. Consider an arbitrary set of $n$ vectors $x_{1},\ldots,\,x_{n}$ in $\mathbb{R}^{d}$ such that $x_{1}=\ldots=x_{n-f}=\overline{x}_{S}$ and $\left\lVert{\overline{x}_{S}}\right\rVert=2C$. We assume that that ${\mathbf{F}\circ\mathbf{Clip}_{C}}$ is $(f,\kappa^{\prime})$-robust for some $\kappa^{\prime}\geq 0$. This assumption implies that
$\left\lVert{\mathbf{F}\circ\mathbf{Clip}_{C}(x_{1},\ldots,x_{n})-\overline{x}_{S}}\right\rVert^{2}$ $\leq\frac{\kappa^{\prime}}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}=0.$ (2)
Therefore,
$\mathbf{F}\circ\mathbf{Clip}_{C}(x_{1},\ldots,x_{n})=\overline{x}_{S}.$
However, the clipping operation results in $\mathrm{clip}_{C}(x_{1})=\ldots=\mathrm{clip}_{C}(x_{n-f})=\frac{1}{2}\overline{x}_{S}$. Therefore, by $(f,\kappa)$-robustness of $\mathbf{F}$,
${\mathbf{F}\circ\mathbf{Clip}_{C}}(x_{1},\ldots,x_{n})=\mathbf{F}\left(\frac{1}{2}\overline{x}_{S},\ldots,\frac{1}{2}\overline{x}_{S},\mathrm{clip}_{C}(x_{n-f+1}),\ldots,\mathrm{clip}_{C}(x_{n})\right)=\frac{1}{2}\overline{x}_{S}.$
This contradicts (2). As this contradiction holds for any value of $\kappa^{\prime}$, ${\mathbf{F}\circ\mathbf{Clip}_{C}}$ cannot be $(f,\kappa^{\prime})$-robust for any $\kappa^{\prime}$.
The proof for the case when $C=0$ is similar to the above, where we choose $x_{1}=\ldots=x_{n-f}=\overline{x}_{S}$ such that $\overline{x}_{S}$ is any vector with strictly positive norm. ∎
B.2 Proof of Theorem 4.2
Our proof relies on the following lemma, proof of which is deferred to B.2.1.
Lemma B.1.
Let $C\in\mathbb{R}^{+}$ and $\mathbf{F}$ be an $(f,\kappa)$-robust aggregation rule. Consider an arbitrary set of $n$ vectors $x_{1},x_{2},\ldots,x_{n}\in\mathbb{R}^{d}$ and an arbitrary $S\subseteq[n]$ with $\left\lvert{S}\right\rvert=n-f$. Let $S_{c}$ denote the set of indices of clipped vectors in $S$, i.e., $S_{c}\coloneqq\{i\in S,\left\lVert{x_{i}}\right\rVert>C\}$. If $\left\lvert{S\setminus S_{c}}\right\rvert\geq 1$, then
$\left\lVert{\mathbf{F}\circ\mathbf{Clip}_{C}(x_{1},\ldots,x_{n})-\overline{x}_{S}}\right\rVert^{2}\leq\frac{\tilde{\kappa}}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}\enspace,$
where $\tilde{\kappa}=\kappa+\frac{\left\lvert{S_{c}}\right\rvert}{\left\lvert{S\setminus S_{c}}\right\rvert}$.
Lemma B.1 shows that $\mathbf{F}\circ\mathbf{Clip}_{C}$ is $(f,\tilde{\kappa})$-robust, provided that $\left\lvert{S\setminus S_{c}}\right\rvert\geq 1$ for all subsets $S$ of size $n-f$. Note that this condition is impossible to guarantee when using a fixed clipping threshold that does not depend on the input vectors. In order to ensure that $\left\lvert{S\setminus S_{c}}\right\rvert\geq 1$ for all subsets $S$ of size $n-f$, the clipping threshold $C$ should be large enough such that less than $n-f$ input vectors are clipped. By construction, ARC satisfies the condition of Lemma B.1. This brings us to the proof of Theorem 4.2, which we recall below for convenience.
See 4.2
Proof.
Since we clip the largest $\left\lfloor{2(f/n)(n-f)}\right\rfloor$ gradients, for a given $S\subseteq[n]$ with $\left\lvert{S}\right\rvert=n-f$ we have
$\left\lvert{S_{c}}\right\rvert\leq\left\lfloor{2(f/n)(n-f)}\right\rfloor\leq 2(f/n)(n-f).$
Therefore,
$\left\lvert{S\setminus S_{c}}\right\rvert=\left\lvert{S}\right\rvert-\left\lvert{S_{c}}\right\rvert\geq(n-f)-\frac{2f}{n}(n-f)=(n-f)\frac{n-2f}{n}.$
Since it is assumed that $f<n/2$, we have
$\left\lvert{S\setminus S_{c}}\right\rvert>\left(n-\frac{n}{2}\right)\frac{n-n}{n}=0.$
Thus, the condition $\left\lvert{S\setminus S_{c}}\right\rvert\geq 1$ is always verified. Hence, from Lemma B.1 we obtain that $\mathbf{F}\circ\mathbf{ARC}$ is $\left(f,\left(\kappa+\frac{\left\lvert{S_{c}}\right\rvert}{\left\lvert{S\setminus S_{c}}\right\rvert}\right)\right)$-robust where
$\frac{\left\lvert{S_{c}}\right\rvert}{\left\lvert{S\setminus S_{c}}\right\rvert}\leq\frac{2(f/n)(n-f)}{(n-f)\frac{n-2f}{n}}=\frac{2f}{n-2f}.$
This concludes the proof. ∎
B.2.1 Proof of Lemma B.1
We start by giving bounds on the variance (Lemma B.2) and the bias (Lemma B.3 and Lemma B.4) due to the clipping operation. We combine these bounds to prove the theorem.
Variance reduction due to clipping
We start by giving a bound on the variance of the clipped vectors. Recall from (1) that $S_{c}\coloneqq\{i\in S,\left\lVert{x_{i}}\right\rVert>C\}$.
Lemma B.2.
Let $C\geq 0$, $n>0$ and $f<\frac{n}{2}$. For all $S\subseteq[n]$ with $\left\lvert{S}\right\rvert=n-f$, the following holds true:
  1. 1.
    If $\left\lVert{\overline{x}_{S}}\right\rVert\leq C$ then
    $\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{y_{i}-\overline{y}_{S}}\right\rVert^{2}\leq\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}-\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S_{c}}(\left\lVert{x_{i}}\right\rVert-C)^{2}.$
  2. 2.
    If $\left\lVert{\overline{x}_{S}}\right\rVert>C$ then
    $\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{y_{i}-\overline{y}_{S}}\right\rVert^{2}\leq\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}-\frac{\left\lvert{S\setminus S_{c}}\right\rvert}{\left\lvert{S}\right\rvert}(\left\lVert{\overline{x}_{S}}\right\rVert-C)^{2}-\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S_{c}}(\left\lVert{x_{i}}\right\rVert-\left\lVert{\overline{x}_{S}}\right\rVert)^{2}.$
Proof.
Note that
$\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{y_{i}-\overline{y}_{S}}\right\rVert^{2}$ $=\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{y_{i}-\overline{x}_{S}+\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}$
$=\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left(\left\lVert{y_{i}-\overline{x}_{S}}\right\rVert^{2}+\left\lVert{\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}+2\langle y_{i}-\overline{x}_{S},\overline{x}_{S}-\overline{y}_{S}\rangle\right)$
$=\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{y_{i}-\overline{x}_{S}}\right\rVert^{2}+\left\lVert{\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}+2\left\langle\underbrace{\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}y_{i}}_{\overline{y}_{S}}-\overline{x}_{S},\,\overline{x}_{S}-\overline{y}_{S}\right\rangle$
$=\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{y_{i}-\overline{x}_{S}}\right\rVert^{2}+\left\lVert{\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}-2\left\langle\overline{x}_{S}-\overline{y}_{S},\,\overline{x}_{S}-\overline{y}_{S}\right\rangle$
$=\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{y_{i}-\overline{x}_{S}}\right\rVert^{2}-\left\lVert{\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}.$ (3)
By the definition of $S_{c}$ (in (1)), for all $i\in S\setminus S_{c}$, $y_{i}=x_{i}$. Therefore,
$\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{y_{i}-\overline{x}_{S}}\right\rVert^{2}$ $=\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S\setminus S_{c}}\left\lVert{y_{i}-\overline{x}_{S}}\right\rVert^{2}+\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S_{c}}\left\lVert{y_{i}-\overline{x}_{S}}\right\rVert^{2}$
$=\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S\setminus S_{c}}\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}+\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S_{c}}\left\lVert{y_{i}-\overline{x}_{S}}\right\rVert^{2}.$
The above can be written as
$\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{y_{i}-\overline{x}_{S}}\right\rVert^{2}=\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}+\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S_{c}}(\left\lVert{y_{i}-\overline{x}_{S}}\right\rVert^{2}-\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}).$ (4)
For $i\in S_{c}$, we have $y_{i}=\frac{C}{\left\lVert{x_{i}}\right\rVert}x_{i}$. Thus, for all $i\in S_{c}$, we obtain that
$\left\lVert{y_{i}-\overline{x}_{S}}\right\rVert^{2}-\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}$ $=\underbrace{\left\lVert{y_{i}}\right\rVert^{2}}_{=C^{2}}+\left\lVert{\overline{x}_{S}}\right\rVert^{2}-2\langle y_{i},\overline{x}_{S}\rangle-\left\lVert{x_{i}}\right\rVert^{2}-\left\lVert{\overline{x}_{S}}\right\rVert^{2}+2\langle x_{i},\overline{x}_{S}\rangle$
$=C^{2}-\left\lVert{x_{i}}\right\rVert^{2}+2\left(1-\frac{C}{\left\lVert{x_{i}}\right\rVert}\right)\langle x_{i},\overline{x}_{S}\rangle$
$=-(\left\lVert{x_{i}}\right\rVert-C)(\left\lVert{x_{i}}\right\rVert+C)+2(\left\lVert{x_{i}}\right\rVert-C)\frac{\langle x_{i},\overline{x}_{S}\rangle}{\left\lVert{x_{i}}\right\rVert}$
$=(\left\lVert{x_{i}}\right\rVert-C)\left(2\frac{\langle x_{i},\overline{x}_{S}\rangle}{\left\lVert{x_{i}}\right\rVert}-\left\lVert{x_{i}}\right\rVert-C\right).$
By Cauchy-Schwarz inequality, we have $\left\langle x_{i},\,\overline{x}_{S}\right\rangle\leq\left\lVert{x_{i}}\right\rVert\left\lVert{\overline{x}_{S}}\right\rVert$ Therefore,
$\left\lVert{y_{i}-\overline{x}_{S}}\right\rVert^{2}-\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}$ $\leq(\left\lVert{x_{i}}\right\rVert-C)\left(2\left\lVert{\overline{x}_{S}}\right\rVert-\left\lVert{x_{i}}\right\rVert-C\right).$
Substituting from the above in (4), we obtain that
$\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{y_{i}-\overline{x}_{S}}\right\rVert^{2}\leq\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}+\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S_{c}}(\left\lVert{x_{i}}\right\rVert-C)\left(2\left\lVert{\overline{x}_{S}}\right\rVert-\left\lVert{x_{i}}\right\rVert-C\right).$
Substituting from the above in (3), we obtain that
$\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{y_{i}-\overline{y}_{S}}\right\rVert^{2}\leq\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}+\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S_{c}}(\left\lVert{x_{i}}\right\rVert-C)\left(2\left\lVert{\overline{x}_{S}}\right\rVert-\left\lVert{x_{i}}\right\rVert-C\right)-\left\lVert{\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}.$ (5)
We now consider below the two cases: $\left\lVert{\overline{x}_{S}}\right\rVert\leq C$ and $\left\lVert{\overline{x}_{S}}\right\rVert>C$.
In the first case, i.e,. when $\left\lVert{\overline{x}_{S}}\right\rVert\leq C$, we have
$\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S_{c}}(\left\lVert{x_{i}}\right\rVert-C)\left(2\left\lVert{\overline{x}_{S}}\right\rVert-\left\lVert{x_{i}}\right\rVert-C\right)$ $\leq\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S_{c}}(\left\lVert{x_{i}}\right\rVert-C)\left(2C-\left\lVert{x_{i}}\right\rVert-C\right)$
$\leq-\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S_{c}}(\left\lVert{x_{i}}\right\rVert-C)^{2}.$
Substituting from the above in (5) yields the following
$\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{y_{i}-\overline{y}_{S}}\right\rVert^{2}$ $\leq\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}-\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S_{c}}(\left\lVert{x_{i}}\right\rVert-C)^{2}-\left\lVert{\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}$
$\leq\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}-\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S_{c}}(\left\lVert{x_{i}}\right\rVert-C)^{2}.$
This proves the first part of the lemma.
Consider the second case, i.e., $\left\lVert{\overline{x}_{S}}\right\rVert>C$. Note that
$\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S_{c}}(\left\lVert{x_{i}}\right\rVert-C)\left(2\left\lVert{\overline{x}_{S}}\right\rVert-\left\lVert{x_{i}}\right\rVert-C\right)$ $=\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S_{c}}\left((\left\lVert{\overline{x}_{S}}\right\rVert-C)^{2}-(\left\lVert{x_{i}}\right\rVert-\left\lVert{\overline{x}_{S}}\right\rVert)^{2}\right)$
$=\frac{\left\lvert{S_{c}}\right\rvert}{\left\lvert{S}\right\rvert}(\left\lVert{\overline{x}_{S}}\right\rVert-C)^{2}-\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S_{c}}(\left\lVert{x_{i}}\right\rVert-\left\lVert{\overline{x}_{S}}\right\rVert)^{2}.$ (6)
Since $\left\lVert{\overline{x}_{S}}\right\rVert>C$, and $\left\lVert{\overline{y}_{S}}\right\rVert\leq C$, we have $\left\lVert{\overline{x}_{S}}\right\rVert-\left\lVert{\overline{y}_{S}}\right\rVert\geq\left\lVert{\overline{x}_{S}}\right\rVert-C\geq 0$ This, in conjunction with the reverse triangle inequality, implies that
$\left\lVert{\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}\geq\left(\left\lVert{\overline{x}_{S}}\right\rVert-\left\lVert{\overline{y}_{S}}\right\rVert\right)^{2}\geq(\left\lVert{\overline{x}_{S}}\right\rVert-C)^{2}.$ (7)
Substituting from (6) and (7) in (5) yields the following:
$\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{y_{i}-\overline{y}_{S}}\right\rVert^{2}$ $\leq\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}+\frac{\left\lvert{S_{c}}\right\rvert}{\left\lvert{S}\right\rvert}(\left\lVert{\overline{x}_{S}}\right\rVert-C)^{2}-\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S_{c}}(\left\lVert{x_{i}}\right\rVert-\left\lVert{\overline{x}_{S}}\right\rVert)^{2}-(\left\lVert{\overline{x}_{S}}\right\rVert-C)^{2}$
$\leq\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}+\left(\frac{\left\lvert{S_{c}}\right\rvert}{\left\lvert{S}\right\rvert}-1\right)(\left\lVert{\overline{x}_{S}}\right\rVert-C)^{2}-\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S_{c}}(\left\lVert{x_{i}}\right\rVert-\left\lVert{\overline{x}_{S}}\right\rVert)^{2}$
$=\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}-\frac{\left\lvert{S\setminus S_{c}}\right\rvert}{\left\lvert{S}\right\rvert}(\left\lVert{\overline{x}_{S}}\right\rVert-C)^{2}-\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S_{c}}(\left\lVert{x_{i}}\right\rVert-\left\lVert{\overline{x}_{S}}\right\rVert)^{2}.$
This proves the second part, which concludes the proof of the lemma. ∎
Bias due to clipping
We now bound the bias induced by clipping the input vectors.
Lemma B.3.
Let $C\geq 0$, $n>0$, $f<\frac{n}{2}$, and $S\subseteq[n]$, $\left\lvert{S}\right\rvert=n-f$. Then,
$\left\lVert{\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}$ $\leq\frac{\left\lvert{S_{c}}\right\rvert}{\left\lvert{S}\right\rvert^{2}}\sum_{i\in S_{c}}(\left\lVert{x_{i}}\right\rVert-C)^{2}.$
Proof.
Note that
$\left\lVert{\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}$ $=\left\lVert{\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}x_{i}-\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}y_{i}}\right\rVert^{2}=\left\lVert{\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}(x_{i}-y_{i})}\right\rVert^{2}.$
As $x_{i}=y_{i}$ for all $i\in S\setminus S_{c}$, we have
$\left\lVert{\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}$ $=\frac{1}{\left\lvert{S}\right\rvert^{2}}\left\lVert{\sum_{i\in S_{c}}(x_{i}-y_{i})}\right\rVert^{2}.$
Due to Jensen’s inequality,
$\left\lVert{\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}$ $=\frac{\left\lvert{S_{c}}\right\rvert^{2}}{\left\lvert{S}\right\rvert^{2}}\left\lVert{\frac{1}{\left\lvert{S_{c}}\right\rvert}\sum_{i\in S_{c}}(x_{i}-y_{i})}\right\rVert^{2}\leq\frac{\left\lvert{S_{c}}\right\rvert}{\left\lvert{S}\right\rvert^{2}}\sum_{i\in S_{c}}\left\lVert{x_{i}-y_{i}}\right\rVert^{2}.$
As $y_{i}=\frac{C}{\left\lVert{x_{i}}\right\rVert}x_{i}$ for all $i\in S_{c}$, substituting this in the above proves the lemma. ∎
We now show that the bias is upper bounded by a multiplicative factor of the variance of the input vectors, as long as there is at least one unclipped honest vector.
Lemma B.4 (Bias due to clipping).
$C\geq 0$, $n>0$, $f<\nicefrac{{n}}{{2}}$, and $S\subseteq[n]$, $\left\lvert{S}\right\rvert=n-f$, if $\left\lvert{S\setminus S_{c}}\right\rvert\geq 1$ then
$\left\lVert{\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}\leq\frac{\left\lvert{S_{c}}\right\rvert}{\left\lvert{S\setminus S_{c}}\right\rvert}\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}.$
Proof.
We assume throughout the proof that $\left\lvert{S_{c}}\right\rvert>0$. Otherwise, if $\left\lvert{S_{c}}\right\rvert=0$, the bias is 00 and the statement is trivially true.
By Lemma B.3, we have
$\left\lVert{\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}$ $\leq\frac{\left\lvert{S_{c}}\right\rvert}{\left\lvert{S}\right\rvert^{2}}\sum_{i\in S_{c}}(\left\lVert{x_{i}}\right\rVert-C)^{2}.$ (8)
We distinguish two cases: $\left\lVert{\overline{x}_{S}}\right\rVert\leq C$ and $\left\lVert{\overline{x}_{S}}\right\rVert>C$. In the first case we have that
$0\leq\left\lVert{x_{i}}\right\rVert-C\leq\left\lVert{x_{i}}\right\rVert-\left\lVert{\overline{x}_{S}}\right\rVert.$
Substituting the above in (8), we find
$\left\lVert{\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}\leq\frac{\left\lvert{S_{c}}\right\rvert}{\left\lvert{S}\right\rvert^{2}}\sum_{i\in S_{c}}(\left\lVert{x_{i}}\right\rVert-\left\lVert{\overline{x}_{S}}\right\rVert)^{2}\leq\frac{\left\lvert{S_{c}}\right\rvert}{\left\lvert{S}\right\rvert^{2}}\sum_{i\in S}(\left\lVert{x_{i}}\right\rVert-\left\lVert{\overline{x}_{S}}\right\rVert)^{2}.$
Using the reverse triangle inequality, we have that $(\left\lVert{x_{i}}\right\rVert-\left\lVert{\overline{x}_{S}}\right\rVert)^{2}\leq\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}$. This implies that
$\left\lVert{\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}\leq\frac{\left\lvert{S_{c}}\right\rvert}{\left\lvert{S}\right\rvert^{2}}\sum_{i\in S}(\left\lVert{x_{i}}\right\rVert-\left\lVert{\overline{x}_{S}}\right\rVert)^{2}\leq\frac{\left\lvert{S_{c}}\right\rvert}{\left\lvert{S\setminus S_{c}}\right\rvert}\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}.$
This proves the result for the first case.
Consider now the second case, i.e $\left\lVert{\overline{x}_{S}}\right\rVert>C$. Noting that we assume that $\left\lvert{S\setminus S_{c}}\right\rvert\geq 1$ and using Young’s inequality with $c=\frac{\left\lvert{S_{c}}\right\rvert}{\left\lvert{S\setminus S_{c}}\right\rvert}$, we find, for any $i\in S_{c}$,
$(\left\lVert{x_{i}}\right\rVert-C)^{2}$ $=(\left\lVert{x_{i}}\right\rVert-\left\lVert{\overline{x}_{S}}\right\rVert+\left\lVert{\overline{x}_{S}}\right\rVert-C)^{2}$
$\leq(1+c)(\left\lVert{x_{i}}\right\rVert-\left\lVert{\overline{x}_{S}}\right\rVert)^{2}+(1+1/c)(\left\lVert{\overline{x}_{S}}\right\rVert-C)^{2})$
$=\frac{\left\lvert{S_{c}}\right\rvert+\left\lvert{S\setminus S_{c}}\right\rvert}{\left\lvert{S\setminus S_{c}}\right\rvert}(\left\lVert{x_{i}}\right\rVert-\left\lVert{\overline{x}_{S}}\right\rVert)^{2}+\frac{\left\lvert{S_{c}}\right\rvert+\left\lvert{S\setminus S_{c}}\right\rvert}{\left\lvert{S_{c}}\right\rvert}(\left\lVert{\overline{x}_{S}}\right\rVert-C)^{2}.$
Recall from (1) that $S_{c}\coloneqq\{i\in S,\left\lVert{x_{i}}\right\rVert>C\}$. This implies that $\left\lvert{S_{c}}\right\rvert+\left\lvert{S\setminus S_{c}}\right\rvert=\left\lvert{S}\right\rvert$. Therefore
$(\left\lVert{x_{i}}\right\rVert-C)^{2}$ $\leq\left\lvert{S}\right\rvert\left(\frac{1}{\left\lvert{S\setminus S_{c}}\right\rvert}(\left\lVert{x_{i}}\right\rVert-\left\lVert{\overline{x}_{S}}\right\rVert)^{2}+\frac{1}{\left\lvert{S_{c}}\right\rvert}(\left\lVert{\overline{x}_{S}}\right\rVert-C)^{2}\right).$ (9)
Substituting (9) in (8), we find
$\left\lVert{\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}$ $\leq\frac{\left\lvert{S_{c}}\right\rvert}{\left\lvert{S}\right\rvert^{2}}\sum_{i\in S_{c}}\left\lvert{S}\right\rvert\left(\frac{1}{\left\lvert{S\setminus S_{c}}\right\rvert}(\left\lVert{x_{i}}\right\rVert-\left\lVert{\overline{x}_{S}}\right\rVert)^{2}+\frac{1}{\left\lvert{S_{c}}\right\rvert}(\left\lVert{\overline{x}_{S}}\right\rVert-C)^{2}\right)$
$=\frac{\left\lvert{S_{c}}\right\rvert}{\left\lvert{S}\right\rvert}\left(\frac{1}{\left\lvert{S\setminus S_{c}}\right\rvert}\sum_{i\in S_{c}}(\left\lVert{x_{i}}\right\rVert-\left\lVert{\overline{x}_{S}}\right\rVert)^{2}+(\left\lVert{\overline{x}_{S}}\right\rVert-C)^{2}\right).$ (10)
Since $\left\lVert{\overline{x}_{S}}\right\rVert>C$, we have for any $i\in S\setminus S_{c}$
$0\leq\left\lVert{\overline{x}_{S}}\right\rVert-C\leq\left\lVert{\overline{x}_{S}}\right\rVert-\left\lVert{x_{i}}\right\rVert.$
Therefore
$\left\lvert{S\setminus S_{c}}\right\rvert(\left\lVert{\overline{x}_{S}}\right\rVert-C)^{2}$ $\leq\sum_{i\in S\setminus S_{c}}(\left\lVert{\overline{x}_{S}}\right\rVert-\left\lVert{x_{i}}\right\rVert)^{2}$
Which gives the bound
$(\left\lVert{\overline{x}_{S}}\right\rVert-C)^{2}$ $\leq\frac{1}{\left\lvert{S\setminus S_{c}}\right\rvert}\sum_{i\in S\setminus S_{c}}(\left\lVert{\overline{x}_{S}}\right\rVert-\left\lVert{x_{i}}\right\rVert)^{2}.$ (11)
Substituting (11) in (10), we find
$\left\lVert{\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}$ $\leq\frac{\left\lvert{S_{c}}\right\rvert}{\left\lvert{S}\right\rvert}\left(\frac{1}{\left\lvert{S\setminus S_{c}}\right\rvert}\sum_{i\in S_{c}}(\left\lVert{x_{i}}\right\rVert-\left\lVert{\overline{x}_{S}}\right\rVert)^{2}+\frac{1}{\left\lvert{S\setminus S_{c}}\right\rvert}\sum_{i\in S\setminus S_{c}}(\left\lVert{x_{i}}\right\rVert-\left\lVert{\overline{x}_{S}}\right\rVert)^{2}\right)$
$=\frac{\left\lvert{S_{c}}\right\rvert}{\left\lvert{S\setminus S_{c}}\right\rvert}\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}(\left\lVert{x_{i}}\right\rVert-\left\lVert{\overline{x}_{S}}\right\rVert)^{2}\leq\frac{\left\lvert{S_{c}}\right\rvert}{\left\lvert{S\setminus S_{c}}\right\rvert}\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}.$
This proves the result for the second case and concludes the proof. ∎
Proof of LemmaB.1
Let us recall the lemma below.
See B.1
Proof.
We assume throughout the proof that $\left\lvert{S_{c}}\right\rvert>0$. Otherwise, the statement is trivially true by $(f,\kappa)$-robustness of $\mathbf{F}$ and the fact that the vectors in $S$ remain unchanged after the clipping operation.
Consider first the case $\kappa=0$. Since $\kappa\geq\frac{f}{n-2f}$  [^6] Proposition 6, [3], we have that $f=0$. Recall that we denote $y_{i}\coloneq\mathrm{clip}_{C}(x_{i})$ and $\overline{y}_{S}\coloneqq\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}y_{i}$. Since $\kappa=0$, the output of $\mathbf{F}$ will correspond to the average of its input, which implies that
$\left\lVert{\mathbf{F}\circ\mathbf{Clip}_{C}(x_{1},\ldots,x_{n})-\overline{y}_{S}}\right\rVert^{2}=\left\lVert{\mathbf{F}(y_{1},\ldots,y_{n})-\overline{y}_{S}}\right\rVert^{2}=0.$
Therefore
$\mathbf{F}\circ\mathbf{Clip}_{C}(x_{1},\ldots,x_{n})=\overline{y}_{S}.$
We find then
$\left\lVert{\mathbf{F}\circ\mathbf{Clip}_{C}(x_{1},\ldots,x_{n})-\overline{x}_{S}}\right\rVert^{2}=\left\lVert{\overline{y}_{S}-\overline{x}_{S}}\right\rVert^{2}.$
The result for the case $\kappa=0$ follows from lemma B.4.
Suppose now that $\kappa>0$. Using Young’s inequality with $c=\frac{\left\lvert{S_{c}}\right\rvert}{\kappa\left\lvert{S\setminus S_{c}}\right\rvert}$ we find
$\left\lVert{\mathbf{F}\circ\mathbf{Clip}_{C}(x_{1},\ldots,x_{n})-\overline{x}_{S}}\right\rVert^{2}$ $=\left\lVert{\mathbf{F}\circ\mathbf{Clip}_{C}(x_{1},\ldots,x_{n})-\overline{y}_{S}+\overline{y}_{S}-\overline{x}_{S}}\right\rVert^{2}$
$\leq\left(1+c\right)\left\lVert{\mathbf{F}\circ\mathbf{Clip}_{C}(x_{1},\ldots,x_{n})-\overline{y}_{S}}\right\rVert^{2}+\left(1+\frac{1}{c}\right)\left\lVert{\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}.$ (12)
On the one hand, by $(f,\kappa)$-robustness of $\mathbf{F}$, we have
$(1+c)\left\lVert{\mathbf{F}\circ\mathbf{Clip}_{C}(x_{1},\ldots,x_{n})-\overline{y}_{S}}\right\rVert^{2}$ $=(1+c)\left\lVert{\mathbf{F}(y_{1},\ldots,y_{n})-\overline{y}_{S}}\right\rVert^{2}$
$\leq(1+c)\frac{\kappa}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{y_{i}-\overline{y}_{S}}\right\rVert^{2}$
$=\left(\kappa+\frac{\left\lvert{S_{c}}\right\rvert}{\left\lvert{S\setminus S_{c}}\right\rvert}\right)\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{y_{i}-\overline{y}_{S}}\right\rVert^{2}.$ (13)
On the other hand, we have
$(1+\nicefrac{{1}}{{c}})\left\lVert{\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}=\left(1+\frac{\kappa\left\lvert{S\setminus S_{c}}\right\rvert}{\left\lvert{S_{c}}\right\rvert}\right)\left\lVert{\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}=\left(\kappa+\frac{\left\lvert{S_{c}}\right\rvert}{\left\lvert{S\setminus S_{c}}\right\rvert}\right)\frac{\left\lvert{S\setminus S_{c}}\right\rvert}{\left\lvert{S_{c}}\right\rvert}\left\lVert{\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}.$ (14)
Substituting (13) and (14) in (12) we obtain that
$\left\lVert{\mathbf{F}\circ\mathbf{Clip}_{C}(x_{1},\ldots,x_{n})-\overline{x}_{S}}\right\rVert^{2}$ $\leq\left(\kappa+\frac{\left\lvert{S_{c}}\right\rvert}{\left\lvert{S\setminus S_{c}}\right\rvert}\right)\left(\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{y_{i}-\overline{y}_{S}}\right\rVert^{2}+\frac{\left\lvert{S\setminus S_{c}}\right\rvert}{\left\lvert{S_{c}}\right\rvert}\left\lVert{\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}\right).$ (15)
We consider below the two cases $\left\lVert{\overline{x}_{S}}\right\rVert\leq C$ and $\left\lVert{\overline{x}_{S}}\right\rVert>C$ separately.
In the first case, we use the the first part of lemma B.2 and lemma B.3 to obtain that
$\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{y_{i}-\overline{y}_{S}}\right\rVert^{2}+\frac{\left\lvert{S\setminus S_{c}}\right\rvert}{\left\lvert{S}\right\rvert}\left\lVert{\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}$
$\leq\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}-\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S_{c}}(\left\lVert{x_{i}}\right\rVert-C)^{2}+\frac{\left\lvert{S\setminus S_{c}}\right\rvert}{\left\lvert{S}\right\rvert}\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S_{c}}(\left\lVert{x_{i}}\right\rVert-C)^{2}$
$=\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}-\frac{\left\lvert{S_{c}}\right\rvert}{\left\lvert{S}\right\rvert}\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}(\left\lVert{x_{i}}\right\rVert-C)^{2}\leq\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}.$
This proves the result for this case.
Consider the second case, i.e $\left\lVert{\overline{x}_{S}}\right\rVert>C$. Following the proof of corollary B.4 until (10) we have
$\left\lVert{\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}\leq\frac{\left\lvert{S_{c}}\right\rvert}{\left\lvert{S}\right\rvert}\left(\frac{1}{\left\lvert{S\setminus S_{c}}\right\rvert}\sum_{i\in S_{c}}(\left\lVert{x_{i}}\right\rVert-\left\lVert{\overline{x}_{S}}\right\rVert)^{2}+(\left\lVert{\overline{x}_{S}}\right\rVert-C)^{2}\right).$
Therefore,
$\frac{\left\lvert{S\setminus S_{c}}\right\rvert}{\left\lvert{S_{c}}\right\rvert}\left\lVert{\overline{x}_{S}-\overline{y}_{S}}\right\rVert^{2}$ $\leq\frac{\left\lvert{S\setminus S_{c}}\right\rvert}{\left\lvert{S}\right\rvert}\left(\frac{1}{\left\lvert{S\setminus S_{c}}\right\rvert}\sum_{i\in S_{c}}(\left\lVert{x_{i}}\right\rVert-\left\lVert{\overline{x}_{S}}\right\rVert)^{2}+(\left\lVert{\overline{x}_{S}}\right\rVert-C)^{2}\right)$
$\leq\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S_{c}}(\left\lVert{x_{i}}\right\rVert-\left\lVert{\overline{x}_{S}}\right\rVert)^{2}+\frac{\left\lvert{S\setminus S_{c}}\right\rvert}{\left\lvert{S}\right\rvert}(\left\lVert{\overline{x}_{S}}\right\rVert-C)^{2}.$
Using the above in conjunction with lemma B.2 for the case $\left\lVert{\overline{x}_{S}}\right\rVert>C$ in (15) proves the result for the second case, which concludes the proof. ∎
Appendix C Proofs for Results in Section 5
Proofs for Theorem 5.1, Corollary 5.2, Theorem 5.4 and Corollary 5.5 are presented in appendices C.2C.3C.4 and C.5, respectively.
C.1 Assuming $\left\lVert{\mathbf{F}\left(x_{1},\ldots,x_{n}\right)}\right\rVert\leq\max_{i\in[n]}\left\lVert{x_{i}}\right\rVert$ is Without Loss of Generality
Recall that we assume $\left\lVert{\mathbf{F}\left(x_{1},\ldots,x_{n}\right)}\right\rVert\leq\max_{i\in[n]}\left\lVert{x_{i}}\right\rVert$ for all set of $n$ vectors $x_{1},\ldots,x_{n}\in\mathbb{R}^{d}$. In case this is not true, we can instead use the aggregation rule $\mathbf{F}^{\dagger}$ given by
$\mathbf{F}^{\dagger}\left(x_{1},\ldots,x_{n}\right)\coloneqq\mathrm{clip}_{C}\left(\mathbf{F}\left(x_{1},\ldots,x_{n}\right)\right),~{}\text{where}~{}C=\max_{i\in[n]}\left\lVert{x_{i}}\right\rVert.$
This modification to the aggregation rule does not affect the learning guarantee. Specifically, due to the non-expansion property of $\mathrm{clip}_{C}(\cdot)$, for any non-empty set $S\subseteq[n]$, we have
$\left\lVert{\mathbf{F}^{\dagger}\left(x_{1},\ldots,x_{n}\right)-\overline{x}_{S}}\right\rVert\leq\left\lVert{\mathbf{F}\left(x_{1},\ldots,x_{n}\right)-\overline{x}_{S}}\right\rVert,$
where $\overline{x}_{S}\coloneqq\frac{1}{\left\lvert{S}\right\rvert}\sum\limits_{i\in S}x_{i}$. Thus, if $\mathbf{F}$ is $(f,\kappa)$-robust, then $\mathbf{F}^{\dagger}$ is also $(f,\kappa)$-robust. Hence, we can make the above assumption on $\left\lVert{\mathbf{F}\left(x_{1},\ldots,x_{n}\right)}\right\rVert$ without loss of generality.
C.2 Proof of Theorem 5.1
Our proof relies on the following lemma.
Lemma C.1.
For all $t\in[T]$, we obtain that
$\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{t}})}\right\rVert\leq\exp{\left(\gamma LT\right)}~{}\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{1}})}\right\rVert.$
Proof.
Consider an arbitrary $t\in[T]$. Let $(x_{1},\ldots,\,x_{n})\coloneqq\mathbf{ARC}\left({{g}^{\left({1}\right)}_{t}},\ldots,~{}{{g}^{\left({n}\right)}_{t}}\right)$. As $n>2f$ and $\left\lvert{\mathcal{H}}\right\rvert=n-f$, the clipping threshold $C$ used in ARC (see Algorithm 2) is bounded by $\max_{i\in\mathcal{H}}\left\lVert{{{g}^{\left({i}\right)}_{t}}}\right\rVert$. Therefore,
$\max_{i\in[n]}\left\lVert{x_{i}}\right\rVert\leq\max_{i\in\mathcal{H}}\left\lVert{{{g}^{\left({i}\right)}_{t}}}\right\rVert=\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{t}})}\right\rVert.$ (16)
Since we assume (without loss of generality) that, for any set of $n$ vectors $z_{1},\ldots,z_{n}\in\mathbb{R}^{d}$, $\left\lVert{\mathbf{F}\left(z_{1},\ldots,z_{n}\right)}\right\rVert\leq\max_{i\in[n]}\left\lVert{z_{i}}\right\rVert$ (see the discussion in Section 5.1), from (16) we obtain that
$\left\lVert{R_{t}}\right\rVert=\left\lVert{\mathbf{F}\circ\mathbf{ARC}\left({{g}^{\left({1}\right)}_{t}},\ldots,~{}{{g}^{\left({n}\right)}_{t}}\right)}\right\rVert\leq\max_{i\in[n]}\left\lVert{x_{i}}\right\rVert\leq\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{t}})}\right\rVert.$
Recall that ${{\theta}_{t+1}}\coloneqq{{\theta}_{t}}-\gamma R_{t}$. Thus, from above we obtain that
$\left\lVert{{{\theta}_{t+1}}-{{\theta}_{t}}}\right\rVert\leq\gamma~{}\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{t}})}\right\rVert.$ (17)
Due to $L$-Lipschitz smoothness of $\mathcal{L}_{i}({{\theta}})$ for all $i\in\mathcal{H}$, the above implies that
$\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{t+1}})-\nabla\mathcal{L}_{i}({{\theta}_{t}})}\right\rVert\leq\gamma L~{}\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{t}})}\right\rVert,\quad\forall i\in\mathcal{H}.$
This, in conjunction with the triangle inequality, implies that
$\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{t+1}})}\right\rVert\leq\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{t}})}\right\rVert+\gamma L~{}\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{t}})}\right\rVert,\quad\forall i\in\mathcal{H}.$
Therefore,
$\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{t+1}})}\right\rVert\leq\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{t}})}\right\rVert+\gamma L\max_{i\in\mathcal{H}}~{}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{t}})}\right\rVert=\left(1+\gamma L\right)\max_{i\in\mathcal{H}}~{}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{t}})}\right\rVert.$
As $t$ was chosen arbitrarily from $[T]$, the above holds true for all $t\in[T]$. For an arbitrary $\tau\in[T]$, using the inequality recursively for $t=\tau,\ldots,1$ we obtain that
$\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{\tau+1}})}\right\rVert\leq\left(1+\gamma L\right)^{\tau}\max_{i\in\mathcal{H}}~{}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{1}})}\right\rVert\leq\left(1+\gamma L\right)^{T}\max_{i\in\mathcal{H}}~{}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{1}})}\right\rVert.$
Since $\left(1+\gamma L\right)^{T}\leq\exp{(\gamma LT)}$, the above concludes the proof. ∎
We are now ready to present our proof of Theorem 5.1, recalled below.
See 5.1
Proof.
For simplicity we write $\mathcal{L}_{\mathcal{H}}$ as $\mathcal{L}$ throughout the proof.
Consider an arbitrary $t\in[T]$. Due to $L$-Lipschitz smoothness of $\mathcal{L}({{\theta}_{)}}$, we obtain that
$\mathcal{L}({{\theta}_{t+1}})\leq\mathcal{L}({{\theta}_{t}})+\left\langle{{\theta}_{t+1}}-{{\theta}_{t}},\,\nabla\mathcal{L}({{{\theta}_{t}}})\right\rangle+\frac{L}{2}\left\lVert{{{\theta}_{t+1}}-{{\theta}_{t}}}\right\rVert^{2}.$
Substituting ${{\theta}_{t+1}}={{\theta}_{t}}-\gamma R_{t}$, and using the identity: $2\left\langle a,\,b\right\rangle=\left\lVert{a}\right\rVert^{2}+\left\lVert{b}\right\rVert^{2}-\left\lVert{a-b}\right\rVert^{2}$, we obtain that
$\mathcal{L}({{\theta}_{t+1}})\leq\mathcal{L}({{\theta}_{t}})-\frac{\gamma}{2}\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}+\frac{\gamma}{2}\left\lVert{R_{t}-\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}-\frac{\gamma}{2}\left(1-\gamma L\right)\left\lVert{R_{t}}\right\rVert^{2}.$
Since $\gamma\leq\frac{1}{L}$, $\left(1-\gamma L\right)\geq 0$. Therefore, from above we obtain that
$\mathcal{L}({{\theta}_{t+1}})\leq\mathcal{L}({{\theta}_{t}})-\frac{\gamma}{2}\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}+\frac{\gamma}{2}\left\lVert{R_{t}-\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}.$
Recall that $t$ is chosen arbitrarily from $[T]$. Thus, the above holds true for all $t\in[T]$. Taking summation on both sides from $t=1$ to $t=T$ we obtain that
$\frac{\gamma}{2}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}\leq\mathcal{L}({{\theta}_{t}})-\mathcal{L}({{\theta}_{T+1}})+\frac{\gamma}{2}\sum_{t=1}^{T}\left\lVert{R_{t}-\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}.$
Multiplying both sides by $2/(\gamma T)$ we obtain that
$\frac{1}{T}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}\leq\frac{2\left(\mathcal{L}({{\theta}_{1}})-\mathcal{L}({{\theta}_{T+1}})\right)}{\gamma T}+\frac{1}{T}\sum_{t=1}^{T}\left\lVert{R_{t}-\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}.$
Note that $\mathcal{L}({{\theta}_{1}})-\mathcal{L}({{\theta}_{T+1}})=\mathcal{L}({{\theta}_{1}})-\mathcal{L}^{*}-\left(\mathcal{L}({{\theta}_{T+1}})-\mathcal{L}^{*}\right)\leq\mathcal{L}({{\theta}_{1}})-\mathcal{L}^{*}\leq\Delta_{o}$. Using this above we obtain that
$\frac{1}{T}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}\leq\frac{2\Delta_{o}}{\gamma T}+\frac{1}{T}\sum_{t=1}^{T}\left\lVert{R_{t}-\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}.$
Recall, by Theorem 4.2, that $\mathbf{F}\circ\mathbf{ARC}$ is $(f,3\kappa)$-robust. Therefore, by Definition 3.2, $\left\lVert{R_{t}-\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}\leq\frac{3\kappa}{\left\lvert{\mathcal{H}}\right\rvert}\sum_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{t}})-\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}$ for all $t$. Using this above we obtain that
$\frac{1}{T}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}\leq\frac{2\Delta_{o}}{\gamma T}+\frac{3\kappa}{T}~{}\frac{1}{\left\lvert{\mathcal{H}}\right\rvert}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{t}})-\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}.$ (18)
Note that, since $\gamma\leq\frac{\Delta_{o}}{\kappa G^{2}T}$ and $\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{1}})}\right\rVert\leq\exp\left(-\frac{\Delta_{o}L}{\kappa G^{2}}\right)\rho$, by Lemma C.1 we have
$\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{t}})}\right\rVert\leq\exp\left(\frac{\Delta_{o}L}{\kappa G^{2}}\right)\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{1}})}\right\rVert\leq\rho,\quad\forall t\in[T].$
This, in conjunction with triangle inequality, implies that
$\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert\leq\frac{1}{\left\lvert{\mathcal{H}}\right\rvert}\sum_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{t}})}\right\rVert\leq\rho,\quad\forall t\in[T].$
Therefore, under $(G,B)$-gradient dissimilarity, for all $t\in[T]$,
$\frac{1}{\left\lvert{\mathcal{H}}\right\rvert}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{t}})-\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}\leq G^{2}+B^{2}~{}\rho^{2}.$
Using this in (18) we obtain that
$\frac{1}{T}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}\leq\frac{2\Delta_{o}}{\gamma T}+3\kappa\left(G^{2}+B^{2}\rho^{2}\right).$ (19)
Consider the two cases: (i) $T\geq\frac{\Delta_{o}L}{\kappa G^{2}}$ and (ii) $T<\frac{\Delta_{o}L}{\kappa G^{2}}$. In case (i), $\gamma=\frac{\Delta_{o}}{\kappa G^{2}T}$. Using this in (19) implies that
$\frac{1}{T}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}\leq 2\kappa G^{2}+3\kappa\left(G^{2}+B^{2}\rho^{2}\right)\leq 5\kappa\left(G^{2}+B^{2}\rho^{2}\right).$ (20)
In case (ii), $\gamma=\frac{1}{L}$. Using this in (19) implies that
$\frac{1}{T}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}\leq\frac{2\Delta_{o}L}{T}+3\kappa\left(G^{2}+B^{2}\rho^{2}\right).$ (21)
Combining (20) and (21) concludes the proof. ∎
C.3 Proof of Corollary 5.2
See 5.2
Proof.
Since $T\geq\frac{\Delta_{o}L}{\kappa G^{2}}$, $\gamma=\frac{\Delta_{o}}{\kappa G^{2}T}\leq\frac{1}{L}$. Therefore, from Theorem 5.1 we obtain that
$\frac{1}{T}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}\leq 5\left(G^{2}+B^{2}\rho^{2}\right).$ (22)
In the specific case when $\kappa B^{2}<1$, as $\gamma=\frac{\Delta_{o}}{\kappa G^{2}T}$, from Lemma 3.4 we obtain that
$\frac{1}{T}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}\leq\frac{2\Delta_{o}}{(1-\kappa B^{2})\gamma T}+\frac{\kappa G^{2}}{1-\kappa B^{2}}=\frac{3\kappa G^{2}}{1-\kappa B^{2}}.$ (23)
Combining (22) and (23) we obtain that
$\frac{1}{T}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}\leq 5\kappa\min\left\{G^{2}+B^{2}\rho^{2},~{}\frac{G^{2}}{1-\kappa B^{2}}\right\}.$
Since $\mathop{{}\mathbb{E}}\left[{\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}}\right]=\frac{1}{T}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}$ (see Algorithm 1), the above concludes the proof. ∎
C.4 Proof of Theorem 5.4
Recall that we denote
$\mathsf{BP}\coloneqq\frac{1}{2+B^{2}},\quad\varepsilon_{o}\left(f/n\right)\coloneqq\frac{1}{4}\cdot\frac{G^{2}(f/n)}{1-(2+B^{2})(f/n)},\quad\text{ and }$
$\Psi(G,B,\rho)\coloneqq 640\left(1+\frac{1}{B^{2}}\right)^{2}\left(1+\frac{B^{2}\rho^{2}}{G^{2}}\right).$
We recall the theorem below for convenience.
See 5.4
Proof.
In the proof, we substitute $\mathsf{BP}=\frac{1}{2+B^{2}}$.
First, we consider $\frac{f}{n}=\frac{1-\xi}{2+B^{2}}$. Note that, due to Lemma 3.3, $\mathbf{F}$ is $(f,\kappa)$-robust with
$\kappa\leq\frac{16f}{n-f}\left(1+\frac{1-\xi}{B^{2}}\right)^{2}\leq\frac{16(1-\xi)}{1+B^{2}}\left(1+\frac{1-\xi}{B^{2}}\right)^{2}\leq\kappa_{o}(1-\xi),$
where $\kappa_{o}\coloneqq\frac{16}{1+B^{2}}\left(1+\frac{1}{B^{2}}\right)^{2}$. Also, recall from the limitations of $(f,\kappa)$-robustness in Section 3 that
$\kappa\geq\frac{f}{n-2f}\geq\frac{f}{n}\geq\frac{1-\xi_{o}}{2+B^{2}}.$
To summarize the above, we have
$\frac{1-\xi_{o}}{2+B^{2}}\leq\kappa\leq\kappa_{o}(1-\xi).$ (24)
This implies that
$\rho\coloneqq\exp\left(\frac{(2+B^{2})\Delta_{o}}{(1-\xi_{o})G^{2}}L\right)\zeta\geq\exp\left(\frac{\Delta_{o}}{\kappa G^{2}}L\right)\zeta.$
Therefore, since $\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{1}})}\right\rVert\leq\zeta$ for all $i\in\mathcal{H}$ under Assumption 5.3, in Theorem 5.1 we always satisfy the condition: $\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{1}})}\right\rVert\leq\exp\left(-\frac{\Delta_{o}}{\kappa G^{2}}L\right)\rho$, for the defined value of $\rho$ Thus, by Corollary 5.2 and (24) we obtain that
$\mathop{{}\mathbb{E}}\left[{\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}}\right]\leq 5\kappa_{o}(1-\xi)(G^{2}+B^{2}\rho^{2})=\frac{5\kappa_{o}(1-\xi)(G^{2}+B^{2}\rho^{2})}{\frac{fG^{2}}{n-(2+B^{2})f}}\,\frac{fG^{2}}{n-(2+B^{2})f}.$ (25)
Since $\frac{f}{n}=\frac{1-\xi}{2+B^{2}}$, we obtain that
$\frac{5\kappa_{o}(1-\xi)(G^{2}+B^{2}\rho^{2})}{\frac{fG^{2}}{n-(2+B^{2})f}}=5\kappa_{o}(1-\xi)(G^{2}+B^{2}\rho^{2})\frac{\xi(2+B^{2})}{G^{2}(1-\xi)}=5\kappa_{o}\left(1+\frac{B^{2}\rho^{2}}{G^{2}}\right)\left(2+B^{2}\right)\xi.$
Substituting $\kappa_{o}$ from above, we obtain that
$\frac{5\kappa_{o}(1-\xi)(G^{2}+B^{2}\rho^{2})}{\frac{fG^{2}}{n-(2+B^{2})f}}$ $=80\left(\frac{2+B^{2}}{1+B^{2}}\right)\left(1+\frac{1}{B^{2}}\right)^{2}\left(1+\frac{B^{2}\rho^{2}}{G^{2}}\right)\xi$
$\leq 160\left(1+\frac{1}{B^{2}}\right)^{2}\left(1+\frac{B^{2}\rho^{2}}{G^{2}}\right)\xi.$
Substituting from above in (25) concludes the proof for the first part of the theorem.
The second part of the theorem follows immediately from the first inequality in (25), using the fact that $\kappa_{o}(1-\xi)\leq\kappa_{o}$. This concludes the proof. ∎
C.5 Proof of Corollary 5.5
See 5.5
Proof.
The proof follows immediately from upon using $\xi\leq\frac{1}{\Psi(G,B,\rho)}\upsilon$ in Part 1 of Theorem 5.4. ∎
Appendix D Comprehensive Experimental Setup
D.1 Datasets and Heterogeneity
In our experiments, we consider three standard image classification datasets, namely MNIST [11], Fashion-MNIST [42], and CIFAR-10 [31]. To simulate data heterogeneity in our experiments, we make the honest workers sample from the datasets using a Dirichlet [23] distribution of parameter $\alpha$, as done in [24, 3, 13]. The smaller the $\alpha$, the more heterogeneous the setting. In our empirical evaluation, we set $\alpha\in\{0.1,0.5,1\}$ on (Fashion-)MNIST and $\alpha\in\{0.05,0.075,0.1,0.2,0.5\}$ on CIFAR-10 (refer to Figure 5). Furthermore, on MNIST and Fashion-MNIST, we also consider an extreme heterogeneity setting where the datapoints are sorted by increasing labels (0 to 9) and sequentially split equally among the honest workers.
The input images of MNIST are normalized with mean $0.1307$ and standard deviation $0.3081$, while the images of Fashion-MNIST are horizontally flipped. Moreover, CIFAR-10 is expanded with horizontally flipped images, followed by a per channel normalization with means 0.4914, 0.4822, 0.4465 and standard deviations 0.2023, 0.1994, 0.2010.
(a) $\alpha=0.1$
(b) $\alpha=0.5$
(c) $\alpha=1$
(d) $\alpha=0.075$
(e) $\alpha=0.2$
Figure 5: Distribution of labels across honest workers on MNIST (row 1) and CIFAR-10 (row 2).
D.2 Algorithm, Distributed System, ML Models, and Hyperparameters
We perform our experiments using Robust-DSGD (see Algorithm 3), which is known to be an order-optimal robust variant of Robust-DGD [3]. On MNIST and Fashion-MNIST, we execute Robust-DSGD in a distributed system composed of $n-f=10$ honest workers, and $f\in\{1,2,3,4,5,7,9\}$ adversarial workers. Furthermore, we train a convolutional neural network (CNN) of 431,080 parameters with batch size $b=25$, $T=1000$, $\gamma=0.1$, and momentum parameter $\beta=0.9$. Moreover, the negative log likelihood (NLL) loss function is used, along with an $\ell_{2}$-regularization of $10^{-4}$. On CIFAR-10, we execute Robust-DSGD using a CNN of 1,310,922 parameters, in a distributed system comprising $n-f=16$ honest workers and $f=1$ adversarial worker. We set $b=50$, $T=2000$, $\beta=0.9$, and $\gamma=0.05$ decaying once at step 1500. Finally, we use the NLL loss function with an $\ell_{2}$ regularization of $10^{-2}$.
Algorithm 3 Robust Stochastic Distributed Gradient Descent (Robust-DSGD)
  Input: Server chooses an initial model ${{\theta}_{1}}\in\mathbb{R}^{d}$, learning rate $\gamma\in\mathbb{R}^{+}$, robust aggregation $\mathbf{F}:\mathbb{R}^{n\times d}\to\mathbb{R}^{d}$, momentum parameter $\beta\in\mathbb{R}^{+}$. Each honest worker ${{w}_{i}}$ sets an initial momentum $m_{0}^{(i)}=0\in\mathbb{R}^{d}$.
  for $t=1$ to $T$ do
     Server broadcasts ${{\theta}_{t}}$ to all workers.
     for each honest worker ${{w}_{i}}$ in parallel do
        Sample a random data point $z^{(i)}$ uniformly from $\mathcal{D}_{i}$.
        Compute stochastic gradient $g_{t}^{(i)}\coloneqq\nabla{\ell(\theta_{t},z^{(i)})}$.
        Update momentum: $m_{t}^{(i)}\coloneqq(1-\beta)g_{t}^{(i)}+\beta m_{t-1}^{(i)}$.
        Send $m_{t}^{(i)}$ to Server.
         // An adversarial worker ${{w}_{j}}$ can send an arbitrary vector in $\mathbb{R}^{d}$ for $m_{t}^{(j)}$
     end for
     Server computes $R_{t}\coloneqq\mathbf{F}\left(m_{t}^{(1)},\dots,\,m_{t}^{(n)}\right)$.
     Server updates the model: ${{\theta}_{t+1}}\coloneqq{{\theta}_{t}}-\gamma\,R_{t}$
  end for
  Output: Server outputs $\hat{{{\theta}}}$ chosen uniformly at random from $\{{{\theta}_{1}},\ldots,\,{{\theta}_{T}}\}$.
In order to present the architectures of the ML models used in our experiments, we adopt the following compact terminology introduced in [3].
L(#outputs) represents a fully-connected linear layer, R stands for ReLU activation, S stands for log-softmax, C(#channels) represents a fully-connected 2D-convolutional layer (kernel size 5, padding 0, stride 1), M stands for 2D-maxpool (kernel size 2), B stands for batch-normalization, and D represents dropout (with fixed probability 0.25).
The comprehensive experimental setup, as well as the architecture of the models, are presented in Table 2.
Dataset (Fashion-)MNIST CIFAR-10
Data heterogeneity $\alpha\in\{0.1,0.5,1\}$ and extreme $\alpha\in\{0.05,0.075,0.1,0.2,0.5\}$
Model type CNN CNN
Model architecture C(20)-R-M-C(20)-R-M-L(500)-R-L(10)-S (3,32×32)-C(64)-R-B-C(64)-R-B-M-D-C(128)-R-B-C(128)-R-B-M-D-L(128)-R-D-L(10)-S
Number of parameters 431,080 1,310,922
Loss NLL NLL
$\ell_{2}$-regularization $10^{-4}$ $10^{-2}$
Number of steps $T=1000$ $T=2000$
Learning rate $\gamma=0.1$ $\gamma_{t}=\begin{cases}0.05&t\leq 1500\\ 0.00082&1500<t\leq 2000\end{cases}$
Momentum parameter $\beta=0.9$ $\beta=0.9$
Batch size $b=25$ $b=50$
Honest workers $n-f=10$ $n-f=16$
Adversarial workers $f\in\{1,2,3,4,5,7,9\}$ $f=1$
Table 2: Experimental setup on (Fashion-)MNIST and CIFAR-10
D.3 Byzantine Attacks
In our experiments, the adversarial workers execute five state-of-the-art adversarial attacks from the Byzantine ML literature, namely sign-flipping (SF) [2], label-flipping (LF) [2], mimic [27], fall of empires (FOE) [43], and a little is enough (ALIE) [6].
The exact functionality of the attacks is detailed below. In every step $t$, let $\overline{m}_{t}$ be an estimation of the true honest momentum at step $t$. In our experiments, we estimate $\overline{m}_{t}$ by averaging the momentums sent by the honest workers in step $t$ of Robust-DSHB. In other words, $\overline{m}_{t}=\frac{1}{|\mathcal{H}|}\sum\limits_{i\in\mathcal{H}}m_{t}^{(i)}$, where $m_{t}^{(i)}$ is the momentum computed by honest worker $w_{i}$ in step $t$.
  • SF: the adversarial workers send the vector $-\overline{m}_{t}$ to the server.
  • LF: the adversarial workers compute their gradients on flipped labels, and send the flipped gradients to the server. Since the original labels $l$ for (Fashion-)MNIST and CIFAR-10 are in $\{0,...,9\}$, the adversarial workers execute a label flip/rotation by computing their gradients on the modified labels $l^{\prime}=9-l$.
  • Mimic: the adversarial workers mimic a certain honest worker by sending its gradient to the server. In order to determine the optimal honest worker for the adversarial workers to mimic, we use the heuristic in [27].
  • FOE: the adversarial workers send $(1-\tau)\overline{m}_{t}$ in step $t$ to the server, where $\tau\geq 0$ is a fixed real number representing the attack factor. When $\tau=2$, this attack is equivalent to SF.
  • ALIE: the adversarial workers send $\overline{m}_{t}+\tau\sigma_{t}$ in step $t$ to the server, where $\tau\geq 0$ is a fixed real number representing the attack factor, and $\sigma_{t}$ is the coordinate-wise standard deviation of $\overline{m}_{t}$.
Since the FOE and ALIE attacks have $\tau$ as parameter, we implement in our experiments enhanced and adaptive versions of these attacks, where the attack factor is not constant and may be different in every iteration. In every step $t$, we determine the optimal attack factor $\tau_{t}$ through a grid search over a predefined range of values. More specifically, in every step $t$, $\tau_{t}$ takes the value that maximizes the damage inflicted by the adversarial workers, i.e., that maximizes the $l^{2}$ norm of the difference between the average of the honest momentums $\overline{m}_{t}$ and the output of the aggregation $R_{t}$ at the server.
D.4 Benchmarking and Reproducibility
[…]
Appendix E Additional Experimental Results
In this section, we complete the experimental results that could not be placed in the main paper.
E.1 MNIST
(a) extreme
(b) $\alpha=0.5$
(c) $\alpha=0.1$
(d) $\alpha=0.1$
(e) $\alpha=1$
(f) $\alpha=1$
Figure 6: Worst-case maximal accuracies achieved by Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed MNIST with $10$ honest workers. We fix the heterogeneity level, and vary the the number of Byzantine workers $f$.
(a) $f=1$
(b) $f=3$
(c) $f=3$
(d) $f=5$
(e) $f=5$
Figure 7: Worst-case maximal accuracies achieved by Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed MNIST with $10$ honest workers. We fix the number of Byzantine workers $f$, and vary the heterogeneity level.
(a) ALIE
(b) SF
(c) LF
(d) Mimic
Figure 8: Performance of Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed MNIST (extreme heterogeneity) with $10$ honest workers and $f=1$, under several attacks. This complements Figure 2c of the main paper.
(a) FOE
(b) SF
(c) ALIE
(d) LF
(e) Mimic
Figure 9: Performance of Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed MNIST ($\alpha=0.1$) with $10$ honest workers and $f=3$, under several attacks.
(a) $f=3$
(b) $f=4$
(c) $f=5$
Figure 10: Performance of Robust-DSGD under the LF attack when using ARC compared to no clipping, on heterogeneously-distributed MNIST ($\alpha=0.1$) with $10$ honest workers and varying number of adversarial workers $f$.
E.2 Fashion-MNIST
(a) $\alpha=0.1$
(b) $\alpha=0.1$
(c) $\alpha=0.05$
(d) $\alpha=0.05$
(e) extreme
(f) extreme
Figure 11: Worst-case maximal accuracies achieved by Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed Fashion-MNIST with $10$ honest workers. We fix the heterogeneity level and vary the number of adversarial workers $f$.
(a) $f=1$
(b) $f=1$
(c) $f=3$
(d) $f=3$
(e) $f=4$
(f) $f=4$
Figure 12: Worst-case maximal accuracies achieved by Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed Fashion-MNIST with $10$ honest workers. We fix the number of adversarial workers $f$ and vary the heterogeneity level.
E.3 CIFAR-10
(a) $f=1$
(b) $f=1$
Figure 13: Worst-case maximal accuracies achieved by Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed CIFAR-10 with $16$ honest workers. We fix the number of adversarial workers $f=1$ and vary the heterogeneity level.
(a) ALIE with CWTM and GM
(b) FOE with CWMED and MK
(c) ALIE with CWMED and CWMED
Figure 14: Worst-case maximal accuracies achieved by Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed CIFAR-10 ($\alpha=0.075$) with $16$ honest workers and $f=1$ executing ALIE and FOE. This figure complements Figure 4 in the main paper.
Appendix F Static Clipping vs ARC
In this section, we present empirical results on static clipping when used as a pre-aggregation technique in Robust-DSGD, and compare its performance to our proposed adaptive algorithm ARC.
(a) Sign-flipping (SF) [2]
(b) Label-flipping (LF) [2]
Figure 15: Impact of pre-aggregation clipping, specifically static clipping with $C=2$ and ARC, on robust-DGD with aggregation scheme $\mathbf{CWTM}\circ\mathbf{NNM}$, under 2 adversarial attacks: (a) SF and (b) LF. We set $n=13$, $f=3$, and consider the MNIST [11] dataset distributed amongst $n-f=10$ honest workers with extreme heterogeneity.
F.1 Limitations of Static Clipping
We perform our experiments using Robust-DSGD (see Algorithm 3) on (Fashion-)MNIST and CIFAR-10. We consider three different levels of heterogeneity, that we call moderate (corresponding to sampling from a Dirichlet distribution of parameter $\alpha=1$), high ($\alpha=0,1$), and extreme (as explained in Appendix D). On MNIST and Fashion-MNIST, we execute Robust-DSGD in a distributed system composed of $n=15$ workers, among which $f\in\{3,4\}$ are Byzantine. Furthermore, we train a convolutional neural network of 431,080 parameters with batch size $b=25$, $T=1000$, $\gamma=0.1$, and momentum parameter $\beta=0.9$. Moreover, the negative log likelihood loss function is used, along with an $\ell_{2}$-regularization of $10^{-4}$. On CIFAR-10, we execute Robust-DSGD on ResNet-18 [21], in a distributed system comprising $n=9$ workers among which $f\in\{1,2\}$ are Byzantine. We set $b=128$, $T=2000$, $\beta=0.9$, and $\gamma=0.1$ decaying once by 10$\times$ at step 1500. Finally, we use the cross-entropy loss function with an $\ell_{2}$ regularization of $5\times 10^{-4}$.
In our experiments, we examine a wide range of static clipping parameters. Specifically, we choose $C\in\{0.02,0.2,2,20\}$ on MNIST.
(a) $f=3$
(b) $f=4$
Figure 16: Impact of the clipping strategy and heterogeneity level on the worst-case maximal accuracy achieved by Robust-DSGD against five Byzantine attacks on heterogeneous MNIST with $n=15$ workers. CWTM $\circ$ NNM is used as aggregation. DSGD reaches at least 98.5% in accuracy in all heterogeneity regimes.
Unreliability due to the attack. First, the efficacy of static clipping is significantly influenced by the type of Byzantine attack executed during the learning process. To illustrate, while $C=2$ exhibits the best performance under the SF attack [2], depicted in Figure 15, it leads to a complete collapse of learning under LF [2]. Consequently, identifying a static clipping threshold $C$ that consistently performs well in practice against any attack is difficult (if not impossible). This highlights the fragility of static clipping, as the unpredictable nature of the Byzantine attack, a parameter that cannot be a priori known, can significantly degrade the performance of the chosen static clipping approach.
Unreliability due to the heterogeneity model. Second, the robustness under static clipping is notably influenced by the level of heterogeneity present across the datasets of honest workers. To illustrate this impact, we present in Figure 16 the performances of Robust-DSGD for different levels of heterogeneity (moderate, high, and extreme) and different values of $C$. In the left plot of Figure 16, $C=2$ emerges as the best static clipping threshold when the heterogeneity is high whereas $C=20$ appears to be sub-optimal. On the other hand, under extreme heterogeneity, the accuracy associated to $C=2$ diminishes drastically while $C=20$ becomes a better choice. Intuitively, in heterogeneous scenarios, honest gradients become large in $l^{2}$-norm, due to the increasingly detrimental effect of Byzantine attacks. Consequently, we must increase the static clipping threshold. Failing to do so could introduce a significant bias. More details on this observation can be found in Figure 17. This highlights the intricate dependence of the performance of static clipping on data heterogeneity, and emphasizes the necessity to fine-tune static clipping strategies prior to the learning.
Figure 17: Distribution of the norms of honest gradients during the learning on heterogeneous MNIST. There are $f=3$ Byzantine workers (among $n=15$) executing the LF attack. CWTM $\circ$ NNM is used as aggregation method.
Unreliability due to the number of Byzantine workers. Last but not least, the number of Byzantine workers $f$ also affects the efficacy of static clipping. As seen in Figure 16, $C=2$ leads to the highest accuracy among static clipping strategies when $f=3$ in high heterogeneity, but cannot be used when $f=4$ as its corresponding accuracy drops below 50% (see right plot of Figure 16).
Overall, our empirical findings reveal that there is no single value of $C$ for which static clipping consistently delivers satisfactory performance across the various settings. We have also shown that the performance of static clipping is greatly influenced on typically uncontrollable parameters, such as data heterogeneity and the nature of the Byzantine attack. Indeed, it should be noted that explicitly estimating these parameters is challenging (if not impossible) in a distributed setting. First, since the server does not have direct access to the data, it cannot estimate the degree of heterogeneity. Second, as (by definition) a Byzantine worker can behave in an unpredictable manner, we cannot adjust $C$ to the attack(s) being executed by the Byzantine workers. This highlights the necessity for a robust clipping alternative that can naturally adapt to the setting in which it is deployed.
F.2 Empirical Benefits of ARC
In contrast to static clipping, ARC is adaptive, delivering consistent performance across diverse levels of heterogeneity, Byzantine attacks, and number of Byzantine workers.
Adaptiveness.
ARC dynamically adjusts its clipping parameter $C_{t}$ based on the norms of honest momentums at step $t$, avoiding static over-clipping or under-clipping. This adaptability is evident in Figure 18, where $C_{t}$ consistently decreases with time under all attacks, an expected behavior when reaching convergence. Moreover, Figure 18 also illustrates that any surge in the norm of the honest mean corresponds to a direct increase in $C_{t}$, highlighting the adaptive nature of ARC.
Robust performance across heterogeneity regimes.
The efficacy of our solution is illustrated in Figure 16, showcasing a consistently robust performance for all considered heterogeneity levels. Specifically, in scenarios of moderate heterogeneity where clipping may not be essential, ARC matches the performance of No clipping as well as static strategies $C=2$ and $C=20$. Conversely, the left plot of Figure 16 shows that under high and extreme heterogeneity, ARC surpasses the best static clipping strategy in terms of accuracy, highlighting the effectiveness of our approach.
Robust performance across Byzantine regimes.
ARC exhibits robust performance across diverse Byzantine scenarios, encompassing variations in both the type of Byzantine attack and the number of Byzantine workers $f$. As depicted in Figures 15 and 16, ARC consistently yields robust performance, regardless of the Byzantine attack. In extreme heterogeneity with $f=3$, ARC maintains an accuracy of 75%, while $C=2$ demonstrates a subpar worst-case accuracy of 25% (also refer to the right plot of Figure 15). Furthermore, despite the increase in the number of Byzantine workers to $f=4$, ARC remains the top-performing clipping approach among all considered strategies.
This analysis underscores the empirical superiority of ARC over static clipping methods, eliminating the reliance on data heterogeneity and the specific Byzantine regime. Similar trends are also observed in Figure 19, when executing Robust-DSGD using other aggregation methods such as CWMed $\circ$ NNM, GM $\circ$ NNM, and Multi-Krum $\circ$ NNM.
Figure 18: Evolution of the adaptive clipping parameter $C_{t}$ of ARC compared to the norm of the honest mean during the learning on heterogeneous MNIST ($\alpha=1$). There are $f=3$ Byzantine workers among $n=15$. CWTM $\circ$ NNM is used as aggregation. ARC dynamically adjusts its clipping parameter $C_{t}$ based on the norms of honest momentums at step $t$, avoiding static over-clipping or under-clipping. This adaptability is evident in under all attacks, where $C_{t}$ consistently decreases with time, an expected behavior when reaching convergence.
(a) CWMed $\circ$ NNM
(b) GM $\circ$ NNM
(c) MK $\circ$ NNM
Figure 19: Impact of the clipping strategy on the worst-case maximal accuracy achieved during the learning against five Byzantine attacks on heterogeneous MNIST, with $f=3$ and $n=15$.

### v3

@@
-Boosting Robustness by Clipping Gradients in Distributed Learning
+The Vital Role of Gradient Clipping in Byzantine-Resilient Distributed Learning
 Abstract
-Robust distributed learning consists in achieving good learning performance despite the presence of misbehaving workers.
-State-of-the-art (SOTA) robust distributed gradient descent (Robust-DGD) methods, relying on robust aggregation, have been proven to be optimal: Their learning error matches the lower bound established under the standard heterogeneity model of $(G,B)$-gradient dissimilarity.
-The learning guarantee of SOTA Robust-DGD cannot be further improved when model initialization is done arbitrarily.
-However, we show that it is possible to circumvent the lower bound, and improve the learning performance, when the workers’ gradients at model initialization are assumed to be bounded.
-We prove this by proposing pre-aggregation clipping of workers’ gradients, using a novel scheme called adaptive robust clipping (ARC).
-Incorporating ARC in Robust-DGD provably improves the learning, under the aforementioned assumption on model initialization.
-The factor of improvement is prominent when the tolerable fraction of misbehaving workers approaches the breakdown point.
-ARC induces this improvement by constricting the search space, while preserving the robustness property of the original aggregation scheme at the same time.
-We validate this theoretical finding through exhaustive experiments on benchmark image classification tasks.
+Byzantine-resilient distributed machine learning seeks to achieve robust learning performance in the presence of misbehaving or adversarial workers.
+While state-of-the-art (SOTA) robust distributed gradient descent (Robust-DGD) methods were proven theoretically optimal, their empirical success has often relied on pre-aggregation gradient clipping.
+However, the currently considered static clipping strategy exhibits mixed results: improving robustness against some attacks while being ineffective or detrimental against others.
+We address this gap by proposing a principled adaptive clipping strategy, termed Adaptive Robust Clipping (ARC).
+We show that ARC consistently enhances the empirical robustness of SOTA Robust-DGD methods, while preserving the theoretical robustness guarantees.
+Our analysis shows that ARC provably improves the asymptotic convergence guarantee of Robust-DGD in the case when the model is well-initialized.
+We validate this theoretical insight through an exhaustive set of experiments on benchmark image classification tasks.
+We observe that the improvement induced by ARC is more pronounced in highly heterogeneous and adversarial settings.
 1 Introduction
 Distributed machine learning, a.k.a.
-federated learning, has emerged as a dominant paradigm to cope with the increasing computational cost of learning tasks, mainly due to growing model sizes and datasets [25].
+federated learning, has emerged as a dominant paradigm to cope with the increasing computational cost of learning tasks, mainly due to growing model sizes and datasets (Kairouz et al., 2021).
 Worker machines, holding each a fraction of the training dataset, collaborate over a network to learn an optimal common model over the collection of their datasets.
-Workers typically collaborate with the help of a central coordinator, that we call server [36].
-Besides scalability, distributed learning is also helpful in preserving data privacy, since the workers do not have to share their local datasets during the learning.
-Conventional distributed learning algorithms are known to be vulnerable to misbehaving workers that could behave unpredictably [7, 25, 18].
+Workers typically collaborate with the help of a central coordinator, that we call server (McMahan et al., 2017).
+Besides scalability, distributed learning is also helpful in preserving data ownership and sovereignty, since the workers do not have to share their local datasets during the learning.
+Conventional distributed learning algorithms are known to be vulnerable to misbehaving workers that could behave unpredictably (Blanchard et al., 2017; Kairouz et al., 2021; Guerraoui et al., 2023).
 Misbehavior may result from software and hardware bugs, data poisoning, or malicious players controlling part of the network.
-In the parlance of distributed computing, misbehaving workers are referred to as Byzantine [32].
-Due to the growing influence of distributed learning in critical public-domain applications, such as healthcare [38] and finance [34], the problem of robustness to misbehaving workers, a.k.a.
-robust distributed learning, has received significant attention in the past [7, 44, 12, 14, 27, 17, 3, 13].
-Robust distributed learning algorithms primarily rely on robust aggregation, such as coordinate-wise trimmed mean (CWTM) [44], geometric median (GM) [10] and multi-Krum (MK) [7].
-Specifically, in a robust distributed gradient descent (Robust-DGD) method, the server aggregates the workers’ local gradients using a robust aggregation rule, instead of simply averaging them.
+In the parlance of distributed computing, misbehaving workers are referred to as Byzantine (Lamport et al., 1982).
+Due to the growing influence of distributed learning in critical public-domain applications such as healthcare (Nguyen et al., 2022) and finance (Long et al., 2020), the problem of robustness to misbehaving workers, a.k.a.
+robust distributed learning, has received significant attention (Blanchard et al., 2017; Yin et al., 2018; Farhadkhani et al., 2022; Karimireddy et al., 2022; Gorbunov et al., 2023; Allouah et al., 2023a; Farhadkhani et al., 2023).
+Robust distributed learning algorithms primarily rely on robust aggregation, such as coordinate-wise trimmed mean (CWTM) (Yin et al., 2018), geometric median (GM) (Chen et al., 2017) and multi-Krum (MK) (Blanchard et al., 2017).
+Specifically, in robust distributed gradient descent (Robust-DGD), the server aggregates the workers’ local gradients using a robust aggregation method, instead of simply averaging them.
 This protects the learning procedure from erroneous gradients sent by misbehaving workers.
-Recent work has made significant improvements over these aggregation rules: by incorporating a pre-aggregation step such as bucketing [27, 17] and nearest-neighbor mixing (NNM) [3] to tackle gradient dissimilarity that results from data heterogeneity.
-The learning guarantee of a resulting Robust-DGD method has been proven to be optimal [4], i.e., it cannot be improved unless additional assumptions are made, under the standard heterogeneity model of $(G,B)$-gradient dissimilarity [28].
-Despite the theoretical tightness of state-of-the-art (SOTA) Robust-DGD, we demonstrate that its performance can be improved in pragmatic cases wherein model initialization incurs bounded errors on workers’ local data.
-We do so by proposing clipping of workers’ gradients prior to their aggregation (at the server) using a novel clipping scheme, named adaptive robust clipping (ARC).
-We show that incorporating ARC provably improves the robustness of Robust-DGD when the local gradients at the initial model estimate are bounded.
-The factor of improvement is prominent when the tolerable fraction of misbehaving workers approaches the breakdown point, [^1] The breakdown point is the minimum fraction of misbehaving workers that can break the system, i.e., makes it impossible to guarantee any bound on the learning error [18].
-established in [4] under $(G,B)$-gradient dissimilarity.
-In general, when the local gradients can be arbitrarily large at initialization, ARC recovers the original learning guarantee of Robust-DGD.
-Main results & contributions.
-We consider a system comprising $n$ workers and a server.
-The goal is to tolerate up to $f$ misbehaving (a.k.a.
-adversarial) workers.
-The impact of data heterogeneity is characterized by the standard definition of $(G,B)$-gradient dissimilarity [28], recalled in Section 2.
-Figure 1: Worst-case maximal accuracies across several types of misbehaviour (see Section 6) of CWTM $\circ$ NNM and GM $\circ$ NNM, with ARC and without (i.e., no clipping), for a heterogeneously distributed MNIST dataset with $10$ honest workers.
-Heterogeneity is simulated using Dirichlet [23] distribution of parameter $\alpha=0.5$.
+Recent work has made significant improvements over these aggregation rules by incorporating a pre-aggregation step such as bucketing (Karimireddy et al., 2022; Gorbunov et al., 2023) and nearest-neighbor mixing (NNM) (Allouah et al., 2023a), to tackle gradient dissimilarity resulting from data heterogeneity.
+The learning guarantee of the resulting Robust-DGD has been proven to be optimal (Allouah et al., 2023b), i.e., it cannot be improved without additional assumptions under the standard heterogeneity model of $(G,B)$-gradient dissimilarity (Karimireddy et al., 2020).
+Despite its theoretical tightness, the empirical success of Robust-DGD has unknowingly relied on pre-aggregation gradient clipping (Mhamdi et al., 2021; Farhadkhani et al., 2022; Allouah et al., 2023a).
+Specifically, clipping the gradients of the workers prior to aggregation has been observed to sometimes enhance the algorithm’s empirical performance in the presence of adversarial workers, as evidenced in Figure 1a.
+Yet, this improvement lacks a concrete explanation, raising the question of whether the observed benefits of clipping are merely anecdotal.
+This leads to the natural inquiry: Why, and when, does pre-aggregation clipping improve robustness?
+(a) Sign-flipping (SF) (Allen-Zhu et al., 2020)
+(b) Label-flipping (LF) (Allen-Zhu et al., 2020)
+Figure 1: Impact of pre-aggregation clipping, specifically static clipping with $C=2$ and our adaptive clipping algorithm ARC, on Robust-DGD with aggregation scheme CWTM $\circ$ NNM, under the SF and LF adversarial attacks.
+We consider the MNIST (Deng, 2012) dataset distributed amongst $10$ honest (non-adversarial) workers with extreme heterogeneity, and there are three adversarial workers.
+In particular, using a constant clipping threshold, referred to as static clipping, has exhibited mixed results.
+Figure 1 shows that while static clipping effectively mitigates sign-flipping (SF) attacks, it completely fails under label-flipping (LF).
+This indicates an inherent fragility of static clipping.
+Indeed, we prove in our work that static clipping breaks the standard $(f,\kappa)$-robustness property of an aggregation method (Allouah et al., 2023a).
+This highlights a key shortcoming of existing empirical results that rely on static clipping (Mhamdi et al., 2021; Farhadkhani et al., 2022; Allouah et al., 2023a).
+To overcome the limitations of static clipping but preserve its empirical benefits at the same time, we introduce a novel adaptive clipping scheme, termed Adaptive Robust Clipping (ARC).
+ARC dynamically adjusts the clipping threshold as per the gradients sent by the workers and the fraction of adversarial workers to be tolerated.
+We demonstrate that integrating ARC into Robust-DGD consistently improves its empirical performance (see Figures 1 and 2a), while also preserving the convergence guarantee of the original Robust-DGD algorithm.
+Moreover, we show that when the model initialization is good, ARC provably improves the robustness of Robust-DGD.
+The benefits of ARC are more pronounced as the fraction of misbehaving workers approaches the system’s breakdown point [^1] Breakdown point refers the minimum fraction of adversarial workers that can break the system, making it impossible to guarantee a bound on the learning error Allouah et al.
+(2023b).
+and when the data across the workers is highly heterogeneous.
+Our key results are summarized below.
+Critical comparisons to prior work are deferred to Section 6.
+Main results & contributions. We consider a system comprising $n$ workers and a server. The goal is to tolerate up to $f$ adversarial workers.
 (1) Adaptive robust clipping (ARC).
 We propose ARC, wherein prior to aggregating the gradients, the server clips the largest $k\coloneqq\left\lfloor{2(f/n)(n-f)}\right\rfloor$ gradients using a clipping parameter given by the (Euclidean) norm of the $(k+1)$-th largest gradient.
-We prove 2 crucial properties: ARC (i) preserves the robustness guarantee of original aggregation and (ii) constrains an adversarial gradient’s norm by that of an honest (non-adversarial) worker’s gradient.
-(2) Improved learning error.
-Under $(G,B)$-gradient dissimilarity, [4] proves a lower bound on the learning error in the presence of adversarial workers.
-We show, however, that this lower bound can be circumvented under the assumption that the honest workers’ gradients are bounded at model initialization.
-Specifically, under this assumption, we prove that incorporating ARC in Robust-DGD improves the learning error beyond the lower bound when $f/n$ approaches the breakdown point.
-Thanks to its two aforementioned properties, ARC induces this improvement by effectively reducing the heterogeneity from $(G,B)$ to $(G+B\rho,0)$-gradient dissimilarity, where $\rho$ depends linearly on the norm of workers’ gradients at model initialization.
-We remark that the above assumption on model initialization is often satisfiable in practice [15].
-(3) Empirical validation.
-We conduct experiments on MNIST [11], Fashion-MNIST [42], and CIFAR-10 [31], under several data heterogeneity and adversarial regimes.
-Our results show that ARC boosts the performance of SOTA robust aggregation methods, especially when considering a large number of adversarial workers.
-A preview of our results on MNIST is shown in Figure 1.
-We make similar observations on the effectiveness of ARC in scenarios of high heterogeneity (see Figure 3).
-Related work.
-Gradient clipping is a well-known technique for tackling exploding gradients in deep learning [16].
-It has been extensively analyzed in the centralized setting [46, 45, 30], with applications to differential privacy [39].
-There is also some work analyzing the impact of clipping in non-robust distributed learning [47, 29].
-It has been shown that the clipping of local gradients causes distributed gradient descent to diverge.
-This motivates the need for an analysis of the clipping technique in our adversarial setting, and also hints at the inadequacy of clipping with a fixed threshold.
-In the context of robust distributed learning, prior work has proposed iterative clipping for robust aggregation [26].
-The analysis, however, cannot be applied to study the impact of pre-aggregation clipping that we consider.
-An adaptive gradient clipping scheme was also proposed for robust aggregation in [20, 33].
-However, the robustness guarantees are only provided for strongly convex loss functions and under a specific redundancy condition that is inadequate for our non-convex machine learning setting.
-Other work [37, 14, 3] has only used static pre-aggregation clipping to enhance the empirical robustness of the proposed algorithms, but fail to provide any theoretical guarantees.
-Recent work [35] has proposed clipping of the gradient differences, in conjunction with variance reduction, to tackle partial worker participation when adversarial workers can form a majority.
-Our focus is however different, as we consider the impact of pre-aggregation clipping on the robustness obtained by a general class of aggregation rules.
-Lastly, prior work [19, 48, 5] has also considered the challenge of privacy preservation alongside robustness in distributed learning.
-In this line of work, the gradients are assumed to have bounded norms in order to control the sensitivity of the algorithm, as is common in differential privacy (DP) [1].
-This is often ensured in practice by clipping the gradients using a static clipping parameter.
-Our findings suggest that it is better to use adaptive clipping instead of static clipping.
-Nevertheless, thoroughly analysing ARC in the context of DP remains an interesting research direction.
-Paper outline.
-Section 2 presents the problem statement.
-Section 3 presents the background on robust distributed learning, while Section 4 introduces ARC and its useful properties.
-Section 5 presents the convergence analysis of Robust-DGD with ARC, and characterizes the improvement in its robustness.
-Section 6 presents our empirical results.
-Lastly, Section 7 presents concluding remarks.
-Appendix A presents the computational complexity of ARC.
-Appendices B and C contain proofs for results presented in Sections 4 and 5, respectively.
-Appendix D describes the experimental setup, while Appendices E and F include additional experimental results.
-2 Problem Statement
-We consider the problem of distributed learning in the server-based architecture.
-The system comprises $n$ workers represented by $w_{1},\dots,w_{n}$, that collaborate with the help of a trusted server.
-The workers hold local datasets $\mathcal{D}_{1},\dots,\mathcal{D}_{n}$ respectively, each composed of $m$ data points from an input space $\mathcal{Z}$.
-Specifically, for any $i\in[n]$, $\mathcal{D}_{i}\coloneqq\{z_{1}^{(i)},\dots,z_{m}^{(i)}\}\subset\mathcal{Z}^{m}$.
-For a given model parameterized by vector $\theta\in\mathbb{R}^{d}$, $d$ being the number of trainable parameters in the model, each worker $w_{i}$ incurs a loss given by the loss function $\mathcal{L}_{i}{(\theta)}\coloneqq\frac{1}{m}\sum_{k=1}^{m}\ell(\theta,z_{k}^{(i)})\enspace$, where $\ell:\mathbb{R}^{d}\times\mathcal{Z}\rightarrow\mathbb{R}$ is the point-wise loss.
-We make the following standard assumptions [8]: (i) the point-wise loss function $\ell$ is differentiable with respect to $\theta$.
-(ii) For all $i\in[n]$, $\mathcal{L}_{i}$ is $L$-Lipschitz smooth, i.e., there exists $L\in\mathbb{R}^{+}$ such that all $\theta,\theta^{\prime}\in\mathbb{R}^{d}$, $\left\lVert{\nabla\mathcal{L}_{i}(\theta{})-\nabla\mathcal{L}_{i}(\theta{}^{\prime})}\right\rVert\leq L\left\lVert{\theta{}-\theta{}^{\prime}}\right\rVert$.
-In the ideal setting where all workers are assumed to be honest, i.e., follow the prescribed algorithm correctly, the server aims to minimize the global average loss function given by $\mathcal{L}(\theta{})\coloneqq\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}_{i}(\theta{})$.
-However, this learning objective is rendered vacuous when some workers could be adversarial, described in the following.
-Robust distributed learning.
-A robust distributed learning algorithm aims to output a good model despite the presence of some adversarial (a.k.a., Byzantine [32]) workers in the system [40, 44, 3].
+It is important to note that in contrast to existing adaptive clipping schemes (Diakonikolas et al., 2020; Abdalla & Zhivotovskiy, 2024), ARC does not require additional a priori information on honest workers’ gradients.
+We prove that ARC preserves the robustness guarantee of the original robust aggregation method.
+(2) Improved empirical robustness.
+We conduct experiments on MNIST (Deng, 2012), Fashion-MNIST (Xiao et al., 2017), and CIFAR-10 (Krizhevsky et al., 2014), across various data heterogeneity settings and adversarial regimes.
+Our results demonstrate that ARC significantly enhances the performance of state-of-the-art Robust-DGD methods, particularly in scenarios with high data heterogeneity (Figure 2a) and a large number of adversarial workers (Figure 4b).
+(3) Improved learning guarantee.
+We demonstrate that ARC possesses an additional property, we call Bounded Aggregation Output, that is not satisfied by classical robust aggregation methods.
+Specifically, ARC constrains the norm of an adversarial gradient by that of an honest (non-adversarial) gradient.
+Leveraging this property, we show that the lower bound established under data heterogeneity in Allouah et al.
+(2023b) can be circumvented using ARC, provided the honest workers’ gradients are bounded at model initialization.
+An empirical validation of this insight is shown in Figure 2b.
+Such model initialization is often satisfiable in practice (Glorot & Bengio, 2010), highlighting the practical relevance of ARC.
+When the model is arbitrarily initialized, ARC recovers the original convergence guarantee of Robust-DGD in the worst case (see Figure 2b).
+(a) Improved robustness
+(b) Influence of model initialization
+Figure 2: Worst-case maximal accuracies of Robust-DSGD, with and without ARC, across several types of misbehavior for distributed MNIST with $10$ honest workers and under extreme heterogeneity.
+On the left, we vary the number of adversarial workers.
+On the right, we vary the initialization conditions by scaling a well-chosen set of initial parameters (CWTM $\circ$ NNM is used, and $f=1$).
+More details on the experimental setup can be found in Sections 4 and 5.2, and Appendix D.
+2 Problem Statement and Relevant Background
+[…]
+A robust distributed learning algorithm aims to output a good model despite the presence of adversarial (a.k.a., Byzantine (Lamport et al., 1982)) workers in the system (Su & Vaidya, 2016; Yin et al., 2018; Allouah et al., 2023a).
 Specifically, the goal is to design a distributed learning algorithm that can tolerate up to $f$ adversarial workers, of a priori unknown identity, out of $n$ workers.
 Adversarial workers need not follow a prescribed algorithm, and can send arbitrary information to the server.
-In the context of distributed gradient descent, adversarial workers can send incorrect gradients to the server [6, 2, 43, 27].
-We let $\mathcal{H}\subseteq[n]$, with $\left\lvert{\mathcal{H}}\right\rvert=n-f$, denote the set of honest workers that do not deviate from an algorithm.
-The objective of the server is to minimize the honest workers’ average loss function given by
+In the context of distributed gradient descent, adversarial workers can send incorrect gradients to the server (Baruch et al., 2019; Allen-Zhu et al., 2020; Xie et al., 2020; Karimireddy et al., 2022).
+We let $\mathcal{H}\subseteq[n]$, with $\left\lvert{\mathcal{H}}\right\rvert=n-f$, denote the set of honest workers that do not deviate from the algorithm.
+The objective of the server is to minimize the average loss function of honest workers given by
 $\mathcal{L}_{\mathcal{H}}(\theta)\coloneqq\frac{1}{\left\lvert{\mathcal{H}}\right\rvert}\sum_{i\in\mathcal{H}}\mathcal{L}_{i}(\theta),~{}\forall\theta\in\mathbb{R}^{d}.$
-While we can find a minimum of $\mathcal{L}_{\mathcal{H}}(\theta)$ when $\mathcal{L}_{\mathcal{H}}$ is convex, in general however the loss function is non-convex, and the optimization problem is NP-hard [9].
+While we can find a minimum of $\mathcal{L}_{\mathcal{H}}(\theta)$ when $\mathcal{L}_{\mathcal{H}}$ is convex, in general however the loss function is non-convex, and the optimization problem is NP-hard (Boyd & Vandenberghe, 2004).
 Therefore, we aim to find a stationary point of $\mathcal{L}_{\mathcal{H}}$ instead, i.e., $\theta^{*}$ such that $\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}(\theta^{\ast})}\right\rVert=0$.
-Formally, we define robustness to adversarial workers by $(f,\epsilon)$-resilience [3].
+Formally, we define robustness to adversarial workers by $(f,\epsilon)$-resilience (Allouah et al., 2023a).
 Definition 2.1.
@@
 If a distributed learning algorithm $\mathcal{A}$ is $(f,\varepsilon)$-resilient then it can tolerate up to $f$ adversarial workers.
-Thus, $f/n$ is referred to as the maximum tolerable fraction of adversarial workers for $\mathcal{A}$.
-It is standard in robust distributed learning to design an algorithm using $f$ as a parameter [40, 7, 44, 27, 17, 3].
-Data heterogeneity. In the context of distributed learning, data heterogeneity is characterized by the following standard notion of $(G,B)$-gradient dissimilarity [41, 28, 27, 17, 4].
+It is standard in robust distributed learning to design robust algorithms using $f$ as a parameter (Su & Vaidya, 2016; Blanchard et al., 2017; Yin et al., 2018; Karimireddy et al., 2022; Gorbunov et al., 2023; Allouah et al., 2023a), $f$ being the maximum tolerable number of adversarial workers.
+In the context of distributed learning, data heterogeneity is characterized by the following standard notion of $(G,B)$-gradient dissimilarity (Vaswani et al., 2019; Karimireddy et al., 2020; 2022; Gorbunov et al., 2023; Allouah et al., 2023b).
 Definition 2.2.
@@
 $\frac{1}{\left\lvert{\mathcal{H}}\right\rvert}\sum_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}(\theta)-\nabla\mathcal{L}_{\mathcal{H}}(\theta)}\right\rVert^{2}\leq G^{2}+B^{2}\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}(\theta)}\right\rVert^{2},\quad\forall\theta\in\mathbb{R}^{d}.$
-3 Relevant Background
-We present in this section useful precursors on robust distributed learning.
 General limitations on robustness.
-Note that $(f,\varepsilon)$-resilience is impossible (for any $\varepsilon$) when $f/n\geq 1/2$ [20].
-Moreover, even for $f/n<1/2$, it is generally impossible to achieve $(f,\varepsilon)$-resilience for arbitrarily small $\varepsilon$ due to the disparity amongst the local datasets $\mathcal{D}_{i},\;i\in\mathcal{H}$ (a.k.a., data heterogeneity) [33].
-Henceforth, we assume that $n>2f$ by default.
+Note that $(f,\varepsilon)$-resilience is impossible (for any $\varepsilon$) when $f/n\geq 1/2$ (Gupta & Vaidya, 2020).
+Moreover, even for $f/n<1/2$, it is generally impossible to achieve $(f,\varepsilon)$-resilience for arbitrarily small $\varepsilon$ due to the disparity amongst the local datasets $\mathcal{D}_{i},\;i\in\mathcal{H}$ (a.k.a., data heterogeneity) (Liu et al., 2021).
+Henceforth, we assume that $n>2f$.
 Under $(G,B)$-gradient dissimilarity, we have the following lower bound on the achievable resilience.
-Lemma 3.1 (Non-convex extension of Theorem 1 in [4]).
-Under $(G,B)$-gradient dissimilarity, a distributed learning algorithm $\mathcal{A}$ satisfies $(f,\varepsilon)$-resilience only if $\frac{f}{n}<\frac{1}{2+B^{2}}$ and $\varepsilon\geq\frac{1}{4}\cdot\frac{f}{n-(2+B^{2})f}G^{2}$.
-For a given distributed learning problem, the minimum fraction of adversarial workers that renders any distributed learning algorithm ineffective is referred the breakdown point [18, 4].
+Lemma 2.3 (Non-convex extension of Theorem 1 Allouah et al.).
+Under $(G,B)$-gradient dissimilarity, a distributed learning algorithm is $(f,\varepsilon)$-resilient only if $\frac{f}{n}<\frac{1}{2+B^{2}}$ and $\varepsilon\geq\frac{1}{4}\cdot\frac{f}{n-(2+B^{2})f}G^{2}$.
+For a given distributed learning problem, the minimum fraction of adversarial workers that renders any distributed learning algorithm ineffective is called the breakdown point (Guerraoui et al., 2023; Allouah et al., 2023b).
 Thus, according to Lemma 2.3, there exists a distributed learning problem satisfying $(G,B)$-gradient dissimilarity whose breakdown point is given by $1/(2+B^{2})$.
-In the ideal setting, i.e., when all the workers are honest, the server can execute the classic distributed gradient descent (DGD) method in collaboration with the workers to find a stationary point of the global average loss function.
-In DGD, the server maintains the model, which is updated periodically using the average of the gradients computed by the workers on their local datasets.
-The averaging of the gradients is rendered ineffective when some workers are adversarial, i.e., they can send malicious vectors posing as their gradients to the server [18].
-To remedy this, averaging is replaced with robust aggregation, e.g., coordinate-wise trimmed mean (CWTM) and median (CWMed) [44], geometric median (GM) [10], and multi-Krum (MK) [7].
-Useful properties of a resulting Robust-DGD method, which is summarized in Algorithm 1, are recalled below.
-Robust distributed gradient descent (Robust-DGD).
-In Robust-DGD, robust aggregation protects the learning procedure from incorrect gradients sent by the adversarial workers.
-The robustness of an aggregation can be quantified by the following property of $(f,\kappa)$-robustness, proposed in [3].
+Robust distributed gradient descent.
+To tolerate adversarial workers, we replace the averaging operation in the classic distributed gradient descent (DGD) method with a robust aggregation, e.g., coordinate-wise trimmed mean (CWTM) and median (CWMed) (Yin et al., 2018), geometric median (GM) (Chen et al., 2017), and multi-Krum (MK) (Blanchard et al., 2017).
+This yields Robust-DGD, presented in Algorithm 1, where robust aggregation protects the learning from incorrect gradients sent by the adversarial workers.
+The robustness of an aggregation method can be quantified by the following property of $(f,\kappa)$-robustness, proposed by Allouah et al.
+(2023a).
 Algorithm 1 Robust Distributed Gradient Descent (Robust-DGD)
-  Initialization: Server chooses a model ${{\theta}_{1}}\in\mathbb{R}^{d}$, a learning rate $\gamma\in\mathbb{R}^{+}$ and a robust aggregation rule $\mathbf{F}:\mathbb{R}^{n\times d}\to\mathbb{R}^{d}$.
+  Initialization: Server chooses a model ${{\theta}_{1}}\in\mathbb{R}^{d}$, a learning rate $\gamma\in\mathbb{R}^{+}$ and a robust aggregation method $\mathbf{F}:\mathbb{R}^{n\times d}\to\mathbb{R}^{d}$.
   for $t=1$ to $T$ do
@@
 Definition 2.4.
-Aggregation rule ${\mathbf{F}:\mathbb{R}^{n\times d}\to\mathbb{R}^{d}}$ is said to be $(f,\kappa)$-robust if there exists $\kappa\in\mathbb{R}$ such that, for all $x_{1},\ldots,x_{n}\in\mathbb{R}^{d}$ and any set $S\subseteq[n]$, $\left\lvert{S}\right\rvert=n-f$, the following holds true:
+${\mathbf{F}:\mathbb{R}^{n\times d}\to\mathbb{R}^{d}}$ is said to be $(f,\kappa)$-robust if there exists a robustness coefficient $\kappa\in\mathbb{R}$ such that for all $x_{1},\ldots,x_{n}\in\mathbb{R}^{d}$ and any set $S\subseteq[n]$, $\left\lvert{S}\right\rvert=n-f$, the following holds:
 $\left\lVert{\mathbf{F}(x_{1},\ldots,x_{n})-\overline{x}_{S}}\right\rVert^{2}\leq\frac{\kappa}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}\enspace,\quad\text{where }~{}\overline{x}_{S}=\frac{1}{\left\lvert{S}\right\rvert}\sum\limits_{i\in S}x_{i}.$
-We refer to $\kappa$ as the robustness coefficient.
-The property of $(f,\kappa)$-robustness encompasses several robust aggregations, including the ones mentioned above and more (e.g., see [3]).222$(f,\kappa)$-robustness unifies other robustness definitions: $(\delta_{\text{max}},c)$-robustness [26] and $(f,\lambda)$-resilience [14].
-It has been shown that an aggregation rule is $(f,\kappa)$-robust only if $\kappa\geq\frac{f}{n-2f}$ [3].
-Importantly, the asymptotic error of Robust-DGD with $(f,\kappa)$-robust aggregation, where $\kappa\in\mathcal{O}(f/(n-2f))$, matches the lower bound (recalled in Lemma 3.1[4].
-Note that all the aforementioned aggregation rules, i.e., CWTM, CWMed, GM and MK, attain optimal robustness, i.e., are $(f,\mathcal{O}\left(f/(n-2f)\right))$-robust, when composed with the pre-aggregation step: nearest neighbor mixing (NNM) [3].
-[^3] CWTM has been shown to be optimal even without NNM [3].
+$(f,\kappa)$-robustness [^2] Unifies $(\delta_{\text{max}},c)$-robustness (Karimireddy et al., 2021) and $(f,\lambda)$-resilience (Farhadkhani et al., 2022).
+encompasses several robust aggregations, including the ones mentioned above and more (e.g., see Allouah et al.
+(2023a)).
+It has been shown that an aggregation method is $(f,\kappa)$-robust only if $\kappa\geq\frac{f}{n-2f}$ (Allouah et al., 2023a).
+Importantly, the asymptotic error of Robust-DGD with $(f,\kappa)$-robust aggregation, where $\kappa\in\mathcal{O}(f/(n-2f))$, matches the lower bound (recalled in Lemma 2.3(Allouah et al., 2023b).
+Note that the aforementioned aggregation methods (CWTM, CWMed, GM and MK) attain optimal robustness when composed with the pre-aggregation scheme: nearest neighbor mixing (NNM) (Allouah et al., 2023a).
+[^3] CWTM has been shown to be optimal even without NNM.
 Specifically, we recall the following result.
-Lemma 3.3 (Lemma 1 in [3]).
+Lemma 2.5 (Lemma 1 in Allouah et al. (2023a)).
 For $\mathbf{F}\in\left\{\mathbf{CWTM},\mathbf{CWMed},\mathbf{GM},\mathbf{MK}\right\}$, the composition $\mathbf{F}\circ\mathbf{NNM}$ is $(f,\kappa)$-robust with
@@
 We let $\mathcal{L}_{\mathcal{H}}^{*}$ denote the minimum value of $\mathcal{L}_{\mathcal{H}}({{\theta}})$.
-Lemma 3.4 (Theorem 2 in [4]).
+Lemma 2.6 (Theorem 2 in Allouah et al. (2023b)).
 Consider Algorithm 1 with $\gamma\leq 1/L$.
 Let $\Delta_{o}\in\mathbb{R}^{+}$ such that $\mathcal{L}_{\mathcal{H}}({{\theta}_{1}})-\mathcal{L}_{\mathcal{H}}^{*}\leq\Delta_{o}$.
-If the aggregation rule $\mathbf{F}$ is $(f,\kappa)$-robust with $\kappa B^{2}<1$, then
+If $\mathbf{F}$ is $(f,\kappa)$-robust with $\kappa B^{2}<1$, then
 $\frac{1}{T}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}({{\theta}_{t}})}\right\rVert^{2}\leq\frac{2\Delta_{o}}{(1-\kappa B^{2})\gamma T}+\frac{\kappa G^{2}}{1-\kappa B^{2}}.$
@@
 3 Adaptive Robust Clipping (ARC) and its Properties
-In this section, upon presenting a preliminary observation on the fragility of static clipping, we introduce adaptive robust clipping (i.e., ARC). Lastly, we present some useful properties of ARC.
+In this section, we first present a preliminary observation on the fragility of static clipping, and then introduce adaptive robust clipping (i.e., ARC) along with its robustness guarantees.
 For a clipping parameter $C\in\mathbb{R}^{+}$ and a vector $x\in\mathbb{R}^{d}$, we denote $\mathrm{clip}_{C}(x)\coloneqq\min\left(1,\frac{C}{\left\lVert{x}\right\rVert}\right)x$.
@@
 For any fixed $C\in\mathbb{R}^{+}$ and $\kappa^{\prime}\geq 0$, ${\mathbf{F}\circ\mathbf{Clip}_{C}}$ is not $(f,\kappa^{\prime})$-robust.
-This means that if the clipping threshold is a priori fixed, i.e., independent from the input vectors, pre-aggregation clipping does not preserve the robustness of the original aggregation.
-This fragility of such static clipping is also apparent in practice, as shown by our experimental study in Appendix F.
+Thus, if the clipping threshold is fixed, i.e., independent from the input vectors, pre-aggregation clipping does not preserve the robustness of the original aggregation.
+This fragility of such static clipping is also apparent in practice, as shown by our experimental study in Appendix F (and Figure 1).
 Description of ARC.
-In ARC, the clipping threshold is determined according to the input vectors.
-More precisely, ARC clips the largest $k=\left\lfloor{2(f/n)(n-f)}\right\rfloor$ vectors using a clipping parameter given by the norm of the $(k+1)$-th largest input vector.
+ARC is an adaptive clipping scheme that only makes use of the standard robustness parameter $\nicefrac{{f}}{{n}}$, i.e., the tolerable fraction of adversarial workers.
+ARC is adaptive in the sense that the clipping threshold $C$ is not fixed but depends on the input vectors.
+Specifically, ARC clips the largest $k=\left\lfloor{2(f/n)(n-f)}\right\rfloor$ vectors using a clipping parameter given by the norm of the $(k+1)$-th largest input vector.
 The overall scheme is formally presented in Algorithm 2, and its computational complexity is $\mathcal{O}(nd+n\log(n))$ (see Appendix A).
-ARC is adaptive in the sense that the clipping threshold $C$ is not fixed but depends on the input vectors.
+Thus, pre-composing more computationally expensive schemes like NNM and multi-Krum, which have a complexity of $\mathcal{O}(dn^{2})$, with ARC does not introduce a significant overhead.
 Algorithm 2 Adaptive Robust Clipping (ARC)
@@
   Output: $\mathbf{Clip}_{C}(x_{1},\ldots,x_{n})$
-Properties of ARC.
-We present below important properties of ARC.
+Robustness Guarantee.
+We present below the preservation of robustness guaranteed by ARC.
 Let $\mathbf{F}$ be an $(f,\kappa)$-robust aggregation rule and $\mathbf{F}\circ\mathbf{ARC}(x_{1},\dots,x_{n})\coloneqq\mathbf{F}(\mathbf{ARC}(x_{1},\dots,x_{n}))$
@@
 If $\mathbf{F}$ is $(f,\kappa)$-robust, then $\mathbf{F}\circ\mathbf{ARC}$ is $\left(f,\kappa+\frac{2f}{n-2f}\right)$-robust.
+Proofs for the results presented in this section are deferred to Appendix B.
 Since $\kappa\geq\frac{f}{n-2f}$ (recalled in Section 2), Theorem 3.2 implies that $\mathbf{F}\circ\mathbf{ARC}$ is $\left(f,3\kappa\right)$-robust.
 In other words, ARC preserves the robustness of the original aggregation scheme.
-Besides preserving robustness in the context of Robust-DGD, ARC also limits the norms of adversarial workers’ gradients by the norm of an honest worker’s gradient.
-Specifically, when $f/n<1/2$, the clipping threshold $C$ in ARC is bounded by $\max_{i\in\mathcal{H}}\left\lVert{x_{i}}\right\rVert$ (see Appendix C.2).
-This property is useful in proving our convergence result, stated below in Theorem 5.1.
-Proofs for Lemma 4.1 and Theorem 4.2 are deferred to Appendix B.
-5 Learning Guarantee of Robust-DGD with ARC
-In this section, we first present our convergence result for Robust-DGD with ARC (or Robust-DGD $\circ$ ARC), and then characterize its improvement over the lower bound recalled in Lemma 3.1.
-Specifically, throughout this section, we consider Algorithm 1 with aggregation $\mathbf{F}\circ\mathbf{ARC}$, i.e., in each learning step $t$, $R_{t}\coloneqq\mathbf{F}\circ\mathbf{ARC}\left({{g}^{\left({1}\right)}_{t}},\dots,\,{{g}^{\left({n}\right)}_{t}}\right)$ Note that, since we have already shown in Theorem 4.2 that $\mathbf{F}\circ\mathbf{ARC}$ is $(f,3\kappa)$-robust, a general convergence result for Robust-DGD $\circ$ ARC follows verbatim from Lemma 3.4, replacing $\kappa$ with $3\kappa$.
-However, we obtain below an alternative convergence result that enables us to show the improvement in learning error due to ARC.
-5.1 Convergence of Robust-DGD with ARC
-We assume that, for any set of vectors $x_{1},\ldots,x_{n}\in\mathbb{R}^{d}$, $\left\lVert{\mathbf{F}\left(x_{1},\ldots,x_{n}\right)}\right\rVert\leq\max_{i\in[n]}\left\lVert{x_{i}}\right\rVert$.
-This assumption can be made without loss of generality, the reason for which is deferred to Appendix C.1.
-Theorem 5.1.
-Let $\mathbf{F}$ be $(f,\kappa)$-robust.
-Let $\Delta_{o}$ and $\rho$ be real values such that $\mathcal{L}_{\mathcal{H}}({{\theta}_{1}})-\mathcal{L}_{\mathcal{H}}^{*}\leq\Delta_{o}$ and $\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{1}})}\right\rVert\leq\exp\left(-\frac{\Delta_{o}}{\kappa G^{2}}L\right)\rho$.
-If $\gamma=\min\left\{\left(\frac{\Delta_{o}}{\kappa G^{2}}\right)\frac{1}{T},~{}\frac{1}{L}\right\}$, then
-$\frac{1}{T}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}\leq\max\left\{5\kappa\left(G^{2}+B^{2}\rho^{2}\right),~{}\frac{2\Delta_{o}L}{T}+3\kappa\left(G^{2}+B^{2}\rho^{2}\right)\right\}.$
-We have the following corollary of Theorem 5.1 and Lemma 3.4.
-Corollary 5.2.
-For parameters given in Theorem 5.1, if $T\geq\frac{\Delta_{o}L}{\kappa G^{2}}$, then Robust-DGD $\circ$ ARC achieves
-$\mathop{{}\mathbb{E}}\left[{\left\lVert{\nabla\mathcal{L}\left(\hat{{{\theta}}}\right)}\right\rVert^{2}}\right]\leq 5\kappa~{}\min\left\{G^{2}+B^{2}\rho^{2},~{}\frac{G^{2}}{\max\{(1-\kappa B^{2}),0\}}\right\},$
-where $\mathop{{}\mathbb{E}}\left[{\cdot}\right]$ denotes the expectation over the choice of $\hat{{{\theta}}}$.
-5.2 Improvement in Robustness
-In the following, we show that if the local gradients at the initial model ${{\theta}_{1}}$ are bounded, then Robust-DGD $\circ$ ARC overcomes the lower bound recalled in 3.1.
-Specifically, we make this assumption:
-Assumption 5.3.
-There exists $\zeta\in\mathbb{R}^{+}$ such that, at the initial model ${{\theta}_{1}}$, $\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}({{\theta}_{1}})}\right\rVert\leq\zeta$.
-For ease of presentation, we denote $\mathsf{BP}\coloneqq\frac{1}{2+B^{2}},$
-$\quad\varepsilon_{o}\left(f/n\right)\coloneqq\frac{1}{4}\cdot\frac{G^{2}(f/n)}{1-(2+B^{2})(f/n)},\quad\text{ and }$ $\Psi(G,B,\rho)\coloneqq 640\left(1+\frac{1}{B^{2}}\right)^{2}\left(1+\frac{B^{2}\rho^{2}}{G^{2}}\right).$
-Recall from Lemma 3.1 that $\mathsf{BP}$ and $\varepsilon_{o}(f/n)$ are the breakdown point and the achievable stationarity error when $f/n<\mathsf{BP}$, respectively, for the worst-case scenario under $(G,B)$-gradient dissimilarity.
-We obtain the following theorem for Robust-DGD $\circ$ ARC.
-Theorem 5.4.
-Suppose Assumption 5.3 and $B>0$.
-Let $\mathbf{F}\in\left\{\mathbf{CWTM},\mathbf{CWMed},\mathbf{GM},\mathbf{MK}\right\}$, and let the remaining parameters for Robust-DGD $\circ$ ARC be as defined in Corollary 5.2.
-Let $\xi_{o}<1$ be a (strictly) positive real value, and let $\rho\coloneqq\exp\left(\frac{(2+B^{2})\Delta_{o}}{(1-\xi_{o})G^{2}}L\right)\zeta$.
-Then, the following holds true:
-  1. 1.
-    If $\frac{f}{n}\coloneqq(1-\xi)\mathsf{BP}$, where $\xi\in(0,\xi_{o}]$, then, $\mathop{{}\mathbb{E}}\left[{\left\lVert{\nabla\mathcal{L}\left(\hat{{{\theta}}}\right)}\right\rVert^{2}}\right]\leq\xi\,\Psi(G,B,\rho)\cdot\varepsilon_{o}\left(f/n\right)$
-  2. 2.
-    If $\mathsf{BP}\leq f/n<1/2$, then $\mathop{{}\mathbb{E}}\left[{\left\lVert{\nabla\mathcal{L}\left(\hat{{{\theta}}}\right)}\right\rVert^{2}}\right]\leq\Psi(G,B,\rho)\cdot\frac{G^{2}}{8}$
-According to Theorem 5.4, if the local gradients at model initialization are bounded, then Robust-DGD $\circ$ ARC can tolerate a larger fraction of adversarial workers than the breakdown point $\mathsf{BP}$ under $(G,B)$-gradient dissimilarity.
-Moreover, even if $f/n<\mathsf{BP}$, Robust-DGD $\circ$ ARC can improve over the worst-case stationarity error under $(G,B)$-gradient dissimilarity, formally shown in the following corollary of Theorem 5.4.
-For the corollary, we consider the same parameters as in Theorem 5.4.
-Corollary 5.5.
-Suppose Assumption 5.3 and that $f/n\coloneqq(1-\xi)~{}\mathsf{BP}$, where $\xi\in(0,\xi_{o}]$.
-For $\upsilon\in(0,1)$, if $\xi\leq\frac{1}{\Psi(G,B,\rho)}\,\upsilon$, then $\mathop{{}\mathbb{E}}\left[{\left\lVert{\nabla\mathcal{L}\left(\hat{{{\theta}}}\right)}\right\rVert^{2}}\right]\leq\upsilon\cdot\varepsilon_{o}\left(f/n\right)$.
-Corollary 5.5 implies that when local gradients at model initialization are bounded, Robust-DGD $\circ$ ARC improves over the lower bound $\varepsilon_{o}\left(f/n\right)$ for sufficiently small $\xi$ (or $f/n$ sufficiently close to $\mathsf{BP}$).
-[^5] An improvement over the lower bound automatically implies improvement over the original learning guarantee of Robust-DGD (recalled in Lemma 3.4), i.e., when model initialization is arbitrary.
-The closer $f/n$ is to $\mathsf{BP}$, the smaller is the factor of reduction $\upsilon$, and thus higher is the improvement.
-The threshold on $\xi$ approaches 00 when $B\rightarrow 0$, i.e., the space of improvement is quite narrow.
-This is expected, since the lower bound approaches $G^{2}f/(4(n-2f))$ in this case, which is already smaller than $5\kappa G^{2}$ (the error for Robust-DGD $\circ$ ARC in this case) for any $(f,\kappa)$-robust $\mathbf{F}$.
-Reconciliation with the lower bound.
-When the workers’ gradients at model initialization can be arbitrarily large, the convergence guarantee of Robust-DGD $\circ$ ARC reduces to that of Robust-DGD.
-On the contrary, if the model initialization is good, i.e., $\zeta$ is small, then ARC improves the learning error, as soon as the tolerable fraction of adversarial workers is in the proximity of $\mathsf{BP}$.
-We remark that Assumption 5.3, pertaining to model initialization, is not too strong and can be satisfied in practice [15], also validated by our experimental study shown in Figure 17 of Appendix F.
-Effectively, under Assumption 5.3, ARC reduces the heterogeneity model to $(G+B\rho,0)$-gradient dissimilarity, with $\rho$ defined in Theorem 5.4, since $\max_{t\in[T]}\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}({{\theta}_{t}})}\right\rVert\leq\rho$ in Robust-DGD $\circ$ ARC for the parameters defined in Theorem 5.1, shown implicitly in the proof of Theorem 5.1 in Appendix C.
-Proofs for all results presented in this section are deferred to Appendix C.
+Therefore, a convergence result for Robust-DGD with ARC follows verbatim from Lemma 2.6, replacing $\kappa$ with $3\kappa$.
+Despite this multiplicative factor of 3, we observe in the next section that incorporating ARC consistently improves the empirical performance of classical aggregation methods.
+A more detailed theoretical explanation will be provided later in Section 5.1.
 4 Empirical Evaluation
-In this section, we delve into the practical performance of ARC when incorporated in Robust-DSGD (see Algorithm 3 in Appendix D), an order-optimal stochastic variant of Robust-DGD [3].
+In this section, we delve into the practical performance of ARC when incorporated in Robust-DSGD (Algorithm 3 in Appendix D), an order-optimal stochastic variant of Robust-DGD (Allouah et al., 2023a).
 We conduct experiments on standard image classification tasks, covering different adversarial scenarios.
-We empirically test four SOTA aggregation rules when pre-composed with ARC.
+We empirically test four aggregation methods when pre-composed with ARC.
 We also contrast these outcomes against the performance of Robust-DSGD when no gradient clipping is used.
-Our findings underscore that clipping workers’ gradients using ARC prior to the aggregation is crucial to ensure the robustness of existing algorithms, especially in extreme scenarios, i.e., when either the data heterogeneity is high or there is a large fraction of adversarial workers to be tolerated.
-We plot in Table 1, and Figures 1 and 3, the metric of worst-case maximal accuracy.
-For each of the five Byzantine attacks executed (see Appendix D.3), we record the maximal accuracy achieved by Robust-DSGD during the learning procedure under that attack.
-The worst-case maximal accuracy is thus the smallest maximal accuracy reached across the five attacks.
-As the adversarial attack cannot be known in advance, this metric is critical to accurately evaluate the robustness of aggregation methods, since it provides an estimate of the potential worst-case performance of the algorithm.
-Furthermore, we use the Dirichlet [23] distribution of parameter $\alpha$ to simulate data heterogeneity in the workers’ datasets.
-The comprehensive experimental setup can be found in Appendix D.
-In this section, we present results on MNIST [11] and CIFAR-10 [31], and defer our findings on Fashion-MNIST [42] to Appendix E.2.
-Finally, for the sake of clarity, we mainly consider here the aggregations CWTM and GM, and show the remaining results on CWMed and MK in Appendix E.
+Our findings underscore that clipping workers’ gradients using ARC prior to the aggregation is crucial to ensure the robustness of existing algorithms, especially in extreme scenarios, i.e., when either the data heterogeneity is high or the fraction of adversarial workers is large.
 (a) $\alpha=0.5$
@@
 (c) extreme heterogeneity
-Figure 2: Performance of Robust-DSGD when using ARC and without clipping on MNIST. Among $n=11$, we consider $f=1$ adversarial worker executing FOE [43], and vary the heterogeneity.
-ARC boosts robustness in high heterogeneity.
-Figure 2 shows the performance of Robust-DSGD against the FOE [43] attack when using ARC opposed to no clipping with $f=1$ adversarial worker among $n=11$.When the heterogeneity is low (i.e., $\alpha=0.5$, see Figure 5), the performance of ARC and no clipping are comparable for both aggregations.
-When the heterogeneity increases ($\alpha=0.1$), the benefits of using ARC are more visible.
-CWTM $\circ$ NNM and GM $\circ$ NNM significantly struggle to learn (with a very large variance), while the same aggregations composed with ARC almost match the accuracy of DSGD towards the end of the training.
-In extreme heterogeneity, the improvement induced by our method is the most pronounced, as ARC enables both aggregations to reach a final accuracy close to 90%.
-Contrastingly, the same aggregations without clipping stagnate at an accuracy close to 10% throughout the training.
-Interestingly, only $f=1$ adversarial worker among $n=11$ (i.e., less than 10%) suffices to completely deteriorate the learning when no clipping is applied.
+Figure 3: Performance of Robust-DSGD when using ARC and without clipping on MNIST. There are 10 honest workers and $f=1$ adversarial worker executing FOE (Xie et al., 2020).
+We show in Table 1, and Figures 2a and 4, the metric of worst-case maximal accuracy.
+For each of the five Byzantine attacks executed (see Appendix D.3), we record the maximal accuracy achieved by Robust-DSGD during the learning procedure under that attack.
+The worst-case maximal accuracy is thus the smallest maximal accuracy reached across the five attacks.
+As the attack cannot be known in advance, this metric is critical to accurately evaluate the robustness of aggregation methods, since it provides an estimate of the potential worst-case performance of the algorithm.
+Furthermore, we use the Dirichlet (Hsu et al., 2019a) distribution of parameter $\alpha$ to simulate data heterogeneity.
+The comprehensive experimental setup can be found in Appendix D.
+In this section, we focus on presenting our results for the CWTM and GM aggregation methods applied to MNIST and CIFAR-10.
+Results for Fashion-MNIST, as well as for the CWMed and MK aggregations, are provided in Appendix E.
+All aggregation methods in our experiments are pre-composed with NNM (Allouah et al., 2023a), but we omit explicit mention of NNM in their names throughout the text for simplicity.
 (a) $f=1$
-(b) extreme heterogeneity
+(b) $\alpha=0.5$
 Figure 4: Worst-case maximal accuracies achieved by Robust-DSGD, with and without ARC, on heterogeneously-distributed MNIST with $10$ honest workers.
 Left: $f=1$ adversarial worker among $n=11$ for varying levels of heterogeneity.
-Right: extreme heterogeneity for varying $f$.
+Right: $\alpha=0.5$ for varying $f$.
+ARC boosts robustness in high heterogeneity.
+Figure 3 shows the performance of Robust-DSGD against the FOE attack when using ARC opposed to no clipping.
+In low heterogeneity (i.e., $\alpha=0.5$, see Figure 7), the performances of ARC and no clipping are comparable for both aggregations.
+When the heterogeneity increases ($\alpha=0.1$), the benefits of using ARC are more visible.
+CWTM and GM significantly struggle to learn (with a very large variance), while the same aggregations composed with ARC almost match the accuracy of DSGD towards the end of the training.
+In extreme heterogeneity, the improvement induced by our method is the most pronounced, as ARC enables both aggregations to reach a final accuracy close to 90%.
+Contrastingly, the same aggregations without clipping stagnate at around 10% throughout the training.
+Interestingly, only $f=1$ adversarial worker among $n=11$ (i.e., less than 10%) suffices to completely deteriorate the learning when no clipping is applied, highlighting the necessity of ARC in extreme heterogeneity.
 Varying heterogeneity for fixed $f$.
 Figure 4a compares the worst-case maximal accuracies achieved by Robust-DSGD when $f=1$, with ARC and without, for varying levels of heterogeneity.
-While CWTM $\circ$ NNM and GM $\circ$ NNM yield comparable accuracies with ARC in low heterogeneity ($\alpha\geq 0.5$), their performance significantly degrades when $\alpha$ drops below that threshold.
+While CWTM and GM yield comparable accuracies with ARC in low heterogeneity ($\alpha\geq 0.5$), their performance significantly degrades when $\alpha$ drops below that threshold.
 Indeed, when $\alpha=0.1$, their accuracies drop below 65%, while ARC enables the same aggregations to maintain their accuracy at just below 98%.
@@
 Similar plots for $f=3,5,7$ convey similar observations (see Appendix E.1).
-Essentially, in settings where gradient clipping is not needed (e.g., low heterogeneity or small $f$), ARC performs at least as well as no clipping.
-Nonetheless, ARC induces a significant improvement near the breakdown point of SOTA robust aggregation schemes (e.g., high heterogeneity or large $f$).
+Essentially, in low heterogeneity or when the fraction of adversarial workers is small, no clipping yields good empirical results and ARC performs at least as well as no clipping.
+However, ARC induces a significant improvement near the breakdown point of SOTA robust aggregation schemes (e.g., high heterogeneity or large $f$).
+ARC increases the breakdown point in adverse settings.
+Figure 2a of Section 1 shows that for $f\in\{1,...,5\}$, Robust-DSGD with CWTM and GM completely fails to learn, consistently yielding worst-case maximal accuracies close to 15%.
+This suggests that $f=1$ constitutes the breakdown point for these aggregations in extreme heterogeneity.
+However, composing them with ARC increases the breakdown point of these aggregations to $f=3$.
+Indeed, for $f=1$ and 2, ARC enables Robust-DSGD to achieve accuracies greater than 85% in the worst case.
+However, when $f\geq 3$, the performance degrades, although CWTM with ARC is still able to reach a satisfactory accuracy close to 70%.
+Moreover, even when the heterogeneity is not large, ARC still produces a significant improvement when the fraction of adversarial workers increases in the system.
+Indeed, in Figure 4b , the performances of ARC and no clipping are comparable for $f\leq 3$.
+However, the improvement induced by ARC is much more visible when $f\geq 5$.
+Particularly when $f=7$, ARC enables CWTM and GM to reach accuracies greater than 80% in the worst case, whereas the same aggregations yield accuracies below 40% without clipping .
+This observation indicates the raise of the breakdown point due to ARC in practice.
+Plots for $\alpha=0.1$ and 1 convey the same observations in Appendix E.1.
+Figure 5: FOE on heterogeneous CIFAR-10 with $\alpha=0.075$. $f=1$ among $n=17$.
+Improved robustness on CIFAR-10.
+We also conduct experiments on the more challenging CIFAR-10 task with $n=17$ and $f=1$.
+Table 2 shows the worst-case maximal accuracies achieved by ARC and no clipping for four robust aggregations.
+For $\alpha=0.2$, ARC consistently outputs accuracies greater than 67% for all aggregations, while the same aggregations without clipping yield lower accuracies (with a larger variance across seeds).
+For instance, GM achieves 41.2% on average, i.e., 26% less than its counterpart with ARC.
+In the more heterogeneous setting $\alpha=0.075$, ARC enables all aggregations to reach worst-case maximal accuracies close to 60% (with a small variance).
+However, executing the same methods without clipping significantly deteriorates the performance of Robust-DSGD, with GM and CWMed achieving 16% and 13.7% in worst-case accuracy, respectively.
+This can also be seen in Figure 5, where FOE completely degrades the learning when ARC is not used.
 $\alpha=0.2$ $\alpha=0.075$
 Aggregation No Clipping ARC No Clipping ARC
-CWTM $\circ$ NNM 51.6 $\pm$ 5.1 67.8 $\pm$ 0.9 40.7 $\pm$ 0.5 60.5 $\pm$ 1.2
-GM $\circ$ NNM 41.2 $\pm$ 3.5 67.0 $\pm$ 1.0 16.0 $\pm$ 2.3 60.0 $\pm$ 2.0
-CWMed $\circ$ NNM 43.4 $\pm$ 4.2 69.4 $\pm$ 0.8 13.7 $\pm$ 12.1 62.7 $\pm$ 1.2
-MK $\circ$ NNM 50.9 $\pm$ 5.1 68.7 $\pm$ 0.5 40.5 $\pm$ 0.7 59.9 $\pm$ 1.9
+CWTM 51.6 $\pm$ 5.1 67.8 $\pm$ 0.9 40.7 $\pm$ 0.5 60.5 $\pm$ 1.2
+GM 41.2 $\pm$ 3.5 67.0 $\pm$ 1.0 16.0 $\pm$ 2.3 60.0 $\pm$ 2.0
+CWMed 43.4 $\pm$ 4.2 69.4 $\pm$ 0.8 13.7 $\pm$ 12.1 62.7 $\pm$ 1.2
+MK 50.9 $\pm$ 5.1 68.7 $\pm$ 0.5 40.5 $\pm$ 0.7 59.9 $\pm$ 1.9
 Table 1: Worst-case maximal accuracies (%) achieved by Robust-DSGD on heterogeneously-distributed CIFAR-10 with ARC and without.
@@
 As a baseline, DSGD ($f=0$) reaches 76.5% and 70% when $\alpha=0.2$ and 0.075, respectively.
-SOTA robust aggregations fail in extreme heterogeneity.
-Figure 3b shows that for all values of $f\in\{1,...,5\}$, Robust-DSGD with CWTM $\circ$ NNM and GM $\circ$ NNM completely fails to learn, consistently yielding worst-case maximal accuracies close to 15%.
-This suggests that $f=1$ constitutes the breakdown point for these aggregations in extreme heterogeneity.
-However, composing them with ARC increases the breakdown point of these aggregations to $f=3$.
-Indeed, for $f=1$ and 2, ARC enables CWTM $\circ$ NNM and GM $\circ$ NNM to achieve accuracies greater than 85% in the worst case.
-However, when $f\geq 3$, their performance degrades, although CWTM $\circ$ NNM $\circ$ ARC reaches a satisfactory accuracy close to 70%.
-Moreover, even when the heterogeneity is not large, ARC still produces a significant improvement when the fraction of adversarial workers increases in the system.
-In Figure 1 of Section 1 where $\alpha=0.5$, the performances of ARC and no clipping are comparable for $f\leq 3$ for both aggregations.
-Figure 4: FOE [43] on heterogeneous CIFAR-10 with $\alpha=0.075$. $f=1$ among $n=17$.
-However, when $f\geq 5$, the improvement induced by ARC is much more visible.
-Particularly when $f=7$, ARC enables CWTM $\circ$ NNM and GM $\circ$ NNM to reach accuracies greater than 80% in the worst case, whereas the same aggregations without clipping yield accuracies below 40%.
-This observation confirms the raise of the breakdown point caused by ARC to robust aggregations in practice.
-Plots for $\alpha=0.1$ and 1 convey the same observations (see Appendix E.1).
-Improved robustness on CIFAR-10.
-We also conduct experiments on the more challenging CIFAR-10 task with $n=17$ and $f=1$.
-Table 2 shows the worst-case maximal accuracies achieved by ARC and no clipping for four robust aggregations.
-For $\alpha=0.2$, ARC consistently outputs accuracies greater than 67% for all aggregations, while the same aggregations without clipping yield lower accuracies (with a large variance across seeds).
-For instance, GM $\circ$ NNM achieves 41.2% on average, i.e., 26% less than its counterpart with ARC.
-In the more heterogeneous setting $\alpha=0.075$, ARC enables all aggregations to reach worst-case maximal accuracies close to 60% (with a small variance).
-On the contrary, executing the same methods without clipping deteriorates the performance of Robust-DSGD, with GM $\circ$ NNM and CWMed $\circ$ NNM reaching 16% and 13.7% in worst-case accuracy, respectively.
-This can also be seen in Figure 4, where FOE completely degrades the learning when ARC is not used, despite CWTM $\circ$ NNM reaching a peak of 40% at the start of the learning.
-7 Conclusion
-We have introduced a pre-aggregation clipping scheme, named adaptive robust clipping (ARC), that helps to circumvent the general limitations on the robustness to adversarial workers in distributed learning, under the standard heterogeneity model of $(G,B)$-gradient dissimilarity.
-We have obtained formal guarantees on the improvement in the learning error of state-of-the-art robust distributed gradient descent (Robust-DGD) when incorporating ARC.
-We have validated our theoretical findings through an exhaustive set of experiments on benchmark image classification tasks.
+5 Improved Guarantee of Robust-DGD with ARC
+In Section 4, we empirically demonstrated that ARC outperforms SOTA aggregation methods without clipping, despite these methods being theoretically proven to be optimal.
+This naturally raises the question: why does ARC, which shares similar convergence guarantees, significantly outperform plain Robust-DSGD?
+To address this, we now focus on the influence of model initialization on the robustness of aggregation methods, considering both theoretical results and empirical insights.
+5.1 Improvement of convergence guarantees
+In this section, we characterize the improvement of Robust-DGD with ARC over the lower bound recalled in Lemma 2.3.
+Specifically, we consider Algorithm 1 with aggregation $\mathbf{F}\circ\mathbf{ARC}$, i.e., in each learning step $t$, $R_{t}\coloneqq\mathbf{F}\circ\mathbf{ARC}\left({{g}^{\left({1}\right)}_{t}},\dots,\,{{g}^{\left({n}\right)}_{t}}\right)$.
+First, we establish in Lemma 5.1 a key property of ARC, which is crucial for deriving the improved stationarity error of Robust-DGD with ARC in Theorem 5.2.
+Lemma 5.1 (Bounded Aggregation Output).
+Let $\mathbf{F}$ be an $(f,\kappa)$-robust aggregation method. For any vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ and set $S\subset[n]$ such that $|S|=n-f$, the following holds true:
+$\left\lVert{\mathbf{F}\circ\mathbf{ARC}(x_{1},\dots,x_{n})}\right\rVert\leq\max_{i\in S}\left\lVert{x_{i}}\right\rVert$ (1)
+This result indicates that incorporating ARC bounds the norm of the aggregated output by the largest norm of the honest gradients, i.e., $\max_{i\in\mathcal{H}}\left\lVert{x_{i}}\right\rVert$.
+In Appendix C, we provide the proof of Lemma 5.1 and further demonstrate in Lemma C.1 that this property is specific to ARC, i.e., classic $(f,\kappa)$-robust aggregation methods do not inherently exhibit this behavior.
+This property enables us to obtain the following improvement on the asymptotic convergence guarantee of Robust-DGD with ARC.
+In the following, we show that when the local gradients of the honest workers at the initial model ${{\theta}_{1}}$ are sufficiently small, then Robust-DGD with ARC overcomes the lower bound recalled in 2.3.
+For ease of presentation, we denote $\mathsf{BP}\coloneqq\frac{1}{2+B^{2}}$, let
+$\quad\varepsilon_{o}\coloneqq\frac{1}{4}\cdot\frac{G^{2}(f/n)}{1-(2+B^{2})(f/n)},\quad\text{ and }$ $\Psi(G,B,\rho)\coloneqq 640\left(1+\frac{1}{B^{2}}\right)^{2}\left(1+\frac{B^{2}\rho^{2}}{G^{2}}\right),$
+where $\rho$ denotes a real value.
+Recall from Lemma 2.3 that $\mathsf{BP}$ and $\varepsilon_{o}$ are the breakdown point and the lower bound on the stationarity error when $f/n<\mathsf{BP}$, respectively, under $(G,B)$-gradient dissimilarity.
+We obtain the following theorem for Robust-DGD with ARC.
+Let $\Delta_{o}$ be a real value such that $\mathcal{L}_{\mathcal{H}}({{\theta}_{1}})-\mathcal{L}_{\mathcal{H}}^{*}\leq\Delta_{o}$.
+The proof of Theorem 5.2 is deferred to Appendix C.
+Theorem 5.2.
+Suppose $B>0$ and there exists $\zeta\in\mathbb{R}^{+}$ such that $\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{1}})}\right\rVert\leq\zeta$.
+Let $\mathbf{F}\in\left\{\mathbf{CWTM},\mathbf{CWMed},\mathbf{GM},\mathbf{MK}\right\}$ $\circ\ \mathbf{NNM}$ , $\gamma=\min\left\{\left(\frac{\Delta_{o}}{\kappa G^{2}}\right)\frac{1}{T},~{}\frac{1}{L}\right\}$ and $T\geq\frac{\Delta_{o}L}{\kappa G^{2}}$.
+Consider an arbitrary real value $\xi_{o}\in(0,1)$.
+Let $\rho\coloneqq\exp\left(\frac{(2+B^{2})\Delta_{o}}{(1-\xi_{o})G^{2}}L\right)\zeta$.
+For any $\boldsymbol{\upsilon}\in(0,1)$, if $\frac{f}{n}\coloneqq(1-\xi)\mathsf{BP}$, where $0<\xi\leq\min\left\{\frac{\boldsymbol{\upsilon}}{\Psi(G,B,\rho)},\xi_{o}\right\}$, then $\mathop{{}\mathbb{E}}\left[{\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}\left(\hat{{{\theta}}}\right)}\right\rVert^{2}}\right]\leq\boldsymbol{\upsilon}~{}\varepsilon_{o}$.
+Theorem 5.2 demonstrates that Robust-DGD with ARC improves over the lower bound $\varepsilon_{o}$ when the ratio $f/n$ is sufficiently close to the breakdown point ($\mathsf{BP}$), provided the local gradients at model initialization are bounded.
+This, in turn, leads to an improvement over the original learning guarantee of Robust-DGD (as stated in Lemma 2.6), which applies for arbitrary model initialization.
+Influence of model initialization and data heterogeneity.
+Note that the smaller the bound on the initial gradients (i.e., the better the model initialization), the greater the improvement induced by ARC.
+Specifically, a decrease in $\zeta$ leads to a reduction in $\rho$, which subsequently lowers $\Psi(G,B,\rho)$.
+As a result, for a fixed value of $\xi$ (i.e., for a fixed ratio of adversarial workers), the condition $\xi\leq\boldsymbol{\upsilon}/\Psi(G,B,\rho)$ is satisfied for a smaller $\boldsymbol{\upsilon}$, thereby yielding a lower stationarity error.
+This theoretical deduction is empirically validated in Section 5.2.
+A similar improvement occurs when increasing data heterogeneity, while keeping all other factors unchanged.
+Specifically, as $G$ increases, $\Psi(G,B,\rho)$ decreases, allowing for a smaller reduction factor $\boldsymbol{\upsilon}$ and thus a larger improvement in performance.
+This trend is also validated empirically in Figures 3 and 4a.
+Influence of $\boldsymbol{f/n}$.
+Additionally, for a fixed model initialization, an increase of $f/n$ towards the $\mathsf{BP}$ (i.e., as $\xi\to 0$) allows the reduction factor $\boldsymbol{\upsilon}$ to become smaller, leading to a greater improvement, as evidenced in Figures 2a and 4b.
+Indeed, we show in Theorem C.4 (a more complete version of Theorem 5.2) that ARC effectively increases the $\mathsf{BP}$ of ($f,\kappa$)-robust methods from $\frac{1}{2+B^{2}}$ to $\frac{1}{2}$, provided the initial honest gradients are bounded in norm.
+Lastly, we would like to recall that when the workers’ gradients at model initialization are arbitrarily large, the convergence guarantee of Robust-DGD with ARC reduces to that of classic Robust-DGD (see Section 3).
+In other words, there is no downside to incorporating ARC.
+5.2 Influence of model initialization on empirical robustness
+Figure 6: Worst-case maximal accuracies achieved by Robust-DSGD with CWTM, both with and without ARC, on MNIST with $\alpha=0.1$, 10 honest worker, and $f=1$ Byzantine worker.
+The x-axis represents settings with progressively worse model initialization.
+This figure complements Figure 2b.
+Figures 2b and 6 illustrate the effect of model initialization on the performance of Robust-DSGD with and without ARC, evaluated on MNIST with 10 honest workers.
+We investigate two regimes of heterogeneity: extreme heterogeneity and $\alpha=0.1$, and consider $f=1$ adversarial worker.
+The experiment proceeds as follows: we begin with the default model parameters initialized by PyTorch (i.e., the same ones used in Section 4), which represent a well-chosen set of initial parameters.
+These parameters are then scaled multiplicatively by a factor $\mu$ where larger values of $\mu$ correspond to progressively worse initialization, and vary $\mu\in\{1,...,5\}$.
+The results show that, under well-initialized conditions ($\mu=1$), ARC significantly enhances the performance of Robust-DSGD, achieving a substantial improvement in worst-case maximal accuracy, particularly under extreme heterogeneity.
+In this regime, ARC boosts accuracy by about 70% compared to plain Robust-DSGD (Figure 2b).
+This ($\mu=1$) corresponds to the initialization conditions of the empirical results presented in Section 4.
+As $\mu$ increases (i.e., the initialization worsens), the performance of ARC-enhanced Robust-DSGD gradually declines, with noticeable degradation starting from $\mu=4$.
+Nevertheless, even at this point, ARC still offers a performance advantage over Robust-DSGD without clipping.
+By $\mu=5$, the performance of Robust-DSGD with ARC converges to that of plain Robust-DSGD, both achieving an accuracy of around 10%.
+Another key observation from Figures 2b and 6 is that the behavior of Robust-DSGD with ARC closely mirrors that of Byzantine-free DSGD.
+Both exhibit similarly poor performance when $\mu=5$, and their accuracies are comparable when $\mu\leq 3$, especially when $\alpha=0.1$.
+This suggests that ARC is particularly effective at exploiting good model initialization, similar to the performance of DSGD in the absence of Byzantine workers.
+In contrast, plain Robust-DSGD struggles to fully leverage well-initialized models, as evidenced by its consistently lower accuracy (around 20%) in Figure 2b.
+These findings highlight the important influence of model initialization on the robustness of aggregation methods, and empirically validate our theoretical findings in Section 5.1.
+6 Related Work
+Gradient clipping is a well-known technique for tackling exploding gradients in deep learning (Goodfellow et al., 2016).
+It has been extensively analyzed in the centralized setting (Zhang et al., 2020b; a; Koloskova et al., 2023), with applications to differential privacy (Pichapati et al., 2019).
+Moreover, we would like to note that gradient clipping has also been shown useful to obtain tighter convergence guarantees for stochastic gradient-descent methods in the case where the gradients have heavy-tailed noise Gorbunov et al.
+(2020; 2022); Danilova (2023).
+The considered proof techniques, however, do not apply directly to our adversarial distributed learning setting.
+A similar study on the impact of clipping also exists for distributed settings (Zhang et al., 2022; Khirirat et al., 2023).
+In the context of robust distributed learning, prior work has proposed iterative clipping for robust aggregation (Karimireddy et al., 2021).
+The clipping scheme, however, has not been shown to induce any improvement over other SOTA robustness schemes.
+Recent work (Malinovsky et al., 2023) has proposed pre-aggregation clipping using temporal gradient differences, in conjunction with variance reduction, to tackle partial worker participation when adversarial workers can form a majority.
+We, however, propose and study pre-aggregation clipping as a tool to improve the robustness obtained by a general class of aggregation rules.
+Other prior work (Mhamdi et al., 2021; Farhadkhani et al., 2022; Allouah et al., 2023a) that used static pre-aggregation clipping to enhance the empirical robustness of Robust-DGD, did not provide any formal explanation.
+While adaptive clipping schemes, similar to ARC, have been proposed and studied in the context of robust aggregation (Gupta & Vaidya, 2020; Liu et al., 2021; Diakonikolas et al., 2020; Abdalla & Zhivotovskiy, 2024), critical differences should be noted.
+The robustness guarantees in (Gupta & Vaidya, 2020; Liu et al., 2021) only apply to strongly convex loss functions, under a specific redundancy condition (namely, $2f$-redundancy).
+We consider non-convex loss functions, as well as a generic heterogeneous setting of $(G,B)$-gradient dissimilarity.
+In contrast to the clipping scheme proposed in Diakonikolas et al.
+(2020); Abdalla & Zhivotovskiy (2024), ARC only uses the $f/n$ parameter to tune the clipping threshold and does not rely on any additional a priori knowledge of the distribution of honest workers’ gradients.
+Moreover, we prove a deterministic robustness property of ARC (see Theorem 3.2), whereas Diakonikolas et al.
+(2020); Abdalla & Zhivotovskiy (2024) only provide probability guarantees assuming the non-adversarial inputs to be i.i.d.
+with the distribution satisfying certain special properties.
+7 Conclusion & Discussion
+We introduced Adaptive Robust Clipping (ARC), a pre-aggregation clipping scheme designed to harness the empirical benefits of gradient clipping, while preserving the worst-case optimal convergence guarantees of Robust-DGD.
+Unlike existing adaptive clipping schemes, ARC does not require additional tuning since it only relies on the standard robustness parameter, i.e., the tolerable fraction of adversarial workers.
+Through theoretical analysis, we explained ARC’s ability to enhance the robustness of Robust-DGD, particularly when the model is well-initialized.
+This phenomenon was validated through comprehensive experiments on standard image classification datasets.
+In short, our experiments showed that ARC consistently boosts the performance of Robust-DGD, especially in scenarios with high heterogeneity and large fraction of adversarial workers.
+Future direction.
+Our work also reveals a gap between theory and practice in Byzantine machine learning (ML).
+While Robust-DGD and ARC-enhanced Robust-DGD share the same worst-case convergence guarantees, their empirical performances are drastically different.
+In fact, and unsurprisingly, Byzantine ML theory focuses on worst-case guarantees, which, although essential, may not fully capture the practical realities of many learning settings.
+In practice, we often operate under favorable conditions (e.g., well-initialized models), where worst-case guarantees have limited relevance.
+This gap opens the door for future work that prioritizes practically-driven research in Byzantine ML, under conditions that are realizable in real-world scenarios.
+It encourages the development of robust distributed learning methods, like ARC, that can take full advantage of favorable practical conditions, thereby yielding relevant theoretical guarantees and superior empirical performance.
 Appendix A Computational Complexity of ARC.
@@
 The sorting operation of the norms takes $\mathcal{O}(n\log(n))$ time.
-Note that if $n$ is large compared to ${k=\left\lfloor{2(f/n)(n-f)}\right\rfloor}$, one can use more efficient algorithms for finding the ${(k+1)}$-th largest element, such as quick-select [22] which has an average-case complexity of $\mathcal{O}(n)$.
+Note that if $n$ is large compared to ${k=\left\lfloor{2(f/n)(n-f)}\right\rfloor}$, one can use more efficient algorithms for finding the ${(k+1)}$-th largest element, such as quick-select Hoare (1961) which has an average-case complexity of $\mathcal{O}(n)$.
 Finally, clipping the vectors requires $\mathcal{O}(nd)$ time.
@@
 B.1 Proof of Lemma 3.1
-See 4.1
+See 3.1
 Proof.
@@
 This brings us to the proof of Theorem 3.2, which we recall below for convenience.
-See 4.2
+See 3.2
 Proof.
@@
 Consider first the case $\kappa=0$.
-Since $\kappa\geq\frac{f}{n-2f}$  [^6] Proposition 6, [3], we have that $f=0$.
+Since $\kappa\geq\frac{f}{n-2f}$  [^5] Proposition 6, Allouah et al.
+(2023a), we have that $f=0$.
 Recall that we denote $y_{i}\coloneq\mathrm{clip}_{C}(x_{i})$ and $\overline{y}_{S}\coloneqq\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}y_{i}$.
@@
 Using the above in conjunction with lemma B.2 for the case $\left\lVert{\overline{x}_{S}}\right\rVert>C$ in (16) proves the result for the second case, which concludes the proof. ∎
-Appendix C Proofs for Results in Section 5
-Proofs for Theorem 5.1, Corollary 5.2, Theorem 5.4 and Corollary 5.5 are presented in appendices C.2C.3C.4 and C.5, respectively.
-C.1 Assuming $\left\lVert{\mathbf{F}\left(x_{1},\ldots,x_{n}\right)}\right\rVert\leq\max_{i\in[n]}\left\lVert{x_{i}}\right\rVert$ is Without Loss of Generality
-Recall that we assume $\left\lVert{\mathbf{F}\left(x_{1},\ldots,x_{n}\right)}\right\rVert\leq\max_{i\in[n]}\left\lVert{x_{i}}\right\rVert$ for all set of $n$ vectors $x_{1},\ldots,x_{n}\in\mathbb{R}^{d}$.
-In case this is not true, we can instead use the aggregation rule $\mathbf{F}^{\dagger}$ given by
-$\mathbf{F}^{\dagger}\left(x_{1},\ldots,x_{n}\right)\coloneqq\mathrm{clip}_{C}\left(\mathbf{F}\left(x_{1},\ldots,x_{n}\right)\right),~{}\text{where}~{}C=\max_{i\in[n]}\left\lVert{x_{i}}\right\rVert.$
-This modification to the aggregation rule does not affect the learning guarantee.
-Specifically, due to the non-expansion property of $\mathrm{clip}_{C}(\cdot)$, for any non-empty set $S\subseteq[n]$, we have
-$\left\lVert{\mathbf{F}^{\dagger}\left(x_{1},\ldots,x_{n}\right)-\overline{x}_{S}}\right\rVert\leq\left\lVert{\mathbf{F}\left(x_{1},\ldots,x_{n}\right)-\overline{x}_{S}}\right\rVert,$
-where $\overline{x}_{S}\coloneqq\frac{1}{\left\lvert{S}\right\rvert}\sum\limits_{i\in S}x_{i}$.
-Thus, if $\mathbf{F}$ is $(f,\kappa)$-robust, then $\mathbf{F}^{\dagger}$ is also $(f,\kappa)$-robust.
-Hence, we can make the above assumption on $\left\lVert{\mathbf{F}\left(x_{1},\ldots,x_{n}\right)}\right\rVert$ without loss of generality.
-C.2 Proof of Theorem 5.1
-Our proof relies on the following lemma.
+Appendix C Proofs of the Results in Section 5
+In this section, we first prove the Bounded Aggregation Output property of ARC in Lemma 5.1, and show that the other SOTA aggregation rules do not satisfy it in Lemma C.1.
+Then, to prove Theorem 5.2, we first prove Lemma C.2.
+We assume throughout this appendix that for any set of vectors $x_{1},\ldots,x_{n}\in\mathbb{R}^{d}$, $\left\lVert{\mathbf{F}\left(x_{1},\ldots,x_{n}\right)}\right\rVert\leq\max_{i\in[n]}\left\lVert{x_{i}}\right\rVert$.
+This assumption can be made without loss of generality, see Appendix C.1.
+See 5.1
+Proof.
+Without loss of generality and for the sake of simplicity, let us suppose that the vectors $x_{1},\dots,x_{n}$ are indexed such that $\left\lVert{x_{1}}\right\rVert\geq\left\lVert{x_{2}}\right\rVert\geq\dots\geq\left\lVert{x_{n}}\right\rVert$.
+Let $(\tilde{x}_{1},\dots,\tilde{x}_{n})=ARC{}(x_{1},\dots,x_{n})$ be the clipped version of these vectors after applying ARC.
+Case 1: When $f=0$, $S=[n]$ and we clip all the vectors by $\left\lVert{x_{1}}\right\rVert=\max_{i\in S}\left\lVert{x_{i}}\right\rVert$, hence for any $j$, $\left\lVert{\tilde{x}_{j}}\right\rVert\leq\max_{i\in S}\left\lVert{x_{i}}\right\rVert$.
+Using the fact that for any set of vectors $x_{1},\ldots,x_{n}\in\mathbb{R}^{d}$, $\left\lVert{\mathbf{F}\left(x_{1},\ldots,x_{n}\right)}\right\rVert\leq\max_{i\in[n]}\left\lVert{x_{i}}\right\rVert$, we have
+$\left\lVert{\mathbf{F}\circ ARC{}(x_{1},\dots,x_{n})}\right\rVert=\left\lVert{\mathbf{F}(\tilde{x}_{1},\dots,\tilde{x}_{n})}\right\rVert\leq\max_{j\in[n]}\left\lVert{\tilde{x}_{j}}\right\rVert\leq\max_{i\in S}\left\lVert{x_{i}}\right\rVert\enspace.$ (17)
+Case 2: When $0<f<\frac{n}{2}$, as presented in Algorithm 2, ARC clips all the vectors $x_{1},\dots,x_{n}$ to $\left\lVert{x_{k+1}}\right\rVert$ where $k=\lfloor 2\frac{f}{n}(n-f)\rfloor$.
+Hence, for any $j\in[n]$,
+$\left\lVert{\tilde{x}_{j}}\right\rVert\leq\left\lVert{x_{k+1}}\right\rVert\enspace.$ (18)
+We have that $k+1=\lfloor 2\frac{f}{n}(n-f)\rfloor+1\geq 2\frac{f}{n}(n-f)$.
+One can show that
+$\frac{2f}{n}(n-f)-f$ $=f(1-\frac{2f}{n})$
+$>0,\ \ \ \forall f\in\left]0,\frac{n}{2}\right[$ (19)
+Hence, $\frac{2f}{n}(n-f)>f$, which gives $k+1>f$, and since $k+1$ is an integer, we have $k+1\geq f+1$.
+Hence we know that at least $f+1$ vectors have their norm greater or equal to the clipping threshold $C=\left\lVert{x_{k+1}}\right\rVert$,
+$\underbrace{\left\lVert{x_{1}}\right\rVert\leq\dots\leq\left\lVert{x_{k+1}}\right\rVert}_{\text{at least $f+1$ vectors}}\leq\underbrace{\left\lVert{x_{k+2}}\right\rVert\leq\dots\leq\left\lVert{x_{n}}\right\rVert}_{\text{at most $n-f-1$ vectors}}$
+Hence, at most $n-f-1$ vectors will not be clipped.
+For any subset of $S\subset[n]$ of size $n-f$, at least one of the vectors has its norm equal to the clipping threshold (i.e.
+$\max_{i\in S}\left\lVert{x_{i}}\right\rVert=\left\lVert{x_{k+1}}\right\rVert$) or will be clipped (i.e.
+$\max_{i\in S}\left\lVert{x_{i}}\right\rVert>\left\lVert{x_{k+1}}\right\rVert$).
+Hence, we have that
+$\max_{i\in S}\left\lVert{x_{i}}\right\rVert\geq\left\lVert{x_{k+1}}\right\rVert\enspace.$ (20)
+Combining (18) and (20), we have that for all $j\in[n]$,
+$\left\lVert{\tilde{x}_{j}}\right\rVert\leq\max_{i\in S}\left\lVert{x_{i}}\right\rVert\enspace.$
+Using the fact that for any set of vectors $x_{1},\ldots,x_{n}\in\mathbb{R}^{d}$, $\left\lVert{\mathbf{F}\left(x_{1},\ldots,x_{n}\right)}\right\rVert\leq\max_{i\in[n]}\left\lVert{x_{i}}\right\rVert$, we have
+$\left\lVert{\mathbf{F}\circ ARC{}(x_{1},\dots,x_{n})}\right\rVert=\left\lVert{\mathbf{F}(\tilde{x}_{1},\dots,\tilde{x}_{n})}\right\rVert\leq\max_{j\in[n]}\left\lVert{\tilde{x}_{j}}\right\rVert\leq\max_{i\in S}\left\lVert{x_{i}}\right\rVert\enspace.$ (21)
+Lemma C.1.
+For any aggregation rule $\mathbf{F}\in\{\mathbf{CWTM},\mathbf{CWMed},\mathbf{GM},\mathbf{MK},\mathbf{CWTM}\circ\mathbf{NNM},\\ \mathbf{CWMed}\circ\mathbf{NNM},\mathbf{GM}\circ\mathbf{NNM},\mathbf{MK}\circ\mathbf{NNM}\}$, there exists $f\in[0,\frac{n}{2})$, a set of vectors $x_{1},\dots,x_{n}$ and a subset $S\subset[n]$, $|S|=n-f$ such that
+$\mathbf{F}(x_{1},\dots,x_{n})\nleq\max_{i\in S}\left\lVert{x_{i}}\right\rVert$ (22)
+Proof.
+Let us consider $(x_{1},x_{2},x_{3})=((0,1),(1,0),(1,1))$, and $S=\{x_{1},x_{2}\}$,
+For $\mathbf{F}{}\in\{\mathbf{CWTM},\mathbf{CWMed},\mathbf{GM}\}$, we have $\left\lVert{\mathbf{F}(x_{1},x_{2},x_{3})}\right\rVert=\left\lVert{(1,1)}\right\rVert=\sqrt{2}>\max_{i\in S}\left\lVert{x_{i}}\right\rVert=\left\lVert{x_{1}}\right\rVert=1$.
+For $\mathbf{F}{}\in\{\mathbf{MK},\mathbf{CWTM}\circ\mathbf{NNM},\mathbf{CWMed}\circ\mathbf{NNM},\mathbf{GM}\circ\mathbf{NNM},\mathbf{MK}\circ\mathbf{NNM}\}$, we have $\left\lVert{\mathbf{F}(x_{1},x_{2},x_{3})}\right\rVert=\left\lVert{(0.5,1)}\right\rVert=\sqrt{1.25}>\max_{i\in S}\left\lVert{x_{i}}\right\rVert=\left\lVert{x_{1}}\right\rVert=1$.
+∎
+Lemma C.2.
+Let $\mathbf{F}$ be $(f,\kappa)$-robust.
+Let $\Delta_{o}$ and $\rho$ be real values such that $\mathcal{L}_{\mathcal{H}}({{\theta}_{1}})-\mathcal{L}_{\mathcal{H}}^{*}\leq\Delta_{o}$ and $\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{1}})}\right\rVert\leq\exp\left(-\frac{\Delta_{o}}{\kappa G^{2}}L\right)\rho$.
+If $\gamma=\min\left\{\left(\frac{\Delta_{o}}{\kappa G^{2}}\right)\frac{1}{T},~{}\frac{1}{L}\right\}$, then
+$\frac{1}{T}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}({{\theta}_{t}})}\right\rVert^{2}\leq\frac{2\Delta_{o}L}{T}+5\kappa\left(G^{2}+B^{2}\rho^{2}\right).$
+Our proof for Lemma C.2 relies on the following sub-result.
 Lemma C.3.
@@
 Let $(x_{1},\ldots,\,x_{n})\coloneqq\mathbf{ARC}\left({{g}^{\left({1}\right)}_{t}},\ldots,~{}{{g}^{\left({n}\right)}_{t}}\right)$.
-As $n>2f$ and $\left\lvert{\mathcal{H}}\right\rvert=n-f$, the clipping threshold $C$ used in ARC (see Algorithm 2) is bounded by $\max_{i\in\mathcal{H}}\left\lVert{{{g}^{\left({i}\right)}_{t}}}\right\rVert$.
+As $n>2f$ and $\left\lvert{\mathcal{H}}\right\rvert=n-f$, the clipping threshold $C$ used in ARC (see Algorithm 2) is bounded by $\max_{i\in\mathcal{H}}\left\lVert{{{g}^{\left({i}\right)}_{t}}}\right\rVert$ (see (20) in Lemma 5.1).
 Therefore,
 $\max_{i\in[n]}\left\lVert{x_{i}}\right\rVert\leq\max_{i\in\mathcal{H}}\left\lVert{{{g}^{\left({i}\right)}_{t}}}\right\rVert=\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{t}})}\right\rVert.$ (23)
-Since we assume (without loss of generality) that, for any set of $n$ vectors $z_{1},\ldots,z_{n}\in\mathbb{R}^{d}$, $\left\lVert{\mathbf{F}\left(z_{1},\ldots,z_{n}\right)}\right\rVert\leq\max_{i\in[n]}\left\lVert{z_{i}}\right\rVert$ (see the discussion in Section 5.1), from (16) we obtain that
+Since for any set of $n$ vectors $z_{1},\ldots,z_{n}\in\mathbb{R}^{d}$, $\left\lVert{\mathbf{F}\left(z_{1},\ldots,z_{n}\right)}\right\rVert\leq\max_{i\in[n]}\left\lVert{z_{i}}\right\rVert$, from (23) we obtain that
 $\left\lVert{R_{t}}\right\rVert=\left\lVert{\mathbf{F}\circ\mathbf{ARC}\left({{g}^{\left({1}\right)}_{t}},\ldots,~{}{{g}^{\left({n}\right)}_{t}}\right)}\right\rVert\leq\max_{i\in[n]}\left\lVert{x_{i}}\right\rVert\leq\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{t}})}\right\rVert.$
@@
 Since $\left(1+\gamma L\right)^{T}\leq\exp{(\gamma LT)}$, the above concludes the proof. ∎
-We are now ready to present our proof of Theorem 5.1, recalled below.
-See 5.1
-Proof.
-For simplicity we write $\mathcal{L}_{\mathcal{H}}$ as $\mathcal{L}$ throughout the proof.
-Consider an arbitrary $t\in[T]$. Due to $L$-Lipschitz smoothness of $\mathcal{L}({{\theta}_{)}}$, we obtain that
+We are now ready to present our proof of Lemma C.2.
+Proof of Lemma C.2.
+For simplicity, we write $\mathcal{L}_{\mathcal{H}}$ as $\mathcal{L}$ throughout the proof.
+Consider an arbitrary $t\in[T]$. Due to $L$-Lipschitz smoothness of $\mathcal{L}({{\theta}})$, we obtain that
 $\mathcal{L}({{\theta}_{t+1}})\leq\mathcal{L}({{\theta}_{t}})+\left\langle{{\theta}_{t+1}}-{{\theta}_{t}},\,\nabla\mathcal{L}({{{\theta}_{t}}})\right\rangle+\frac{L}{2}\left\lVert{{{\theta}_{t+1}}-{{\theta}_{t}}}\right\rVert^{2}.$
@@
 Using this above we obtain that
-$\frac{1}{T}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}\leq\frac{2\Delta_{o}}{\gamma T}+\frac{3\kappa}{T}~{}\frac{1}{\left\lvert{\mathcal{H}}\right\rvert}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{t}})-\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}.$ (18)
+$\frac{1}{T}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}\leq\frac{2\Delta_{o}}{\gamma T}+\frac{3\kappa}{T}~{}\sum_{t=1}^{T}\frac{1}{\left\lvert{\mathcal{H}}\right\rvert}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{t}})-\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}.$ (25)
 Note that, since $\gamma\leq\frac{\Delta_{o}}{\kappa G^{2}T}$ and $\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{1}})}\right\rVert\leq\exp\left(-\frac{\Delta_{o}L}{\kappa G^{2}}\right)\rho$, by Lemma C.3 we have
@@
 Combining (27) and (28) concludes the proof. ∎
-C.3 Proof of Corollary 5.2
-See 5.2
-Proof.
-Since $T\geq\frac{\Delta_{o}L}{\kappa G^{2}}$, $\gamma=\frac{\Delta_{o}}{\kappa G^{2}T}\leq\frac{1}{L}$. Therefore, from Theorem 5.1 we obtain that
-$\frac{1}{T}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}\leq 5\left(G^{2}+B^{2}\rho^{2}\right).$ (22)
+We now prove Theorem C.4, stated below, which immediately implies Theorem 5.2.
+Specifically, Theorem 5.2 follows immediately from Part 1 of Theorem C.4, upon substituting $\xi\leq\frac{1}{\Psi(G,B,\rho)}\upsilon$.
+Theorem C.4.
+Suppose $B>0$ and there exists $\zeta\in\mathbb{R}^{+}$ such that $\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{1}})}\right\rVert\leq\zeta$.
+Let $\mathbf{F}\in\left\{\mathbf{CWTM},\mathbf{CWMed},\mathbf{GM},\mathbf{MK}\right\}$, $\gamma=\min\left\{\left(\frac{\Delta_{o}}{\kappa G^{2}}\right)\frac{1}{T},~{}\frac{1}{L}\right\}$ and $T\geq\frac{\Delta_{o}L}{\kappa G^{2}}$.
+Consider an arbitrary real value $\xi_{o}\in(0,1)$.
+Let $\rho\coloneqq\exp\left(\frac{(2+B^{2})\Delta_{o}}{(1-\xi_{o})G^{2}}L\right)\zeta$.
+Then, the following holds true:
+  1. 1.
+    If $\frac{f}{n}\coloneqq(1-\xi)\mathsf{BP}$, where $\xi\in(0,\xi_{o}]$, then, $\mathop{{}\mathbb{E}}\left[{\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}\left(\hat{{{\theta}}}\right)}\right\rVert^{2}}\right]\leq\xi\,\Psi(G,B,\rho)~{}\varepsilon_{o}$.
+  2. 2.
+    If $\mathsf{BP}\leq f/n<1/2$, then $\mathop{{}\mathbb{E}}\left[{\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}\left(\hat{{{\theta}}}\right)}\right\rVert^{2}}\right]\leq\Psi(G,B,\rho)\cdot\frac{G^{2}}{8}$
+In order to prove Theorem C.4, we first obtain the following corollary of Lemma C.2 and Lemma 2.6.
+Corollary C.5.
+For the parameters given in Lemma C.2, if $T\geq\frac{\Delta_{o}L}{\kappa G^{2}}$, then Robust-DGD $\circ$ ARC achieves
+$\mathop{{}\mathbb{E}}\left[{\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}\left(\hat{{{\theta}}}\right)}\right\rVert^{2}}\right]\leq 5\kappa~{}\min\left\{G^{2}+B^{2}\rho^{2},~{}\frac{G^{2}}{\max\{(1-\kappa B^{2}),0\}}\right\},$
+where $\mathop{{}\mathbb{E}}\left[{\cdot}\right]$ denotes the expectation over the choice of $\hat{{{\theta}}}$.
+Proof of Corollary C.5.
+For simplicity, we write $\mathcal{L}_{\mathcal{H}}$ as $\mathcal{L}$ throughout the proof.
+Since $T\geq\frac{\Delta_{o}L}{\kappa G^{2}}$, $\gamma=\frac{\Delta_{o}}{\kappa G^{2}T}\leq\frac{1}{L}$. Therefore, from Lemma C.2 we obtain that
+$\frac{1}{T}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}\leq 5\kappa\left(G^{2}+B^{2}\rho^{2}\right).$ (29)
 In the specific case when $\kappa B^{2}<1$, as $\gamma=\frac{\Delta_{o}}{\kappa G^{2}T}$, from Lemma 2.6 we obtain that
@@
 ∎
-C.4 Proof of Theorem 5.4
-Recall that we denote
-$\mathsf{BP}\coloneqq\frac{1}{2+B^{2}},\quad\varepsilon_{o}\left(f/n\right)\coloneqq\frac{1}{4}\cdot\frac{G^{2}(f/n)}{1-(2+B^{2})(f/n)},\quad\text{ and }$
+We are now ready to prove Theorem C.4. Recall that
+$\mathsf{BP}\coloneqq\frac{1}{2+B^{2}},\quad\varepsilon_{o}\coloneqq\frac{1}{4}\cdot\frac{G^{2}(f/n)}{1-(2+B^{2})(f/n)},\quad\text{ and }$
 $\Psi(G,B,\rho)\coloneqq 640\left(1+\frac{1}{B^{2}}\right)^{2}\left(1+\frac{B^{2}\rho^{2}}{G^{2}}\right).$
-We recall the theorem below for convenience.
-See 5.4
-Proof.
+Proof of Theorem C.4.
+For simplicity, we write $\mathcal{L}_{\mathcal{H}}$ as $\mathcal{L}$ throughout the proof.
 In the proof, we substitute $\mathsf{BP}=\frac{1}{2+B^{2}}$.
-First, we consider $\frac{f}{n}=\frac{1-\xi}{2+B^{2}}$. Note that, due to Lemma 3.3, $\mathbf{F}$ is $(f,\kappa)$-robust with
-$\kappa\leq\frac{16f}{n-f}\left(1+\frac{1-\xi}{B^{2}}\right)^{2}\leq\frac{16(1-\xi)}{1+B^{2}}\left(1+\frac{1-\xi}{B^{2}}\right)^{2}\leq\kappa_{o}(1-\xi),$
+We first consider the case when $\frac{f}{n}=\frac{1-\xi}{2+B^{2}}$. In this particular case,
+$n=\left(\frac{2}{1-\xi}+\frac{B^{2}}{1-\xi}\right)f\geq\left(2+\frac{B^{2}}{1-\xi}\right)f.$
+Therefore, thanks to Lemma 2.5, $\mathbf{F}$ is $(f,\kappa)$-robust with
+$\kappa\leq\frac{16f}{n-f}\left(1+\frac{1-\xi}{B^{2}}\right)^{2}=\frac{16(1-\xi)}{1+B^{2}+\xi}\left(1+\frac{1-\xi}{B^{2}}\right)^{2}\leq\frac{16(1-\xi)}{1+B^{2}}\left(1+\frac{1}{B^{2}}\right)^{2}\leq\kappa_{o}(1-\xi),$
 where $\kappa_{o}\coloneqq\frac{16}{1+B^{2}}\left(1+\frac{1}{B^{2}}\right)^{2}$. Also, recall from the limitations of $(f,\kappa)$-robustness in Section 2 that
 $\kappa\geq\frac{f}{n-2f}\geq\frac{f}{n}\geq\frac{1-\xi_{o}}{2+B^{2}}.$
-To summarize the above, we have
+Summarizing from above, we have
 $\frac{1-\xi_{o}}{2+B^{2}}\leq\kappa\leq\kappa_{o}(1-\xi).$ (31)
@@
 $\rho\coloneqq\exp\left(\frac{(2+B^{2})\Delta_{o}}{(1-\xi_{o})G^{2}}L\right)\zeta\geq\exp\left(\frac{\Delta_{o}}{\kappa G^{2}}L\right)\zeta.$
-Therefore, since $\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{1}})}\right\rVert\leq\zeta$ for all $i\in\mathcal{H}$ under Assumption 5.3, in Theorem 5.1 we always satisfy the condition: $\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{1}})}\right\rVert\leq\exp\left(-\frac{\Delta_{o}}{\kappa G^{2}}L\right)\rho$, for the defined value of $\rho$ Thus, by Corollary 5.2 and (24) we obtain that
+Therefore, since $\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{1}})}\right\rVert\leq\zeta$ for all $i\in\mathcal{H}$, the condition: $\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{1}})}\right\rVert\leq\exp\left(-\frac{\Delta_{o}}{\kappa G^{2}}L\right)\rho$ in Lemma C.2 is satisfied.
+Thus, by Corollary C.5 and (31) we obtain that
 $\mathop{{}\mathbb{E}}\left[{\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}}\right]\leq 5\kappa_{o}(1-\xi)(G^{2}+B^{2}\rho^{2})=\frac{5\kappa_{o}(1-\xi)(G^{2}+B^{2}\rho^{2})}{\frac{fG^{2}}{n-(2+B^{2})f}}\,\frac{fG^{2}}{n-(2+B^{2})f}.$ (32)
@@
 $\leq 160\left(1+\frac{1}{B^{2}}\right)^{2}\left(1+\frac{B^{2}\rho^{2}}{G^{2}}\right)\xi.$
-Substituting from above in (25) concludes the proof for the first part of the theorem.
-The second part of the theorem follows immediately from the first inequality in (25), using the fact that $\kappa_{o}(1-\xi)\leq\kappa_{o}$. This concludes the proof. ∎
-C.5 Proof of Corollary 5.5
-See 5.5
-Proof.
-The proof follows immediately from upon using $\xi\leq\frac{1}{\Psi(G,B,\rho)}\upsilon$ in Part 1 of Theorem 5.4. ∎
+Substituting from above in (32) concludes the proof for the first part of Theorem C.4.
+The second part of Theorem C.4 follows immediately from the first inequality in (32), using the fact that $\kappa_{o}(1-\xi)\leq\kappa_{o}$. This concludes the proof. ∎
+C.1 Assuming $\left\lVert{\mathbf{F}\left(x_{1},\ldots,x_{n}\right)}\right\rVert\leq\max_{i\in[n]}\left\lVert{x_{i}}\right\rVert$ is Without Loss of Generality
+Recall that we assume $\left\lVert{\mathbf{F}\left(x_{1},\ldots,x_{n}\right)}\right\rVert\leq\max_{i\in[n]}\left\lVert{x_{i}}\right\rVert$ for all set of $n$ vectors $x_{1},\ldots,x_{n}\in\mathbb{R}^{d}$.
+In case this is not true, we can instead use the aggregation rule $\mathbf{F}^{\dagger}$ given by
+$\mathbf{F}^{\dagger}\left(x_{1},\ldots,x_{n}\right)\coloneqq\mathrm{clip}_{C}\left(\mathbf{F}\left(x_{1},\ldots,x_{n}\right)\right),~{}\text{where}~{}C=\max_{i\in[n]}\left\lVert{x_{i}}\right\rVert.$
+This modification to the aggregation rule does not affect the learning guarantee.
+Specifically, due to the non-expansion property of $\mathrm{clip}_{C}(\cdot)$, for any non-empty set $S\subseteq[n]$, we have
+$\left\lVert{\mathbf{F}^{\dagger}\left(x_{1},\ldots,x_{n}\right)-\overline{x}_{S}}\right\rVert\leq\left\lVert{\mathbf{F}\left(x_{1},\ldots,x_{n}\right)-\overline{x}_{S}}\right\rVert,$
+where $\overline{x}_{S}\coloneqq\frac{1}{\left\lvert{S}\right\rvert}\sum\limits_{i\in S}x_{i}$.
+Thus, if $\mathbf{F}$ is $(f,\kappa)$-robust, then $\mathbf{F}^{\dagger}$ is also $(f,\kappa)$-robust.
+Hence, we can make the above assumption on $\left\lVert{\mathbf{F}\left(x_{1},\ldots,x_{n}\right)}\right\rVert$ without loss of generality.
 Appendix D Comprehensive Experimental Setup
+In this section, we present the comprehensive experimental setup considered in our paper.
 D.1 Datasets and Heterogeneity
-In our experiments, we consider three standard image classification datasets, namely MNIST [11], Fashion-MNIST [42], and CIFAR-10 [31].
-To simulate data heterogeneity in our experiments, we make the honest workers sample from the datasets using a Dirichlet [23] distribution of parameter $\alpha$, as done in [24, 3, 13].
+In our experiments, we consider three standard image classification datasets, namely MNIST Deng (2012), Fashion-MNIST Xiao et al.
+(2017), and CIFAR-10 Krizhevsky et al.
+(2014).
+To simulate data heterogeneity in our experiments, we make the honest workers sample from the datasets using a Dirichlet Hsu et al.
+(2019a) distribution of parameter $\alpha$, as done in Hsu et al.
+(2019b); Allouah et al.
+(2023a); Farhadkhani et al.
+(2023).
 The smaller the $\alpha$, the more heterogeneous the setting.
@@
 D.2 Algorithm, Distributed System, ML Models, and Hyperparameters
-We perform our experiments using Robust-DSGD (see Algorithm 3), which is known to be an order-optimal robust variant of Robust-DGD [3].
+We perform our experiments using Robust-DSGD (see Algorithm 3), which is known to be an order-optimal robust variant of Robust-DGD Allouah et al.
+(2023a).
 On MNIST and Fashion-MNIST, we execute Robust-DSGD in a distributed system composed of $n-f=10$ honest workers, and $f\in\{1,2,3,4,5,7,9\}$ adversarial workers.
@@
 Finally, we use the NLL loss function with an $\ell_{2}$ regularization of $10^{-2}$.
-Algorithm 3 Robust Stochastic Distributed Gradient Descent (Robust-DSGD)
+Algorithm 3 Robust Distributed Stochastic Gradient Descent (Robust-DSGD)
   Input: Server chooses an initial model ${{\theta}_{1}}\in\mathbb{R}^{d}$, learning rate $\gamma\in\mathbb{R}^{+}$, robust aggregation $\mathbf{F}:\mathbb{R}^{n\times d}\to\mathbb{R}^{d}$, momentum parameter $\beta\in\mathbb{R}^{+}$.
@@
   Output: Server outputs $\hat{{{\theta}}}$ chosen uniformly at random from $\{{{\theta}_{1}},\ldots,\,{{\theta}_{T}}\}$.
-In order to present the architectures of the ML models used in our experiments, we adopt the following compact terminology introduced in [3].
+In order to present the architectures of the ML models used in our experiments, we adopt the following compact terminology introduced in Allouah et al. (2023a).
 L(#outputs) represents a fully-connected linear layer, R stands for ReLU activation, S stands for log-softmax, C(#channels) represents a fully-connected 2D-convolutional layer (kernel size 5, padding 0, stride 1), M stands for 2D-maxpool (kernel size 2), B stands for batch-normalization, and D represents dropout (with fixed probability 0.25).
@@
 D.3 Byzantine Attacks
-In our experiments, the adversarial workers execute five state-of-the-art adversarial attacks from the Byzantine ML literature, namely sign-flipping (SF) [2], label-flipping (LF) [2], mimic [27], fall of empires (FOE) [43], and a little is enough (ALIE) [6].
+In our experiments, the adversarial workers execute five state-of-the-art adversarial attacks from the Byzantine ML literature, namely sign-flipping (SF) Allen-Zhu et al.
+(2020), label-flipping (LF) Allen-Zhu et al.
+(2020), mimic Karimireddy et al.
+(2022), fall of empires (FOE) Xie et al.
+(2020), and a little is enough (ALIE) Baruch et al.
+(2019).
 The exact functionality of the attacks is detailed below.
@@
   • Mimic: the adversarial workers mimic a certain honest worker by sending its gradient to the server.
-In order to determine the optimal honest worker for the adversarial workers to mimic, we use the heuristic in [27].
+In order to determine the optimal honest worker for the adversarial workers to mimic, we use the heuristic in Karimireddy et al.
+(2022).
   • FOE: the adversarial workers send $(1-\tau)\overline{m}_{t}$ in step $t$ to the server, where $\tau\geq 0$ is a fixed real number representing the attack factor.
@@
 In this section, we present empirical results on static clipping when used as a pre-aggregation technique in Robust-DSGD, and compare its performance to our proposed adaptive algorithm ARC.
-(a) Sign-flipping (SF) [2]
-(b) Label-flipping (LF) [2]
-Figure 15: Impact of pre-aggregation clipping, specifically static clipping with $C=2$ and ARC, on robust-DGD with aggregation scheme $\mathbf{CWTM}\circ\mathbf{NNM}$, under 2 adversarial attacks: (a) SF and (b) LF.
-We set $n=13$, $f=3$, and consider the MNIST [11] dataset distributed amongst $n-f=10$ honest workers with extreme heterogeneity.
 F.1 Limitations of Static Clipping
@@
 Moreover, the negative log likelihood loss function is used, along with an $\ell_{2}$-regularization of $10^{-4}$.
-On CIFAR-10, we execute Robust-DSGD on ResNet-18 [21], in a distributed system comprising $n=9$ workers among which $f\in\{1,2\}$ are Byzantine.
+On CIFAR-10, we execute Robust-DSGD on ResNet-18 He et al.
+(2015), in a distributed system comprising $n=9$ workers among which $f\in\{1,2\}$ are Byzantine.
 We set $b=128$, $T=2000$, $\beta=0.9$, and $\gamma=0.1$ decaying once by 10$\times$ at step 1500.
@@
 First, the efficacy of static clipping is significantly influenced by the type of Byzantine attack executed during the learning process.
-To illustrate, while $C=2$ exhibits the best performance under the SF attack [2], depicted in Figure 15, it leads to a complete collapse of learning under LF [2].
+To illustrate, while $C=2$ exhibits the best performance under the SF attack Allen-Zhu et al.
+(2020), depicted in Figure 1, it leads to a complete collapse of learning under LF Allen-Zhu et al.
+(2020).
 Consequently, identifying a static clipping threshold $C$ that consistently performs well in practice against any attack is difficult (if not impossible).
@@
 ARC exhibits robust performance across diverse Byzantine scenarios, encompassing variations in both the type of Byzantine attack and the number of Byzantine workers $f$.
-As depicted in Figures 15 and 16, ARC consistently yields robust performance, regardless of the Byzantine attack.
+As depicted in Figures 1 and 17, ARC consistently yields robust performance, regardless of the Byzantine attack.
 In extreme heterogeneity with $f=3$, ARC maintains an accuracy of 75%, while $C=2$ demonstrates a subpar worst-case accuracy of 25% (also refer to the right plot of Figure 1).

### v5

@@
-The Vital Role of Gradient Clipping in Byzantine-Resilient Distributed Learning
+Adaptive Gradient Clipping for Robust Federated Learning
 Abstract
-Byzantine-resilient distributed machine learning seeks to achieve robust learning performance in the presence of misbehaving or adversarial workers.
+Robust federated learning aims to maintain reliable performance despite the presence of adversarial or misbehaving workers.
 While state-of-the-art (SOTA) robust distributed gradient descent (Robust-DGD) methods were proven theoretically optimal, their empirical success has often relied on pre-aggregation gradient clipping.
-However, the currently considered static clipping strategy exhibits mixed results: improving robustness against some attacks while being ineffective or detrimental against others.
-We address this gap by proposing a principled adaptive clipping strategy, termed Adaptive Robust Clipping (ARC).
-We show that ARC consistently enhances the empirical robustness of SOTA Robust-DGD methods, while preserving the theoretical robustness guarantees.
-Our analysis shows that ARC provably improves the asymptotic convergence guarantee of Robust-DGD in the case when the model is well-initialized.
-We validate this theoretical insight through an exhaustive set of experiments on benchmark image classification tasks.
-We observe that the improvement induced by ARC is more pronounced in highly heterogeneous and adversarial settings.
+However, existing static clipping strategies yield inconsistent results: enhancing robustness against some attacks while being ineffective or even detrimental against others.
+To address this limitation, we propose a principled adaptive clipping strategy, Adaptive Robust Clipping (ARC), which dynamically adjusts clipping thresholds based on the input gradients.
+We prove that ARC not only preserves the theoretical robustness guarantees of SOTA Robust-DGD methods but also provably improves asymptotic convergence when the model is well-initialized.
+Extensive experiments on benchmark image classification tasks confirm these theoretical insights, demonstrating that ARC significantly enhances robustness, particularly in highly heterogeneous and adversarial settings.
 1 Introduction
@@
 Due to the growing influence of distributed learning in critical public-domain applications such as healthcare (Nguyen et al., 2022) and finance (Long et al., 2020), the problem of robustness to misbehaving workers, a.k.a.
-robust distributed learning, has received significant attention (Blanchard et al., 2017; Yin et al., 2018; Farhadkhani et al., 2022; Karimireddy et al., 2022; Gorbunov et al., 2023; Allouah et al., 2023a; Farhadkhani et al., 2023).
+robust distributed learning, has received significant attention (Yin et al., 2018; Farhadkhani et al., 2022; Karimireddy et al., 2022; Gorbunov et al., 2023; Allouah et al., 2023a; Farhadkhani et al., 2023; El-Mhamdi et al., 2023).
 Robust distributed learning algorithms primarily rely on robust aggregation, such as coordinate-wise trimmed mean (CWTM) (Yin et al., 2018), geometric median (GM) (Chen et al., 2017) and Multi-Krum (MK) (Blanchard et al., 2017).
 Specifically, in robust distributed gradient descent (Robust-DGD), the server aggregates the workers’ local gradients using a robust aggregation method, instead of simply averaging them.
-This protects the learning procedure from erroneous gradients sent by misbehaving workers.
-Recent work has made significant improvements over these aggregation rules by incorporating a pre-aggregation step such as bucketing (Karimireddy et al., 2022; Gorbunov et al., 2023) and nearest-neighbor mixing (NNM) (Allouah et al., 2023a), to tackle gradient dissimilarity resulting from data heterogeneity.
-The learning guarantee of the resulting Robust-DGD has been proven to be optimal (Allouah et al., 2023b), i.e., it cannot be improved without additional assumptions under the standard heterogeneity model of $(G,B)$-gradient dissimilarity (Karimireddy et al., 2020).
+This protects the learning from erroneous gradients sent by misbehaving workers.
+Recent work has made significant improvements over these aggregation techniques by incorporating a pre-aggregation step such as bucketing (Karimireddy et al., 2022; Gorbunov et al., 2023) and nearest-neighbor mixing (NNM) (Allouah et al., 2023a), to tackle gradient dissimilarity resulting from data heterogeneity.
+The learning guarantee of the resulting Robust-DGD has been proven optimal (Allouah et al., 2023b), i.e., it cannot be improved without additional assumptions under the standard heterogeneity model of $(G,B)$-gradient dissimilarity (Karimireddy et al., 2020).
 Despite its theoretical tightness, the empirical success of Robust-DGD has unknowingly relied on pre-aggregation gradient clipping (Mhamdi et al., 2021; Farhadkhani et al., 2022; Allouah et al., 2023a).
-Specifically, clipping the gradients of the workers prior to aggregation has been observed to sometimes enhance the algorithm’s empirical performance in the presence of adversarial workers, as evidenced in Figure 1a.
+Specifically, clipping the gradients of the workers prior to aggregation has been observed to sometimes enhance the algorithm’s empirical performance in the presence of adversarial workers (as evidenced in Figure 7(a)).
 Yet, this improvement lacks a concrete explanation, raising the question of whether the observed benefits of clipping are merely anecdotal.
 This leads to the natural inquiry: Why, and when, does pre-aggregation clipping improve robustness?
-(a) Sign-flipping (SF) (Allen-Zhu et al., 2020)
-(b) Label-flipping (LF) (Allen-Zhu et al., 2020)
-Figure 1: Impact of pre-aggregation clipping, specifically static clipping with $C=2$ and our adaptive clipping algorithm ARC, on Robust-DGD with aggregation scheme CWTM $\circ$ NNM, under the SF and LF adversarial attacks.
-We consider the MNIST (Deng, 2012) dataset distributed amongst $10$ honest (non-adversarial) workers with extreme heterogeneity, and there are three adversarial workers.
 In particular, using a constant clipping threshold, referred to as static clipping, has exhibited mixed results.
@@
 ARC dynamically adjusts the clipping threshold as per the gradients sent by the workers and the fraction of adversarial workers to be tolerated.
-We demonstrate that integrating ARC into Robust-DGD consistently improves its empirical performance (see Figures 1 and 2a), while also preserving the convergence guarantee of the original Robust-DGD algorithm.
+We demonstrate that integrating ARC into Robust-DGD consistently improves its empirical performance (see Figures 7 and 1(a)), while also preserving the convergence guarantee of the original Robust-DGD algorithm.
 Moreover, we show that when the model initialization is good, ARC provably improves the robustness of Robust-DGD.
-The benefits of ARC are more pronounced as the fraction of misbehaving workers approaches the system’s breakdown point [^1] Breakdown point refers the minimum fraction of adversarial workers that can break the system, making it impossible to guarantee a bound on the learning error Allouah et al.
-(2023b).
+The benefits of ARC are more pronounced as the fraction of misbehaving workers approaches the system’s breakdown point [^1] Breakdown point refers to the minimum fraction of adversarial workers that can break the system, making it impossible to guarantee a bound on the learning error (Allouah et al., 2023b).
 and when the data across the workers is highly heterogeneous.
@@
 Critical comparisons to prior work are deferred to Section 6.
+(a) Improved robustness
+(b) Influence of model initialization
+Figure 1: Worst-case maximal accuracies of Robust-DSGD, with and without ARC, across several types of misbehavior for distributed MNIST with $10$ honest workers and under extreme heterogeneity.
+On the left, we vary the number of adversarial workers.
+On the right, we vary the initialization conditions by scaling a well-chosen set of initial parameters (CWTM $\circ$ NNM is used, and $f=1$).
+More details on the experimental setup can be found in Sections 4 and 5.2, and Appendix D.
 Main results & contributions. We consider a system comprising $n$ workers and a server. The goal is to tolerate up to $f$ adversarial workers.
@@
 We conduct experiments on MNIST (Deng, 2012), Fashion-MNIST (Xiao et al., 2017), and CIFAR-10 (Krizhevsky et al., 2014), across various data heterogeneity settings and adversarial regimes.
-Our results demonstrate that ARC significantly enhances the performance of state-of-the-art Robust-DGD methods, particularly in scenarios with high data heterogeneity (Figure 2a) and a large number of adversarial workers (Figure 4b).
+Our results demonstrate that ARC significantly enhances the performance of state-of-the-art Robust-DGD methods, particularly in scenarios with high data heterogeneity (Figure 1(a)) and a large number of adversarial workers (Figure 4(b)).
 (3) Improved learning guarantee.
-We demonstrate that ARC possesses an additional property, we call Bounded Aggregation Output, that is not satisfied by classical robust aggregation methods.
+We demonstrate that ARC possesses an additional property that is not satisfied by classical robust aggregation methods.
 Specifically, ARC constrains the norm of an adversarial gradient by that of an honest (non-adversarial) gradient.
-Leveraging this property, we show that the lower bound established under data heterogeneity in Allouah et al.
-(2023b) can be circumvented using ARC, provided the honest workers’ gradients are bounded at model initialization.
-An empirical validation of this insight is shown in Figure 2b.
+Leveraging this property, we show that ARC circumvents the lower bound established under data heterogeneity in Allouah et al.
+(2023b), provided the honest gradients are bounded at model initialization.
+An empirical validation of this insight is shown in Figure 1(b).
 Such model initialization is often satisfiable in practice (Glorot & Bengio, 2010), highlighting the practical relevance of ARC.
-When the model is arbitrarily initialized, ARC recovers the original convergence guarantee of Robust-DGD in the worst case (see Figure 2b).
-(a) Improved robustness
-(b) Influence of model initialization
-Figure 2: Worst-case maximal accuracies of Robust-DSGD, with and without ARC, across several types of misbehavior for distributed MNIST with $10$ honest workers and under extreme heterogeneity.
-On the left, we vary the number of adversarial workers.
-On the right, we vary the initialization conditions by scaling a well-chosen set of initial parameters (CWTM $\circ$ NNM is used, and $f=1$).
-More details on the experimental setup can be found in Sections 4 and 5.2, and Appendix D.
+When the model is arbitrarily initialized, ARC recovers the original convergence guarantee of Robust-DGD in the worst case.
 2 Problem Statement and Relevant Background
 […]
-A robust distributed learning algorithm aims to output a good model despite the presence of adversarial (a.k.a., Byzantine (Lamport et al., 1982)) workers in the system (Su & Vaidya, 2016; Yin et al., 2018; Allouah et al., 2023a).
+A robust distributed learning algorithm aims to output a good model despite the presence of adversarial (a.k.a., Byzantine) workers in the system (Su & Vaidya, 2016; Yin et al., 2018; Allouah et al., 2023a).
 Specifically, the goal is to design a distributed learning algorithm that can tolerate up to $f$ adversarial workers, of a priori unknown identity, out of $n$ workers.
 Adversarial workers need not follow a prescribed algorithm, and can send arbitrary information to the server.
-In the context of distributed gradient descent, adversarial workers can send incorrect gradients to the server (Baruch et al., 2019; Allen-Zhu et al., 2020; Xie et al., 2020; Karimireddy et al., 2022).
+In the context of distributed gradient descent, adversarial workers can send incorrect gradients to the server (Baruch et al., 2019; Xie et al., 2020).
 We let $\mathcal{H}\subseteq[n]$, with $\left\lvert{\mathcal{H}}\right\rvert=n-f$, denote the set of honest workers that do not deviate from the algorithm.
-The objective of the server is to minimize the average loss function of honest workers given by
-$\mathcal{L}_{\mathcal{H}}(\theta)\coloneqq\frac{1}{\left\lvert{\mathcal{H}}\right\rvert}\sum_{i\in\mathcal{H}}\mathcal{L}_{i}(\theta),~{}\forall\theta\in\mathbb{R}^{d}.$
+The objective of the server is to minimize the average loss function of honest workers given by $\mathcal{L}_{\mathcal{H}}(\theta)\coloneqq\frac{1}{\left\lvert{\mathcal{H}}\right\rvert}\sum_{i\in\mathcal{H}}\mathcal{L}_{i}(\theta),~{}\forall\theta\in\mathbb{R}^{d}$.
 While we can find a minimum of $\mathcal{L}_{\mathcal{H}}(\theta)$ when $\mathcal{L}_{\mathcal{H}}$ is convex, in general however the loss function is non-convex, and the optimization problem is NP-hard (Boyd & Vandenberghe, 2004).
@@
 A distributed learning algorithm $\mathcal{A}$ is said to be $(f,\varepsilon)$-resilient if, despite the presence of $f$ adversarial workers, $\mathcal{A}$ enables the server to output a model $\hat{\theta}$ such that $\mathop{{}\mathbb{E}}\left[{\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}\left(\hat{\theta}\right)}\right\rVert^{2}}\right]\leq\varepsilon$, where the expectation $\mathop{{}\mathbb{E}}\left[{\cdot}\right]$ is over the randomness of the algorithm.
-If a distributed learning algorithm $\mathcal{A}$ is $(f,\varepsilon)$-resilient then it can tolerate up to $f$ adversarial workers.
-It is standard in robust distributed learning to design robust algorithms using $f$ as a parameter (Su & Vaidya, 2016; Blanchard et al., 2017; Yin et al., 2018; Karimireddy et al., 2022; Gorbunov et al., 2023; Allouah et al., 2023a), $f$ being the maximum tolerable number of adversarial workers.
+If a distributed learning algorithm $\mathcal{A}$ is $(f,\varepsilon)$-resilient, then it can tolerate up to $f$ adversarial workers.
+It is standard in robust distributed learning to design robust algorithms using $f$ as a parameter (Blanchard et al., 2017; Yin et al., 2018; Gorbunov et al., 2023; Allouah et al., 2023a).
 In the context of distributed learning, data heterogeneity is characterized by the following standard notion of $(G,B)$-gradient dissimilarity (Vaswani et al., 2019; Karimireddy et al., 2020; 2022; Gorbunov et al., 2023; Allouah et al., 2023b).
@@
 This yields Robust-DGD, presented in Algorithm 1, where robust aggregation protects the learning from incorrect gradients sent by the adversarial workers.
-The robustness of an aggregation method can be quantified by the following property of $(f,\kappa)$-robustness, proposed by Allouah et al.
-(2023a).
+The robustness of an aggregation method can be quantified by the following property of $(f,\kappa)$-robustness (Allouah et al., 2023a).
 Algorithm 1 Robust Distributed Gradient Descent (Robust-DGD)
@@
 $\left\lVert{\mathbf{F}(x_{1},\ldots,x_{n})-\overline{x}_{S}}\right\rVert^{2}\leq\frac{\kappa}{\left\lvert{S}\right\rvert}\sum_{i\in S}\left\lVert{x_{i}-\overline{x}_{S}}\right\rVert^{2}\enspace,\quad\text{where }~{}\overline{x}_{S}=\frac{1}{\left\lvert{S}\right\rvert}\sum\limits_{i\in S}x_{i}.$
-$(f,\kappa)$-robustness [^2] Unifies $(\delta_{\text{max}},c)$-robustness (Karimireddy et al., 2021) and $(f,\lambda)$-resilience (Farhadkhani et al., 2022).
-encompasses several robust aggregations, including the ones mentioned above and more (e.g., see Allouah et al.
+$(f,\kappa)$-robustness encompasses several robust aggregations, including the ones mentioned above and more (e.g., see Allouah et al.
 (2023a)).
 It has been shown that an aggregation method is $(f,\kappa)$-robust only if $\kappa\geq\frac{f}{n-2f}$ (Allouah et al., 2023a).
-Importantly, the asymptotic error of Robust-DGD with $(f,\kappa)$-robust aggregation, where $\kappa\in\mathcal{O}(f/(n-2f))$, matches the lower bound (recalled in Lemma 2.3(Allouah et al., 2023b).
+Importantly, the asymptotic error of Robust-DGD with an $(f,\kappa)$-robust aggregation, where $\kappa\in\mathcal{O}(f/(n-2f))$, matches the lower bound (recalled in Lemma 2.3(Allouah et al., 2023b).
+This testifies to the tightness of the $(f,\kappa)$-robustness property.
 Note that the aforementioned aggregation methods (CWTM, CWMed, GM and MK) attain optimal robustness when composed with the pre-aggregation scheme: nearest neighbor mixing (NNM) (Allouah et al., 2023a).
-[^3] CWTM has been shown to be optimal even without NNM.
 Specifically, we recall the following result.
 Lemma 2.5 (Lemma 1 in Allouah et al. (2023a)).
-For $\mathbf{F}\in\left\{\mathbf{CWTM},\mathbf{CWMed},\mathbf{GM},\mathbf{MK}\right\}$, the composition $\mathbf{F}\circ\mathbf{NNM}$ is $(f,\kappa)$-robust with
-$\kappa\leq\frac{8f}{n-f}\left(1+\left(1+\frac{f}{n-2f}\right)^{2}\right).$
+For $\mathbf{F}\in\left\{\mathbf{CWTM},\mathbf{CWMed},\mathbf{GM},\mathbf{MK}\right\}$, the composition $\mathbf{F}\circ\mathbf{NNM}$ is $(f,\kappa)$-robust with $\kappa\leq\frac{8f}{n-f}\left(1+\left(1+\frac{f}{n-2f}\right)^{2}\right)$.
 Thus, if $n\geq(2+\delta)f$ for $\delta>0$, then $\mathbf{F}\circ\mathbf{NNM}$ is $(f,\kappa)$-robust with $\kappa\leq\frac{16f}{n-f}\left(\frac{\delta+1}{\delta}\right)^{2}\in\mathcal{O}\left(\frac{f}{n-2f}\right)$.
@@
 Thus, when $\kappa\in\mathcal{O}\left(f/(n-2f)\right)$ (and smaller than $1/B^{2}$), Robust-DGD is optimal, i.e., its error matches the lower bound recalled in Lemma 2.3, as soon as $T\geq\Delta_{o}/\gamma\kappa G^{2}$.
-[^4] For $f=0$, Lemma 2.3 recovers the convergence of DGD, provided that $\kappa$ is tight, i.e., $\kappa\in\mathcal{O}(f/(n-2f))$.
+For $f=0$, Lemma 2.6 recovers the convergence of DGD, provided that $\kappa$ is tight, i.e., $\kappa\in\mathcal{O}(\frac{f}{n-2f})$.
 3 Adaptive Robust Clipping (ARC) and its Properties
@@
 Thus, if the clipping threshold is fixed, i.e., independent from the input vectors, pre-aggregation clipping does not preserve the robustness of the original aggregation.
-This fragility of such static clipping is also apparent in practice, as shown by our experimental study in Appendix F (and Figure 1).
+This fragility of such static clipping is also apparent in practice, as shown by our experimental study in Section 4 and Appendix E.
 Description of ARC.
-ARC is an adaptive clipping scheme that only makes use of the standard robustness parameter $\nicefrac{{f}}{{n}}$, i.e., the tolerable fraction of adversarial workers.
+ARC is an adaptive clipping scheme that only makes use of the standard robustness parameter $f$, i.e., the tolerable number of adversarial workers.
 ARC is adaptive in the sense that the clipping threshold $C$ is not fixed but depends on the input vectors.
 Specifically, ARC clips the largest $k=\left\lfloor{2(f/n)(n-f)}\right\rfloor$ vectors using a clipping parameter given by the norm of the $(k+1)$-th largest input vector.
-The overall scheme is formally presented in Algorithm 2, and its computational complexity is $\mathcal{O}(nd+n\log(n))$ (see Appendix A).
-Thus, pre-composing more computationally expensive schemes like NNM and multi-Krum, which have a complexity of $\mathcal{O}(dn^{2})$, with ARC does not introduce a significant overhead.
+The overall scheme is formally presented in Algorithm 2, and its computational complexity is $\mathcal{O}(nd+n\log(n))$ (see Appendix A.2).
+Thus, pre-composing more computationally expensive schemes like NNM and Multi-Krum, which have a complexity of $\mathcal{O}(dn^{2})$, with ARC does not introduce a significant overhead (see Appendix F.4).
+For a detailed explanation of the underlying intuition behind our adaptive clipping strategy, ARC, we refer the reader to Appendix A.1.
 Algorithm 2 Adaptive Robust Clipping (ARC)
@@
 We present below the preservation of robustness guaranteed by ARC.
-Let $\mathbf{F}$ be an $(f,\kappa)$-robust aggregation rule and $\mathbf{F}\circ\mathbf{ARC}(x_{1},\dots,x_{n})\coloneqq\mathbf{F}(\mathbf{ARC}(x_{1},\dots,x_{n}))$
+Let $\mathbf{F}$ be an $(f,\kappa)$-robust aggregation method and $\mathbf{F}\circ\mathbf{ARC}(x_{1},\dots,x_{n})\coloneqq\mathbf{F}(\mathbf{ARC}(x_{1},\dots,x_{n}))$
 Theorem 3.2.
@@
 Proofs for the results presented in this section are deferred to Appendix B.
-Since $\kappa\geq\frac{f}{n-2f}$ (recalled in Section 2), Theorem 3.2 implies that $\mathbf{F}\circ\mathbf{ARC}$ is $\left(f,3\kappa\right)$-robust.
-In other words, ARC preserves the robustness of the original aggregation scheme.
+Since $\kappa\geq\frac{f}{n-2f}$ (recalled in Section 2), Theorem 3.2 implies that $\mathbf{F}\circ\mathbf{ARC}$ is $\left(f,3\kappa\right)$-robust, i.e., ARC preserves the robustness of the original aggregation scheme.
 Therefore, a convergence result for Robust-DGD with ARC follows verbatim from Lemma 2.6, replacing $\kappa$ with $3\kappa$.
-Despite this multiplicative factor of 3, we observe in the next section that incorporating ARC consistently improves the empirical performance of classical aggregation methods.
-A more detailed theoretical explanation will be provided later in Section 5.1.
+Despite this multiplicative factor of 3, we observe in Section 4 that incorporating ARC consistently improves the empirical performance of classical aggregation methods.
+A more detailed theoretical explanation is provided later in Section 5.1.
 4 Empirical Evaluation
@@
 We empirically test four aggregation methods when pre-composed with ARC.
-We also contrast these outcomes against the performance of Robust-DSGD when no gradient clipping is used.
-Our findings underscore that clipping workers’ gradients using ARC prior to the aggregation is crucial to ensure the robustness of existing algorithms, especially in extreme scenarios, i.e., when either the data heterogeneity is high or the fraction of adversarial workers is large.
+We contrast these outcomes against the performance of these aggregations under static clipping and when no gradient clipping is used.
+We show in Table 1, and Figures 2 and 4, the metric of worst-case maximal accuracy.
+For each of the five Byzantine attacks executed (see Appendix D.3), we record the maximal accuracy achieved by Robust-DSGD during the learning procedure under that attack.
+The worst-case maximal accuracy is thus the smallest maximal accuracy reached across the five attacks.
+Furthermore, we use the Dirichlet (Hsu et al., 2019a) distribution of parameter $\alpha$ to simulate data heterogeneity.
+The comprehensive experimental setup can be found in Appendix D.
+In this section, we focus on presenting our results for the CWTM and GM aggregation methods applied to MNIST and CIFAR-10.
+As benchmark, we also execute the standard DSGD algorithm, in the absence of adversarial workers (i.e., without attack and $f=0$).
+Results for Fashion-MNIST, as well as for CWMed and MK, are provided in Appendices E and  F.
+All aggregation methods in our experiments are pre-composed with NNM (Allouah et al., 2023a), but we omit explicit mention of NNM in their names for simplicity.
+4.1 Brittleness of Static Clipping and Superiority of ARC
+(a) $f=3$
+(b) $f=4$
+Figure 2: Impact of the clipping strategy, under varying heterogeneity levels, on the worst-case maximal accuracy achieved by Robust-DSGD on MNIST with $n=15$ workers.
+CWTM $\circ$ NNM is used as aggregation.
+DSGD reaches at least 98.5% in accuracy in all heterogeneity regimes.
+We empirically demonstrate the brittleness of static clipping in robust distributed learning.
+While an exhaustive empirical study can be found in Appendix E, we present findings on MNIST using CWTM, considering three heterogeneity levels: moderate ($\alpha=1$), high ($\alpha=0.1$), and extreme.
+We execute Robust-DSGD with $n=15$ workers, among which $f\in\{3,4\}$ are adversarial, and test static clipping thresholds $C\in\{0.02,0.2,2,20\}$.
+First, we observe that the optimal $C$ depends heavily on data heterogeneity.
+As seen in Figure 2, $C=2$ is ideal in high heterogeneity, whereas $C=20$ performs much better under extreme heterogeneity.
+This dependency requires fine-tuning, which is impractical in distributed settings.
+Moreover, static clipping’s effectiveness varies significantly across Byzantine attacks.
+While $C=2$ performs well under SF (Allen-Zhu et al., 2020), it fails under LF, causing training collapse (see Figure 7 in Appendix E).
+This unpredictability makes static clipping unreliable in practice.
+The number of adversarial workers further affects static clipping’s performance.
+In Figure 2, $C=2$ is optimal for $f=3$ under high heterogeneity, yet its accuracy drops below 50% when $f=4$, making it ineffective.
+Unlike static clipping, ARC consistently matches or outperforms the best static threshold across all configurations, as shown in Figure 2.
+These results underscore the superiority of ARC over static clipping, eliminating the need for manual tuning and ensuring robustness regardless of unpredictable parameters like data heterogeneity and Byzantine attacks.
+4.2 Performance Gains of ARC Over Robust-DSGD
+To quantify the improvement induced by ARC over Robust-DSGD without clipping, we evaluate a distributed system with $n-f=10$ honest workers while varying the number of adversarial workers $f$.
+We compare the performance of Robust-DSGD with and without ARC on the MNIST dataset, highlighting the improvements achieved by our method.
+Additionally, we extend our analysis to larger distributed systems with 30 honest workers and $f\in\{3,6,9\}$, as detailed in Appendix F.1.2.
+ARC boosts robustness in high heterogeneity.
+Figure 3 shows the performance of Robust-DSGD against the FOE attack when using ARC opposed to no clipping.
+In low heterogeneity (i.e., $\alpha=0.5$), the performances of ARC and no clipping are comparable.
+When the heterogeneity increases ($\alpha=0.1$), CWTM and GM significantly struggle to learn, while the same aggregations composed with ARC almost match the final accuracy of DSGD.
+In extreme heterogeneity, the improvement induced by ARC is the most pronounced, enabling both aggregations to reach a final accuracy close to 90%.
+Contrastingly, the same aggregations without clipping stagnate at around 10% throughout the training.
+Interestingly, only $f=1$ adversarial worker among $n=11$ (i.e., less than 10%) suffices to completely deteriorate the learning when no clipping is applied, highlighting the necessity of ARC in extreme heterogeneity.
+These observations are also conveyed in Figure 4(a) which compares the worst-case maximal accuracies achieved by Robust-DSGD when $f=1$, with and without ARC, for varying levels of heterogeneity.
+While CWTM and GM yield comparable accuracies with ARC in low heterogeneity ($\alpha\geq 0.5$), their performance significantly degrades when $\alpha$ drops below that threshold.
+Indeed, when $\alpha=0.1$, their accuracies drop below 65%, while ARC enables the same aggregations to maintain their accuracy at just below 98%.
+In extreme heterogeneity, the performance of the aggregations without clipping completely deteriorates with accuracies close to 15%.
+Contrastingly, ARC efficiently mitigates the Byzantine attacks, resulting in accuracies above 85% in the worst case for both aggregations.
+Similar plots for $f\in\{3,5,7\}$ convey similar observations (see Appendix F.1).
 (a) $\alpha=0.5$
@@
 Figure 3: Performance of Robust-DSGD when using ARC and without clipping on MNIST. There are 10 honest workers and $f=1$ adversarial worker executing FOE (Xie et al., 2020).
-We show in Table 1, and Figures 2a and 4, the metric of worst-case maximal accuracy.
-For each of the five Byzantine attacks executed (see Appendix D.3), we record the maximal accuracy achieved by Robust-DSGD during the learning procedure under that attack.
-The worst-case maximal accuracy is thus the smallest maximal accuracy reached across the five attacks.
-As the attack cannot be known in advance, this metric is critical to accurately evaluate the robustness of aggregation methods, since it provides an estimate of the potential worst-case performance of the algorithm.
-Furthermore, we use the Dirichlet (Hsu et al., 2019a) distribution of parameter $\alpha$ to simulate data heterogeneity.
-The comprehensive experimental setup can be found in Appendix D.
-In this section, we focus on presenting our results for the CWTM and GM aggregation methods applied to MNIST and CIFAR-10.
-Results for Fashion-MNIST, as well as for the CWMed and MK aggregations, are provided in Appendix E.
-All aggregation methods in our experiments are pre-composed with NNM (Allouah et al., 2023a), but we omit explicit mention of NNM in their names throughout the text for simplicity.
+ARC increases the breakdown point in adverse settings.
+Figure 1(a) of Section 1 shows that for $f\in\{1,...,5\}$, Robust-DSGD with CWTM and GM completely fails to learn, consistently yielding worst-case maximal accuracies close to 15%.
+This suggests that $f=1$ constitutes the breakdown point for these aggregations in extreme heterogeneity.
+However, composing them with ARC increases the breakdown point of these aggregations to $f=3$.
+Indeed, for $f=1$ and 2, ARC enables Robust-DSGD to achieve accuracies greater than 85% in the worst case.
+However, when $f\geq 3$, the performance degrades, although CWTM with ARC is still able to reach a satisfactory accuracy close to 70%.
+Moreover, even when the heterogeneity is not large ($\alpha=0.5$), ARC still produces a significant improvement when the fraction of adversarial workers increases in the system.
+Indeed, in Figure 4(b) , the performances of ARC and no clipping are comparable for $f\leq 3$.
+However, the improvement induced by ARC is much more visible when $f\geq 5$.
+Particularly when $f=7$, ARC enables CWTM and GM to reach accuracies greater than 80% in the worst case, whereas the same aggregations yield accuracies below 40% without clipping, indicating the raise of the breakdown point due to ARC.
+Plots for $\alpha=0.1$ and 1 convey the same observations in Appendix F.1.
 (a) $f=1$
@@
 Right: $\alpha=0.5$ for varying $f$.
-ARC boosts robustness in high heterogeneity.
-Figure 3 shows the performance of Robust-DSGD against the FOE attack when using ARC opposed to no clipping.
-In low heterogeneity (i.e., $\alpha=0.5$, see Figure 7), the performances of ARC and no clipping are comparable for both aggregations.
-When the heterogeneity increases ($\alpha=0.1$), the benefits of using ARC are more visible.
-CWTM and GM significantly struggle to learn (with a very large variance), while the same aggregations composed with ARC almost match the accuracy of DSGD towards the end of the training.
-In extreme heterogeneity, the improvement induced by our method is the most pronounced, as ARC enables both aggregations to reach a final accuracy close to 90%.
-Contrastingly, the same aggregations without clipping stagnate at around 10% throughout the training.
-Interestingly, only $f=1$ adversarial worker among $n=11$ (i.e., less than 10%) suffices to completely deteriorate the learning when no clipping is applied, highlighting the necessity of ARC in extreme heterogeneity.
-Varying heterogeneity for fixed $f$.
-Figure 4a compares the worst-case maximal accuracies achieved by Robust-DSGD when $f=1$, with ARC and without, for varying levels of heterogeneity.
-While CWTM and GM yield comparable accuracies with ARC in low heterogeneity ($\alpha\geq 0.5$), their performance significantly degrades when $\alpha$ drops below that threshold.
-Indeed, when $\alpha=0.1$, their accuracies drop below 65%, while ARC enables the same aggregations to maintain their accuracy at just below 98%.
-In extreme heterogeneity, the performance of the aggregations without clipping completely deteriorates with accuracies close to 15%.
-Contrastingly, ARC efficiently mitigates the Byzantine attacks, resulting in accuracies above 85% in the worst case for both aggregations.
-Similar plots for $f=3,5,7$ convey similar observations (see Appendix E.1).
-Essentially, in low heterogeneity or when the fraction of adversarial workers is small, no clipping yields good empirical results and ARC performs at least as well as no clipping.
-However, ARC induces a significant improvement near the breakdown point of SOTA robust aggregation schemes (e.g., high heterogeneity or large $f$).
-ARC increases the breakdown point in adverse settings.
-Figure 2a of Section 1 shows that for $f\in\{1,...,5\}$, Robust-DSGD with CWTM and GM completely fails to learn, consistently yielding worst-case maximal accuracies close to 15%.
-This suggests that $f=1$ constitutes the breakdown point for these aggregations in extreme heterogeneity.
-However, composing them with ARC increases the breakdown point of these aggregations to $f=3$.
-Indeed, for $f=1$ and 2, ARC enables Robust-DSGD to achieve accuracies greater than 85% in the worst case.
-However, when $f\geq 3$, the performance degrades, although CWTM with ARC is still able to reach a satisfactory accuracy close to 70%.
-Moreover, even when the heterogeneity is not large, ARC still produces a significant improvement when the fraction of adversarial workers increases in the system.
-Indeed, in Figure 4b , the performances of ARC and no clipping are comparable for $f\leq 3$.
-However, the improvement induced by ARC is much more visible when $f\geq 5$.
-Particularly when $f=7$, ARC enables CWTM and GM to reach accuracies greater than 80% in the worst case, whereas the same aggregations yield accuracies below 40% without clipping .
-This observation indicates the raise of the breakdown point due to ARC in practice.
-Plots for $\alpha=0.1$ and 1 convey the same observations in Appendix E.1.
-Figure 5: FOE on heterogeneous CIFAR-10 with $\alpha=0.075$. $f=1$ among $n=17$.
 Improved robustness on CIFAR-10.
-We also conduct experiments on the more challenging CIFAR-10 task with $n=17$ and $f=1$.
-Table 2 shows the worst-case maximal accuracies achieved by ARC and no clipping for four robust aggregations.
-For $\alpha=0.2$, ARC consistently outputs accuracies greater than 67% for all aggregations, while the same aggregations without clipping yield lower accuracies (with a larger variance across seeds).
+We also conduct experiments on CIFAR-10 with $n=17$ and $f=1$.
+Table 1 shows the worst-case maximal accuracies achieved by ARC and no clipping for CWTM and GM.
+For $\alpha=0.2$, ARC consistently outputs accuracies greater than 67% for both aggregations, while no clipping yields lower accuracies (with a larger variance).
 For instance, GM achieves 41.2% on average, i.e., 26% less than its counterpart with ARC.
 In the more heterogeneous setting $\alpha=0.075$, ARC enables all aggregations to reach worst-case maximal accuracies close to 60% (with a small variance).
-However, executing the same methods without clipping significantly deteriorates the performance of Robust-DSGD, with GM and CWMed achieving 16% and 13.7% in worst-case accuracy, respectively.
-This can also be seen in Figure 5, where FOE completely degrades the learning when ARC is not used.
+However, executing the same methods without clipping significantly deteriorates the performance of Robust-DSGD, with GM achieving 16% in worst-case accuracy.
+This can also be seen in Figure 23(a), where FOE completely degrades the learning when ARC is not used.
+We further extend this experiment to a larger distributed system with 33 honest workers, as detailed in Appendix F.3, where we observe the same trends as reported here.
 $\alpha=0.2$ $\alpha=0.075$
@@
 GM 41.2 $\pm$ 3.5 67.0 $\pm$ 1.0 16.0 $\pm$ 2.3 60.0 $\pm$ 2.0
-CWMed 43.4 $\pm$ 4.2 69.4 $\pm$ 0.8 13.7 $\pm$ 12.1 62.7 $\pm$ 1.2
-MK 50.9 $\pm$ 5.1 68.7 $\pm$ 0.5 40.5 $\pm$ 0.7 59.9 $\pm$ 1.9
 Table 1: Worst-case maximal accuracies (%) achieved by Robust-DSGD on heterogeneously-distributed CIFAR-10 with ARC and without.
@@
 5.1 Improvement of convergence guarantees
-In this section, we characterize the improvement of Robust-DGD with ARC over the lower bound recalled in Lemma 2.3.
+In this section, we characterize the improvement induced by ARC over the lower bound recalled in Lemma 2.3.
 Specifically, we consider Algorithm 1 with aggregation $\mathbf{F}\circ\mathbf{ARC}$, i.e., in each learning step $t$, $R_{t}\coloneqq\mathbf{F}\circ\mathbf{ARC}\left({{g}^{\left({1}\right)}_{t}},\dots,\,{{g}^{\left({n}\right)}_{t}}\right)$.
-First, we establish in Lemma 5.1 a key property of ARC, which is crucial for deriving the improved stationarity error of Robust-DGD with ARC in Theorem 5.2.
-Lemma 5.1 (Bounded Aggregation Output).
-Let $\mathbf{F}$ be an $(f,\kappa)$-robust aggregation method. For any vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ and set $S\subset[n]$ such that $|S|=n-f$, the following holds true:
-$\left\lVert{\mathbf{F}\circ\mathbf{ARC}(x_{1},\dots,x_{n})}\right\rVert\leq\max_{i\in S}\left\lVert{x_{i}}\right\rVert$ (1)
+We establish in Lemma 5.1 a key property of ARC, crucial to derive the improved stationarity error of Robust-DGD with ARC in Theorem 5.2.
+Lemma 5.1 (Bounded Output).
+Let $\mathbf{F}$ be an $(f,\kappa)$-robust aggregation method.
+For any vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ and set $S\subset[n]$ such that $|S|=n-f$, $\left\lVert{\mathbf{F}\circ\mathbf{ARC}(x_{1},\dots,x_{n})}\right\rVert\leq\max_{i\in S}\left\lVert{x_{i}}\right\rVert$.
 This result indicates that incorporating ARC bounds the norm of the aggregated output by the largest norm of the honest gradients, i.e., $\max_{i\in\mathcal{H}}\left\lVert{x_{i}}\right\rVert$.
@@
 This property enables us to obtain the following improvement on the asymptotic convergence guarantee of Robust-DGD with ARC.
-In the following, we show that when the local gradients of the honest workers at the initial model ${{\theta}_{1}}$ are sufficiently small, then Robust-DGD with ARC overcomes the lower bound recalled in 2.3.
-For ease of presentation, we denote $\mathsf{BP}\coloneqq\frac{1}{2+B^{2}}$, let
-$\quad\varepsilon_{o}\coloneqq\frac{1}{4}\cdot\frac{G^{2}(f/n)}{1-(2+B^{2})(f/n)},\quad\text{ and }$ $\Psi(G,B,\rho)\coloneqq 640\left(1+\frac{1}{B^{2}}\right)^{2}\left(1+\frac{B^{2}\rho^{2}}{G^{2}}\right),$
+In the following, we show that when the local gradients of the honest workers at the initial model ${{\theta}_{1}}$ are sufficiently small, then Robust-DGD with ARC overcomes the lower bound recalled in Lemma 2.3.
+We denote $\mathsf{BP}\coloneqq\frac{1}{2+B^{2}}$, $\varepsilon_{o}\coloneqq\frac{1}{4}\cdot\frac{G^{2}(f/n)}{1-(2+B^{2})(f/n)},\text{ and }\Psi(G,B,\rho)\coloneqq 640\left(1+\frac{1}{B^{2}}\right)^{2}\left(1+\frac{B^{2}\rho^{2}}{G^{2}}\right)$,
 where $\rho$ denotes a real value.
@@
 Specifically, as $G$ increases, $\Psi(G,B,\rho)$ decreases, allowing for a smaller reduction factor $\boldsymbol{\upsilon}$ and thus a larger improvement in performance.
-This trend is also validated empirically in Figures 3 and 4a.
+This trend is also validated empirically in Figures 3 and 4(a).
 Influence of $\boldsymbol{f/n}$.
-Additionally, for a fixed model initialization, an increase of $f/n$ towards the $\mathsf{BP}$ (i.e., as $\xi\to 0$) allows the reduction factor $\boldsymbol{\upsilon}$ to become smaller, leading to a greater improvement, as evidenced in Figures 2a and 4b.
+Additionally, for a fixed model initialization, an increase of $f/n$ towards the $\mathsf{BP}$ (i.e., as $\xi\to 0$) allows the reduction factor $\boldsymbol{\upsilon}$ to become smaller, leading to a greater improvement, as evidenced in Figures 1(a) and 4(b).
 Indeed, we show in Theorem C.4 (a more complete version of Theorem 5.2) that ARC effectively increases the $\mathsf{BP}$ of ($f,\kappa$)-robust methods from $\frac{1}{2+B^{2}}$ to $\frac{1}{2}$, provided the initial honest gradients are bounded in norm.
 Lastly, we would like to recall that when the workers’ gradients at model initialization are arbitrarily large, the convergence guarantee of Robust-DGD with ARC reduces to that of classic Robust-DGD (see Section 3).
-In other words, there is no downside to incorporating ARC.
+Practical scope of Theorem 5.2.
+Since $\xi=\frac{\mathsf{BP}-f/n}{\mathsf{BP}}$ and $\upsilon\geq\Psi(G,B,\rho)\xi$, it follows that the improvement factor $\upsilon$ is at least $\Psi(G,B,\rho)\frac{\mathsf{BP}-f/n}{\mathsf{BP}}$.
+For ARC to induce an improvement, we require $\upsilon<1$, which reduces to having $f>n\mathsf{BP}\left(1-\frac{1}{\Psi(G,B,\rho)}\right)$.
+Since Theorem 5.2 assumes $f<n\mathsf{BP}$, the improvement occurs when $f\in\left(n\mathsf{BP}\left(1-\frac{1}{\Psi(G,B,\rho)}\right),n\mathsf{BP}\right)$.
+The length of this interval, i.e., $n\frac{\mathsf{BP}}{\Psi(G,B,\rho)}$, depends on the size of the system $n$.
+In small systems where $n<\Psi(G,B,\rho)$, the length of the interval is smaller than 1 since $\mathsf{BP}<\nicefrac{{1}}{{2}}$.
+Therefore, at most one value of $f$, i.e., the largest integer smaller than $n\mathsf{BP}$, can lie in the improvement interval.
+Conversely, in large-scale systems where $n\in\Omega\left(\Psi(G,B,\rho)\log{\Psi(G,B,\rho)}\right)$, the improvement interval expands significantly and can be much larger than 1, thereby demonstrating ARC’s effectiveness in large-scale distributed learning.
 5.2 Influence of model initialization on empirical robustness
-Figure 6: Worst-case maximal accuracies achieved by Robust-DSGD with CWTM, both with and without ARC, on MNIST with $\alpha=0.1$, 10 honest worker, and $f=1$ Byzantine worker.
-The x-axis represents settings with progressively worse model initialization.
-This figure complements Figure 2b.
-Figures 2b and 6 illustrate the effect of model initialization on the performance of Robust-DSGD with and without ARC, evaluated on MNIST with 10 honest workers.
+Figure 5: Worst-case maximal accuracies achieved by Robust-DSGD with CWTM, on MNIST ($\alpha=0.1$), with 10 honest workers and 1 Byzantine worker. The x-axis represents worsening model initialization.
+Figures 1(b) and 5 illustrate the effect of model initialization on the performance of Robust-DSGD with and without ARC, evaluated on MNIST with 10 honest workers.
 We investigate two regimes of heterogeneity: extreme heterogeneity and $\alpha=0.1$, and consider $f=1$ adversarial worker.
@@
 The results show that, under well-initialized conditions ($\mu=1$), ARC significantly enhances the performance of Robust-DSGD, achieving a substantial improvement in worst-case maximal accuracy, particularly under extreme heterogeneity.
-In this regime, ARC boosts accuracy by about 70% compared to plain Robust-DSGD (Figure 2b).
+In this regime, ARC boosts accuracy by about 70% compared to plain Robust-DSGD (Figure 1(b)).
 This ($\mu=1$) corresponds to the initialization conditions of the empirical results presented in Section 4.
@@
 By $\mu=5$, the performance of Robust-DSGD with ARC converges to that of plain Robust-DSGD, both achieving an accuracy of around 10%.
-Another key observation from Figures 2b and 6 is that the behavior of Robust-DSGD with ARC closely mirrors that of Byzantine-free DSGD.
+Another key observation from Figures 1(b) and 5 is that the behavior of Robust-DSGD with ARC closely mirrors that of Byzantine-free DSGD.
 Both exhibit similarly poor performance when $\mu=5$, and their accuracies are comparable when $\mu\leq 3$, especially when $\alpha=0.1$.
 This suggests that ARC is particularly effective at exploiting good model initialization, similar to the performance of DSGD in the absence of Byzantine workers.
-In contrast, plain Robust-DSGD struggles to fully leverage well-initialized models, as evidenced by its consistently lower accuracy (around 20%) in Figure 2b.
+In contrast, plain Robust-DSGD struggles to fully leverage well-initialized models, as evidenced by its consistently lower accuracy (around 20%) in Figure 1(b).
 These findings highlight the important influence of model initialization on the robustness of aggregation methods, and empirically validate our theoretical findings in Section 5.1.
@@
 It has been extensively analyzed in the centralized setting (Zhang et al., 2020b; a; Koloskova et al., 2023), with applications to differential privacy (Pichapati et al., 2019).
-Moreover, we would like to note that gradient clipping has also been shown useful to obtain tighter convergence guarantees for stochastic gradient-descent methods in the case where the gradients have heavy-tailed noise Gorbunov et al.
-(2020; 2022); Danilova (2023).
+Moreover, we would like to note that gradient clipping has also been shown useful to obtain tighter convergence guarantees for stochastic gradient-descent methods in the case where the gradients have heavy-tailed noise (Gorbunov et al., 2020; 2022; Danilova, 2023).
 The considered proof techniques, however, do not apply directly to our adversarial distributed learning setting.
@@
 Other prior work (Mhamdi et al., 2021; Farhadkhani et al., 2022; Allouah et al., 2023a) that used static pre-aggregation clipping to enhance the empirical robustness of Robust-DGD, did not provide any formal explanation.
-While adaptive clipping schemes, similar to ARC, have been proposed and studied in the context of robust aggregation (Gupta & Vaidya, 2020; Liu et al., 2021; Diakonikolas et al., 2020; Abdalla & Zhivotovskiy, 2024), critical differences should be noted.
-The robustness guarantees in (Gupta & Vaidya, 2020; Liu et al., 2021) only apply to strongly convex loss functions, under a specific redundancy condition (namely, $2f$-redundancy).
+While adaptive clipping schemes, similar to ARC, have been studied in the context of robust aggregation (Gupta & Vaidya, 2020; Liu et al., 2021; Diakonikolas et al., 2020; Abdalla & Zhivotovskiy, 2024), critical differences should be noted.
+The robustness guarantees in Gupta & Vaidya (2020); Liu et al.
+(2021) only apply to strongly convex loss functions, under the specific $2f$-redundancy condition.
 We consider non-convex loss functions, as well as a generic heterogeneous setting of $(G,B)$-gradient dissimilarity.
 In contrast to the clipping scheme proposed in Diakonikolas et al.
-(2020); Abdalla & Zhivotovskiy (2024), ARC only uses the $f/n$ parameter to tune the clipping threshold and does not rely on any additional a priori knowledge of the distribution of honest workers’ gradients.
+(2020); Abdalla & Zhivotovskiy (2024), ARC only uses the $f$ parameter to tune the clipping threshold and does not rely on any additional a priori knowledge of the distribution of honest workers’ gradients.
 Moreover, we prove a deterministic robustness property of ARC (see Theorem 3.2), whereas Diakonikolas et al.
-(2020); Abdalla & Zhivotovskiy (2024) only provide probability guarantees assuming the non-adversarial inputs to be i.i.d.
+(2020); Abdalla & Zhivotovskiy (2024) only provides probability guarantees assuming the non-adversarial inputs to be i.i.d.
 with the distribution satisfying certain special properties.
+Lastly, prior work (Guerraoui et al., 2021; Zhu & Ling, 2022; Allouah et al., 2023c; Choffrut et al., 2024; Allouah et al., 2024) has considered the challenge of privacy preservation alongside robustness in distributed learning.
+In this context, the gradients are assumed to have bounded norms in order to control the sensitivity of the algorithm, as is common in differential privacy (DP) (Abadi et al., 2016).
+This is often enforced in practice through static gradient clipping.
+Our findings suggest that an adaptive approach, such as ARC, may provide a more reliable alternative.
+However, a rigorous analysis of ARC in the context of DP remains an important avenue for future research.
 7 Conclusion & Discussion
@@
 It encourages the development of robust distributed learning methods, like ARC, that can take full advantage of favorable practical conditions, thereby yielding relevant theoretical guarantees and superior empirical performance.
-Appendix A Computational Complexity of ARC.
-The algorithm starts by computing the norms of all input vectors, which takes $\mathcal{O}(nd)$ time.
-Note that this step is also required for static clipping.
-The sorting operation of the norms takes $\mathcal{O}(n\log(n))$ time.
-Note that if $n$ is large compared to ${k=\left\lfloor{2(f/n)(n-f)}\right\rfloor}$, one can use more efficient algorithms for finding the ${(k+1)}$-th largest element, such as quick-select Hoare (1961) which has an average-case complexity of $\mathcal{O}(n)$.
-Finally, clipping the vectors requires $\mathcal{O}(nd)$ time.
-Therefore, the total complexity of ARC is $\mathcal{O}(nd+n\log(n))$.
-Overall, the complexity is comparable to static clipping, given that the sorting (or search) is crucially dimension independent.
+Acknowledgment.
+This work was supported in part by the FNS Project *Controlling the Spread of Epidemics: A Computing Perspective* (200021_200477) and the FNS Project *TruBrain* (20CH21_218778 / 1).
+We also extend our gratitude to the anonymous reviewers of the ICLR 2025 conference for their valuable insights and feedback.
+Appendix A Additional Details on ARC
+This section provides further insights into the design and computational complexity of ARC, explaining its adaptive clipping strategy and its theoretical foundations.
+We also analyze its efficiency compared to static clipping.
+A.1 Design of ARC
+Lemma B.1 in Appendix B shows that $\mathbf{F}\circ\mathbf{Clip}_{C}$ is $(f,\tilde{\kappa})$-robust, provided that $\left\lvert{S\setminus S_{c}}\right\rvert\geq 1$ for all subsets $S$ of size $n-f$.
+Note that this condition is impossible to guarantee when using a fixed clipping threshold that does not depend on the input vectors.
+In order to ensure that $\left\lvert{S\setminus S_{c}}\right\rvert\geq 1$ for all subsets $S$ of size $n-f$, the clipping threshold $C$ should be large enough such that less than $n-f$ input vectors are clipped.
+This motivates the design of an adaptive clipping strategy.
+Accordingly, we propose to choose a clipping threshold such that the total number of clipped vectors is of the form $\left\lfloor{\lambda(n-f)}\right\rfloor$, where $\lambda<1$.
+Note that it is natural to clip more vectors as the fraction of adversarial workers $\nicefrac{{f}}{{n}}$ increases, to control the impact of Byzantine behavior.
+Therefore, we set $\lambda\coloneqq\zeta\frac{f}{n}$ where $0\leq\zeta\leq 2$.
+Since $\frac{f}{n}<\frac{1}{2}$ and $\lambda<1$, the total number of clipped vectors $\left\lfloor{\zeta\frac{f}{n}(n-f)}\right\rfloor<n-f$ for all $\zeta\in[0,2]$.
+This constitutes the underlying idea behind our adaptive clipping strategy ARC, presented in Algorithm 2.
+Notably, our theoretical results hold for all $\zeta\in[0,2]$.
+However, empirical evaluations show that setting $\zeta=2$ consistently provides the best performance, leading us to adopt this value in Algorithm 2 and throughout the paper.
+A.2 Computational Complexity of ARC
+The computational complexity of ARC is comparable to static clipping, with the primary difference being an additional sorting step. The overall complexity can be broken down as follows:
+  1. 1.
+    Norm computation: Computing the norms of all input vectors requires $\mathcal{O}(nd)$ time. (This step is also required in static clipping.)
+  2. 2.
+    Sorting step: Sorting the norms incurs an additional $\mathcal{O}(n\log n)$ time.
+If $n$ is large relative to $k=\left\lfloor{2(f/n)(n-f)}\right\rfloor$, an efficient selection algorithm such as quick-select (Hoare, 1961) can be used instead, reducing the complexity to $\mathcal{O}(n)$ in the average case.
+  3. 3.
+    Gradient clipping: Clipping vectors requires $\mathcal{O}(nd)$ time.
+Thus, the total time complexity of ARC is:
+$\mathcal{O}(nd+n\log n).$
+Crucially, this complexity remains dimension-independent in the sorting step, making it scalable while maintaining robustness.
 Appendix B Proofs for Results in Section 3
@@
 Consider first the case $\kappa=0$.
-Since $\kappa\geq\frac{f}{n-2f}$  [^5] Proposition 6, Allouah et al.
-(2023a), we have that $f=0$.
+Since $\kappa\geq\frac{f}{n-2f}$  [^2] Proposition 6 (Allouah et al., 2023a), we have that $f=0$.
 Recall that we denote $y_{i}\coloneq\mathrm{clip}_{C}(x_{i})$ and $\overline{y}_{S}\coloneqq\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}y_{i}$.
@@
 Hence we know that at least $f+1$ vectors have their norm greater or equal to the clipping threshold $C=\left\lVert{x_{k+1}}\right\rVert$,
-$\underbrace{\left\lVert{x_{1}}\right\rVert\leq\dots\leq\left\lVert{x_{k+1}}\right\rVert}_{\text{at least $f+1$ vectors}}\leq\underbrace{\left\lVert{x_{k+2}}\right\rVert\leq\dots\leq\left\lVert{x_{n}}\right\rVert}_{\text{at most $n-f-1$ vectors}}$
+$\underbrace{\left\lVert{x_{1}}\right\rVert\geq\dots\geq\left\lVert{x_{k+1}}\right\rVert}_{\text{at least $f+1$ vectors}}\geq\underbrace{\left\lVert{x_{k+2}}\right\rVert\geq\dots\geq\left\lVert{x_{n}}\right\rVert}_{\text{at most $n-f-1$ vectors}}$
 Hence, at most $n-f-1$ vectors will not be clipped.
@@
 D.1 Datasets and Heterogeneity
-In our experiments, we consider three standard image classification datasets, namely MNIST Deng (2012), Fashion-MNIST Xiao et al.
-(2017), and CIFAR-10 Krizhevsky et al.
-(2014).
-To simulate data heterogeneity in our experiments, we make the honest workers sample from the datasets using a Dirichlet Hsu et al.
-(2019a) distribution of parameter $\alpha$, as done in Hsu et al.
-(2019b); Allouah et al.
-(2023a); Farhadkhani et al.
-(2023).
+In our experiments, we consider three standard image classification datasets, namely MNIST (Deng, 2012), Fashion-MNIST (Xiao et al., 2017), and CIFAR-10 (Krizhevsky et al., 2014).
+To simulate data heterogeneity in our experiments, we make the honest workers sample from the datasets using a Dirichlet distribution of parameter $\alpha$ (Hsu et al., 2019a), as done in (Hsu et al., 2019b; Allouah et al., 2023a; Farhadkhani et al., 2023).
 The smaller the $\alpha$, the more heterogeneous the setting.
@@
 Moreover, CIFAR-10 is expanded with horizontally flipped images, followed by a per channel normalization with means 0.4914, 0.4822, 0.4465 and standard deviations 0.2023, 0.1994, 0.2010.
-(a) $\alpha=0.1$
-(b) $\alpha=0.5$
-(c) $\alpha=1$
-(d) $\alpha=0.075$
-(e) $\alpha=0.2$
 Figure 6: Distribution of labels across honest workers on MNIST (row 1) and CIFAR-10 (row 2).
 D.2 Algorithm, Distributed System, ML Models, and Hyperparameters
-We perform our experiments using Robust-DSGD (see Algorithm 3), which is known to be an order-optimal robust variant of Robust-DGD Allouah et al.
-(2023a).
-On MNIST and Fashion-MNIST, we execute Robust-DSGD in a distributed system composed of $n-f=10$ honest workers, and $f\in\{1,2,3,4,5,7,9\}$ adversarial workers.
-Furthermore, we train a convolutional neural network (CNN) of 431,080 parameters with batch size $b=25$, $T=1000$, $\gamma=0.1$, and momentum parameter $\beta=0.9$.
+We perform our experiments using Robust-DSGD, an order-optimal robust variant of Robust-DGD (Allouah et al., 2023a) (see Algorithm 3), in distributed systems of varying sizes.
+On MNIST and Fashion-MNIST, we train a convolutional neural network (CNN) of 431,080 parameters with batch size $b=25$, $T=1000$, $\gamma=0.1$, and momentum parameter $\beta=0.9$.
 Moreover, the negative log likelihood (NLL) loss function is used, along with an $\ell_{2}$-regularization of $10^{-4}$.
-On CIFAR-10, we execute Robust-DSGD using a CNN of 1,310,922 parameters, in a distributed system comprising $n-f=16$ honest workers and $f=1$ adversarial worker.
+On CIFAR-10, we train a CNN of 1,310,922 parameters.
 We set $b=50$, $T=2000$, $\beta=0.9$, and $\gamma=0.05$ decaying once at step 1500.
 Finally, we use the NLL loss function with an $\ell_{2}$ regularization of $10^{-2}$.
+In order to present the architectures of the ML models used in our experiments, we adopt the following compact terminology introduced in Choffrut et al.
+(2024).
+L(#outputs) represents a fully-connected linear layer, R stands for ReLU activation, S stands for log-softmax, C(#channels) represents a fully-connected 2D-convolutional layer (kernel size 5, padding 0, stride 1), M stands for 2D-maxpool (kernel size 2), B stands for batch-normalization, and D represents dropout (with fixed probability 0.25).
+The comprehensive experimental setup and the architecture of the models are presented in Table 2.
 Algorithm 3 Robust Distributed Stochastic Gradient Descent (Robust-DSGD)
@@
   Output: Server outputs $\hat{{{\theta}}}$ chosen uniformly at random from $\{{{\theta}_{1}},\ldots,\,{{\theta}_{T}}\}$.
-In order to present the architectures of the ML models used in our experiments, we adopt the following compact terminology introduced in Allouah et al. (2023a).
-L(#outputs) represents a fully-connected linear layer, R stands for ReLU activation, S stands for log-softmax, C(#channels) represents a fully-connected 2D-convolutional layer (kernel size 5, padding 0, stride 1), M stands for 2D-maxpool (kernel size 2), B stands for batch-normalization, and D represents dropout (with fixed probability 0.25).
-The comprehensive experimental setup, as well as the architecture of the models, are presented in Table 2.
 Dataset (Fashion-)MNIST CIFAR-10
@@
 Batch size $b=25$ $b=50$
-Honest workers $n-f=10$ $n-f=16$
-Adversarial workers $f\in\{1,2,3,4,5,7,9\}$ $f=1$
-Table 2: Experimental setup on (Fashion-)MNIST and CIFAR-10
+Table 2: Comprehensive experimental setup.
 D.3 Byzantine Attacks
-In our experiments, the adversarial workers execute five state-of-the-art adversarial attacks from the Byzantine ML literature, namely sign-flipping (SF) Allen-Zhu et al.
-(2020), label-flipping (LF) Allen-Zhu et al.
-(2020), mimic Karimireddy et al.
-(2022), fall of empires (FOE) Xie et al.
-(2020), and a little is enough (ALIE) Baruch et al.
-(2019).
-The exact functionality of the attacks is detailed below.
+In our experiments, the adversarial workers execute five state-of-the-art adversarial attacks from the Byzantine ML literature, namely sign-flipping (SF) (Allen-Zhu et al., 2020), label-flipping (LF) (Allen-Zhu et al., 2020), mimic (Karimireddy et al., 2022), fall of empires (FOE) (Xie et al., 2020), and a little is enough (ALIE) (Baruch et al., 2019).
+The exact functionality of the attacks is detailed next.
 In every step $t$, let $\overline{m}_{t}$ be an estimation of the true honest momentum at step $t$.
-In our experiments, we estimate $\overline{m}_{t}$ by averaging the momentums sent by the honest workers in step $t$ of Robust-DSHB.
-In other words, $\overline{m}_{t}=\frac{1}{|\mathcal{H}|}\sum\limits_{i\in\mathcal{H}}m_{t}^{(i)}$, where $m_{t}^{(i)}$ is the momentum computed by honest worker $w_{i}$ in step $t$.
+In our experiments, we estimate $\overline{m}_{t}$ by averaging the momentums sent by the honest workers in step $t$ of Robust-DSGD.
+In other words, $\overline{m}_{t}=\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}m_{t}^{(i)}$, where $m_{t}^{(i)}$ is the momentum computed by honest worker $w_{i}$ in step $t$.
   • SF: the adversarial workers send the vector $-\overline{m}_{t}$ to the server.
@@
   • ALIE: the adversarial workers send $\overline{m}_{t}+\tau\sigma_{t}$ in step $t$ to the server, where $\tau\geq 0$ is a fixed real number representing the attack factor, and $\sigma_{t}$ is the coordinate-wise standard deviation of $\overline{m}_{t}$.
-Since the FOE and ALIE attacks have $\tau$ as parameter, we implement in our experiments enhanced and adaptive versions of these attacks, where the attack factor is not constant and may be different in every iteration.
+Since FOE and ALIE are parametrized by $\tau\in\mathbb{R}$, we implement in our experiments enhanced and adaptive versions of these attacks, where the attack factor is not constant and may be different in every iteration.
 In every step $t$, we determine the optimal attack factor $\tau_{t}$ through a grid search over a predefined range of values.
-More specifically, in every step $t$, $\tau_{t}$ takes the value that maximizes the damage inflicted by the adversarial workers, i.e., that maximizes the $l^{2}$ norm of the difference between the average of the honest momentums $\overline{m}_{t}$ and the output of the aggregation $R_{t}$ at the server.
+More specifically, in every step $t$, $\tau_{t}$ takes the value that maximizes the damage inflicted by the adversarial workers, i.e., that maximizes the $\ell_{2}$-norm of the difference between the average of the honest momentums $\overline{m}_{t}$ and the output of the aggregation $R_{t}$ at the server.
 D.4 Benchmarking and Reproducibility
 […]
-Appendix E Additional Experimental Results
-In this section, we complete the experimental results that could not be placed in the main paper.
-E.1 MNIST
-(a) extreme
-(b) $\alpha=0.5$
-(c) $\alpha=0.1$
-(d) $\alpha=0.1$
-(e) $\alpha=1$
-(f) $\alpha=1$
-Figure 8: Worst-case maximal accuracies achieved by Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed MNIST with $10$ honest workers.
-We fix the heterogeneity level, and vary the the number of Byzantine workers $f$.
-(a) $f=1$
-(b) $f=3$
-(c) $f=3$
-(d) $f=5$
-(e) $f=5$
-Figure 9: Worst-case maximal accuracies achieved by Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed MNIST with $10$ honest workers.
-We fix the number of Byzantine workers $f$, and vary the heterogeneity level.
-(a) ALIE
-(b) SF
-(c) LF
-(d) Mimic
-Figure 10: Performance of Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed MNIST (extreme heterogeneity) with $10$ honest workers and $f=1$, under several attacks.
-This complements Figure 3c of the main paper.
-(a) FOE
-(b) SF
-(c) ALIE
-(d) LF
-(e) Mimic
-Figure 11: Performance of Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed MNIST ($\alpha=0.1$) with $10$ honest workers and $f=3$, under several attacks.
-(a) $f=3$
-(b) $f=4$
-(c) $f=5$
-Figure 12: Performance of Robust-DSGD under the LF attack when using ARC compared to no clipping, on heterogeneously-distributed MNIST ($\alpha=0.1$) with $10$ honest workers and varying number of adversarial workers $f$.
-E.2 Fashion-MNIST
-(a) $\alpha=0.1$
-(b) $\alpha=0.1$
-(c) $\alpha=0.05$
-(d) $\alpha=0.05$
-(e) extreme
-(f) extreme
-Figure 13: Worst-case maximal accuracies achieved by Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed Fashion-MNIST with $10$ honest workers.
-We fix the heterogeneity level and vary the number of adversarial workers $f$.
-(a) $f=1$
-(b) $f=1$
-(c) $f=3$
-(d) $f=3$
-(e) $f=4$
-(f) $f=4$
-Figure 14: Worst-case maximal accuracies achieved by Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed Fashion-MNIST with $10$ honest workers.
-We fix the number of adversarial workers $f$ and vary the heterogeneity level.
-E.3 CIFAR-10
-(a) $f=1$
-(b) $f=1$
-Figure 15: Worst-case maximal accuracies achieved by Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed CIFAR-10 with $16$ honest workers.
-We fix the number of adversarial workers $f=1$ and vary the heterogeneity level.
-(a) ALIE with CWTM and GM
-(b) FOE with CWMED and MK
-(c) ALIE with CWMED and CWMED
-Figure 16: Worst-case maximal accuracies achieved by Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed CIFAR-10 ($\alpha=0.075$) with $16$ honest workers and $f=1$ executing ALIE and FOE.
-This figure complements Figure 5 in the main paper.
-Appendix F Static Clipping vs ARC
+Appendix E ARC vs. Static Clipping
 In this section, we present empirical results on static clipping when used as a pre-aggregation technique in Robust-DSGD, and compare its performance to our proposed adaptive algorithm ARC.
-F.1 Limitations of Static Clipping
-We perform our experiments using Robust-DSGD (see Algorithm 3) on (Fashion-)MNIST and CIFAR-10.
-We consider three different levels of heterogeneity, that we call moderate (corresponding to sampling from a Dirichlet distribution of parameter $\alpha=1$), high ($\alpha=0,1$), and extreme (as explained in Appendix D).
-On MNIST and Fashion-MNIST, we execute Robust-DSGD in a distributed system composed of $n=15$ workers, among which $f\in\{3,4\}$ are Byzantine.
-Furthermore, we train a convolutional neural network of 431,080 parameters with batch size $b=25$, $T=1000$, $\gamma=0.1$, and momentum parameter $\beta=0.9$.
-Moreover, the negative log likelihood loss function is used, along with an $\ell_{2}$-regularization of $10^{-4}$.
-On CIFAR-10, we execute Robust-DSGD on ResNet-18 He et al.
-(2015), in a distributed system comprising $n=9$ workers among which $f\in\{1,2\}$ are Byzantine.
-We set $b=128$, $T=2000$, $\beta=0.9$, and $\gamma=0.1$ decaying once by 10$\times$ at step 1500.
-Finally, we use the cross-entropy loss function with an $\ell_{2}$ regularization of $5\times 10^{-4}$.
-In our experiments, we examine a wide range of static clipping parameters. Specifically, we choose $C\in\{0.02,0.2,2,20\}$ on MNIST.
-(a) $f=3$
-(b) $f=4$
-Figure 17: Impact of the clipping strategy and heterogeneity level on the worst-case maximal accuracy achieved by Robust-DSGD against five Byzantine attacks on heterogeneous MNIST with $n=15$ workers.
-CWTM $\circ$ NNM is used as aggregation.
-DSGD reaches at least 98.5% in accuracy in all heterogeneity regimes.
+E.1 Comparison with Static Clipping on MNIST
+On MNIST, we execute Robust-DSGD in a distributed system composed of $n=15$ workers, among which $f\in\{3,4\}$ are adversarial.
+We consider three different levels of heterogeneity: moderate ($\alpha=1$), high ($\alpha=0,1$), and extreme (as explained in Appendix D).
+In our experiments, we examine a wide range of static clipping parameters $C\in\{0.02,0.2,2,20\}$.
+(a) Sign-flipping (SF) (Allen-Zhu et al., 2020)
+(b) Label-flipping (LF) (Allen-Zhu et al., 2020)
+Figure 7: Impact of pre-aggregation clipping, specifically static clipping with $C=2$ and our adaptive clipping algorithm ARC, on Robust-DGD with CWTM $\circ$ NNM, under the SF and LF attacks.
+We consider the MNIST (Deng, 2012) dataset distributed amongst $10$ honest (non-adversarial) workers with extreme heterogeneity, and there are three adversarial workers.
 Unreliability due to the attack.
 First, the efficacy of static clipping is significantly influenced by the type of Byzantine attack executed during the learning process.
-To illustrate, while $C=2$ exhibits the best performance under the SF attack Allen-Zhu et al.
-(2020), depicted in Figure 1, it leads to a complete collapse of learning under LF Allen-Zhu et al.
-(2020).
+To illustrate, while $C=2$ exhibits the best performance under the SF attack (Allen-Zhu et al., 2020), depicted in Figure 7, it leads to a complete collapse of learning under LF (Allen-Zhu et al., 2020).
 Consequently, identifying a static clipping threshold $C$ that consistently performs well in practice against any attack is difficult (if not impossible).
@@
 To illustrate this impact, we present in Figure 2 the performances of Robust-DSGD for different levels of heterogeneity (moderate, high, and extreme) and different values of $C$.
-In the left plot of Figure 17, $C=2$ emerges as the best static clipping threshold when the heterogeneity is high whereas $C=20$ appears to be sub-optimal.
+In Figure 2(a), $C=2$ emerges as the best static clipping threshold when the heterogeneity is high whereas $C=20$ appears to be sub-optimal.
 On the other hand, under extreme heterogeneity, the accuracy associated to $C=2$ diminishes drastically while $C=20$ becomes a better choice.
-Intuitively, in heterogeneous scenarios, honest gradients become large in $l^{2}$-norm, due to the increasingly detrimental effect of Byzantine attacks.
+Intuitively, in heterogeneous scenarios, honest gradients become large in $\ell_{2}$-norm, due to the increasingly detrimental effect of Byzantine attacks.
 Consequently, we must increase the static clipping threshold.
-Failing to do so could introduce a significant bias.
-More details on this observation can be found in Figure 18.
 This highlights the intricate dependence of the performance of static clipping on data heterogeneity, and emphasizes the necessity to fine-tune static clipping strategies prior to the learning.
-Figure 18: Distribution of the norms of honest gradients during the learning on heterogeneous MNIST.
-There are $f=3$ Byzantine workers (among $n=15$) executing the LF attack.
-CWTM $\circ$ NNM is used as aggregation method.
-Unreliability due to the number of Byzantine workers.
-Last but not least, the number of Byzantine workers $f$ also affects the efficacy of static clipping.
-As seen in Figure 17, $C=2$ leads to the highest accuracy among static clipping strategies when $f=3$ in high heterogeneity, but cannot be used when $f=4$ as its corresponding accuracy drops below 50% (see right plot of Figure 17).
+Unreliability due to the number of adversarial workers.
+Last but not least, the number of adversarial workers $f$ also affects the efficacy of static clipping.
+As seen in Figure 2, $C=2$ leads to the highest accuracy among static clipping strategies when $f=3$ in high heterogeneity, but cannot be used when $f=4$ as its corresponding accuracy drops below 50% (see Figure 2(b)).
 Overall, our empirical findings reveal that there is no single value of $C$ for which static clipping consistently delivers satisfactory performance across the various settings.
@@
 First, since the server does not have direct access to the data, it cannot estimate the degree of heterogeneity.
-Second, as (by definition) a Byzantine worker can behave in an unpredictable manner, we cannot adjust $C$ to the attack(s) being executed by the Byzantine workers.
+Second, as (by definition) an adversarial worker can behave in an unpredictable manner, we cannot adjust $C$ to the attack(s) being executed by the adversarial workers.
 This highlights the necessity for a robust clipping alternative that can naturally adapt to the setting in which it is deployed.
-F.2 Empirical Benefits of ARC
-In contrast to static clipping, ARC is adaptive, delivering consistent performance across diverse levels of heterogeneity, Byzantine attacks, and number of Byzantine workers.
+In contrast to static clipping, ARC is adaptive, delivering consistent performance across diverse levels of heterogeneity, Byzantine attacks, and number of adversarial workers.
 Adaptiveness.
@@
 Specifically, in scenarios of moderate heterogeneity where clipping may not be essential, ARC matches the performance of No clipping as well as static strategies $C=2$ and $C=20$.
-Conversely, the left plot of Figure 17 shows that under high and extreme heterogeneity, ARC surpasses the best static clipping strategy in terms of accuracy, highlighting the effectiveness of our approach.
+Conversely, Figure 2(a) shows that under high and extreme heterogeneity, ARC surpasses the best static clipping strategy in terms of accuracy, highlighting the effectiveness of our approach.
 Robust performance across Byzantine regimes.
-ARC exhibits robust performance across diverse Byzantine scenarios, encompassing variations in both the type of Byzantine attack and the number of Byzantine workers $f$.
-As depicted in Figures 1 and 17, ARC consistently yields robust performance, regardless of the Byzantine attack.
-In extreme heterogeneity with $f=3$, ARC maintains an accuracy of 75%, while $C=2$ demonstrates a subpar worst-case accuracy of 25% (also refer to the right plot of Figure 1).
-Furthermore, despite the increase in the number of Byzantine workers to $f=4$, ARC remains the top-performing clipping approach among all considered strategies.
+ARC exhibits robust performance across diverse Byzantine scenarios, encompassing variations in both the type of Byzantine attack and the number of adversarial workers $f$.
+As depicted in Figures 7 and 2, ARC consistently yields robust performance, regardless of the Byzantine attack.
+In extreme heterogeneity with $f=3$, ARC maintains an accuracy of 75%, while $C=2$ demonstrates a subpar worst-case accuracy of 25% (also refer to Figure 7(b)).
+Furthermore, despite the increase in the number of adversarial workers to $f=4$, ARC remains the top-performing clipping approach among all considered strategies.
 This analysis underscores the empirical superiority of ARC over static clipping methods, eliminating the reliance on data heterogeneity and the specific Byzantine regime.
@@
 Figure 8: Evolution of the adaptive clipping parameter $C_{t}$ of ARC compared to the norm of the honest mean during the learning on heterogeneous MNIST ($\alpha=1$).
-There are $f=3$ Byzantine workers among $n=15$.
+There are $f=3$ adversarial workers among $n=15$.
 CWTM $\circ$ NNM is used as aggregation.
-ARC dynamically adjusts its clipping parameter $C_{t}$ based on the norms of honest momentums at step $t$, avoiding static over-clipping or under-clipping.
-This adaptability is evident in under all attacks, where $C_{t}$ consistently decreases with time, an expected behavior when reaching convergence.
-(a) CWMed $\circ$ NNM
-(b) GM $\circ$ NNM
-(c) MK $\circ$ NNM
 Figure 9: Impact of the clipping strategy on the worst-case maximal accuracy achieved during the learning against five Byzantine attacks on heterogeneous MNIST, with $f=3$ and $n=15$.
+E.2 Comparison with Static Clipping on Fashion-MNIST
+On Fashion-MNIST, we execute Robust-DSGD in a distributed system composed of $n=15$ workers, among which $f=3$ are adversarial. See Figure 10.
+Figure 10: Impact of the clipping strategy and heterogeneity level on the worst-case maximal accuracy achieved during the learning against five Byzantine attacks on Fashion-MNIST.
+$f=3$ and $n=15$, and the heterogeneity levels considered are extreme, high ($\alpha=0.1$), and moderate ($\alpha=1)$.
+Aggregation methods used: CWTM $\circ$ NNM (top left), CWMed $\circ$ NNM (top right), GM $\circ$ NNM (bottom left), and Multi-Krum $\circ$ NNM (bottom right).
+DSGD reaches 85% in accuracy in all settings.
+E.3 Comparison with Static Clipping on CIFAR-10
+For this particular experiment, we execute Robust-DSGD using the ResNet-18 (He et al., 2015) model, in a distributed system comprising $n=9$ workers among which $f\in\{1,2\}$ are adversarial.
+We set $b=128$, $T=2000$, $\beta=0.9$, and $\gamma=0.1$ decaying once by 10$\times$ at step 1500.
+Finally, we use the cross-entropy loss function with an $\ell_{2}$-regularization of $5\times 10^{-4}$.
+See Tables 3 and 4.
+Attack $C=0.05$ $C=0.5$ $C=5$ $C=50$ No clipping ARC
+FOE 26.4 $\pm$ 1.2 52.6 $\pm$ 2.0 76.4 $\pm$ 5.4 47.0 $\pm$ 8.0 43.0 $\pm$ 6.2 $79.1\pm 1.6$
+ALIE 34.6 $\pm$ 1.3 43.3 $\pm$ 1.1 47.2 $\pm$ 3.4 47.2 $\pm$ 4.2 47.2 $\pm$ 4.2 $51.8\pm 3.5$
+LF 33.7 $\pm$ 1.5 58.4 $\pm$ 1.8 83.6 $\pm$ 1.2 $85.7\pm 0.5$ $85.9\pm 0.4$ 81.5 $\pm$ 1.4
+SF 32.3 $\pm$ 1.4 57.9 $\pm$ 1.2 77.2 $\pm$ 4.6 68.6 $\pm$ 8.5 66.4 $\pm$ 9.0 $82.0\pm 1.1$
+Mimic 38.2 $\pm$ 1.6 68.3 $\pm$ 1.4 $85.6\pm 0.42$ $85.7\pm 0.4$ $85.7\pm 0.4$ $85.4\pm 0.5$
+Worst-case 26.4 $\pm$ 1.2 43.3 $\pm$ 1.1 47.2 $\pm$ 3.4 47.0 $\pm$ 8.0 43.0 $\pm$ 6.2 $51.8\pm 3.5$
+Table 3: Maximum accuracy (%) achieved by CWTM $\circ$ NNM on CIFAR-10 under moderate heterogeneity ($\alpha=1$), for various clipping strategies and attacks.
+There are $f=2$ adversarial workers among $n=9$.
+We highlight in blue the highest accuracy achieved per attack (i.e., per row).
+Attack $C=0.05$ $C=0.5$ $C=5$ $C=50$ No clipping ARC
+FOE 28.9 $\pm$ 1.7 56.8 $\pm$ 1.9 $75.6\pm 2.8$ 35.5 $\pm$ 11.6 34.9 $\pm$ 12.2 $75.9\pm 3.0$
+ALIE 34.9 $\pm$ 1.3 43.4 $\pm$ 1.0 47.4 $\pm$ 4.1 44.9 $\pm$ 3.8 44.9 $\pm$ 3.8 $54.6\pm 3.6$
+LF 33.7 $\pm$ 1.4 58.4 $\pm$ 1.8 77.4 $\pm$ 12.9 $85.8\pm 0.6$ $85.6\pm 0.7$ 80.5 $\pm$ 1.8
+SF 32.5 $\pm$ 1.4 57.5 $\pm$ 1.4 77.7 $\pm$ 5.0 63.0 $\pm$ 12.0 55.4 $\pm$ 18.9 $82.5\pm 0.3$
+Mimic 38.1 $\pm$ 1.3 68.5 $\pm$ 1.4 $85.5\pm 0.5$ $85.9\pm 0.5$ $85.9\pm 0.5$ $85.4\pm 0.2$
+Worst-case 28.9 $\pm$ 1.7 43.4 $\pm$ 1.0 47.4 $\pm$ 4.1 35.5 $\pm$ 11.6 34.9 $\pm$ 12.2 $54.6\pm 3.6$
+Table 4: Maximum accuracy (%) achieved by CWMed $\circ$ NNM on heterogeneous CIFAR-10 ($\alpha=1$), for various clipping strategies and attacks.
+There are $f=2$ adversarial workers among $n=9$.
+We highlight in blue the highest accuracy achieved per attack (i.e., per row).
+Appendix F Performance Gains of ARC Over Robust-DSGD
+In this section, we present additional experimental results demonstrating the performance improvements achieved by ARC in Robust-DSGD on the MNIST, Fashion-MNIST, and CIFAR-10 datasets.
+These results could not be included in the main paper due to space constraints.
+F.1 MNIST
+F.1.1 Initial System (Main Paper): 10 Honest Workers
+We consider training on MNIST in a system comprised of $n-f=10$ honest workers. These results complement the ones presented in Section 4 of the main paper.
+Figure 11: Worst-case maximal accuracies achieved by Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed MNIST with $10$ honest workers.
+We fix the heterogeneity level, and vary the the number of adversarial workers $f$.
+Figure 12: Worst-case maximal accuracies achieved by Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed MNIST with $10$ honest workers.
+We fix the number of adversarial workers $f$, and vary the heterogeneity level.
+Figure 13: Performance of Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed MNIST (extreme heterogeneity) with $10$ honest workers and $f=1$, under several attacks.
+This complements Figure 3(c) of the main paper.
+Figure 14: Performance of Robust-DSGD when using ARC compared to no clipping, on heterogeneous MNIST ($\alpha=0.1$) with $10$ honest workers and $f=3$, under several attacks.
+Figure 15: Performance of Robust-DSGD under the LF attack when using ARC compared to no clipping, on heterogeneously-distributed MNIST ($\alpha=0.1$) with $10$ honest workers and varying number of adversarial workers $f$.
+F.1.2 Larger System: 30 Honest Workers
+We consider training on the MNIST dataset in a larger system comprised of $n-f=30$ honest workers, and $f\in\{3,6,9\}$ adversarial workers.
+Figure 16: Performance of Robust-DSGD when using ARC and without clipping on distributed MNIST under extreme heterogeneity. There are $f=3$ adversarial workers executing 4 attacks.
+Figure 17: Performance of Robust-DSGD when using ARC and without clipping on heterogeneously-distributed MNIST with $\alpha=0.1$. There are $f=6$ adversarial workers executing 4 attacks.
+Figure 18: Performance of Robust-DSGD when using ARC and without clipping on heterogeneously-distributed MNIST with $\alpha=0.1$. There are $f=9$ adversarial workers executing 4 attacks.
+Figure 19: Worst-case maximal accuracies achieved by Robust-DSGD, with and without ARC, on heterogeneously-distributed MNIST with $30$ honest workers.
+We consider a heterogeneous data distribution with $\alpha=0.1$ (left) and extreme heterogeneity (right), and vary $f\in\{3,6,9\}$.
+F.2 Fashion-MNIST
+Figure 20: Worst-case maximal accuracies achieved by Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed Fashion-MNIST with $10$ honest workers.
+We fix the heterogeneity level and vary the number of adversarial workers $f$.
+Figure 21: Worst-case maximal accuracies achieved by Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed Fashion-MNIST with $10$ honest workers.
+We fix the number of adversarial workers $f$ and vary the heterogeneity level.
+F.3 CIFAR-10
+Initial System (Main Paper): 16 Honest Workers
+$\alpha=0.2$ $\alpha=0.075$
+Aggregation No Clipping ARC No Clipping ARC
+CWMed 43.4 $\pm$ 4.2 69.4 $\pm$ 0.8 13.7 $\pm$ 12.1 62.7 $\pm$ 1.2
+MK 50.9 $\pm$ 5.1 68.7 $\pm$ 0.5 40.5 $\pm$ 0.7 59.9 $\pm$ 1.9
+Table 5: Worst-case maximal accuracies (%) achieved by Robust-DSGD on heterogeneously-distributed CIFAR-10 with ARC and without.
+There is $f=1$ adversarial worker among $n=17$.
+As a baseline, DSGD ($f=0$) reaches 76.5% and 70% when $\alpha=0.2$ and 0.075, respectively.
+This table complements Table 1 in the main paper.
+Figure 22: Worst-case maximal accuracies achieved by Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed CIFAR-10 with $16$ honest workers.
+We fix the number of adversarial workers $f=1$ and vary the heterogeneity level.
+Figure 23: Performance achieved by Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed CIFAR-10 ($\alpha=0.075$) with $16$ honest workers and $f=1$ executing ALIE and FOE.
+We also test the performance of ARC in more adversarial regimes, with $f\in\{2,3\}$ adversarial workers (along with $n-f=16$ honest workers).
+We consider heterogeneity regimes of $\alpha=0.2$ and 0.5, when $f=2$ and 3, respectively.
+The results are presented in Table 6, and we show in Figure 24 the performance of ARC compared to no clipping under the FOE attack.
+$f=2,\alpha=0.2$ $f=3,\alpha=0.5$
+Aggregation No Clipping ARC No Clipping ARC
+CWTM $\circ$ NNM 44.3 $\pm$ 5.2 53.0 $\pm$ 4.7 52.0 $\pm$ 1.3 55.4 $\pm$ 1.1
+GM $\circ$ NNM 33.8 $\pm$ 6.7 50.2 $\pm$ 3.3 50.9 $\pm$ 2.5 49.6 $\pm$ 1.4
+CWMed $\circ$ NNM 32.6 $\pm$ 9.8 49.7 $\pm$ 2.7 48.7 $\pm$ 2.4 62.7 $\pm$ 0.9
+MK $\circ$ NNM 45.0 $\pm$ 4.2 51.9 $\pm$ 2.1 50.3 $\pm$ 1.4 50.2 $\pm$ 2.4
+Table 6: Worst-case maximal accuracies (%) achieved by Robust-DSGD on heterogeneously-distributed CIFAR-10 with ARC and without.
+There are $f\in\{2,3\}$ adversarial workers with $n-f=16$ honest workers.
+Figure 24: Performance of Robust-DSGD when using ARC and without clipping on CIFAR-10.
+There are 16 honest workers, and $f=2$ (left) and $f=3$ (right) adversarial workers executing the FOE attack.
+The aggregations used are CWTM $\circ$ NNM and GM $\circ$ NNM (top), and CWMed $\circ$ NNM and MK $\circ$ NNM (bottom).
+Larger System: 33 Honest Workers
+We also run experiments on CIFAR-10 in a larger system comprised of $n-f=33$ honest workers and $f\in\{2,3\}$ adversarial workers.
+We consider the high heterogeneity regime of $\alpha=0.1$.
+The results are presented in Table 7, and we show in Figure 25 the performance of ARC compared to no clipping under the FOE attack.
+$f=2$ $f=3$
+Aggregation No Clipping ARC No Clipping ARC
+CWTM $\circ$ NNM 45.5 $\pm$ 1.3 54.2 $\pm$ 1.0 41.3 $\pm$ 0.5 47.5 $\pm$ 1.5
+GM $\circ$ NNM 41.8 $\pm$ 1.1 55.2 $\pm$ 2.2 40.0 $\pm$ 1.2 47.5 $\pm$ 1.5
+CWMed $\circ$ NNM 42.5 $\pm$ 1.42 54.7 $\pm$ 1.2 35.7 $\pm$ 8.7 49.7 $\pm$ 3.4
+MK $\circ$ NNM 45.8 $\pm$ 0.8 54.8 $\pm$ 2.5 41.4 $\pm$ 0.6 48.7 $\pm$ 2.8
+Table 7: Worst-case maximal accuracies (%) achieved by Robust-DSGD on heterogeneously-distributed CIFAR-10 with ARC and without.
+There are $f=2$ (left) and $f=3$ (right) adversarial workers, along with $n-f=33$ honest workers.
+We consider the high heterogeneity regime $\alpha=0.1$.
+Figure 25: Performance of Robust-DSGD when using ARC and without clipping on CIFAR-10.
+There are 33 honest workers, and $f=2$ (top) and 3 (bottom) adversarial workers executing the FOE attack.
+We consider the high heterogeneity regime $\alpha=0.1$.
+The aggregations used are CWTM $\circ$ NNM and GM $\circ$ NNM (left), and CWMed $\circ$ NNM and MK $\circ$ NNM (right).
+F.4 Runtime Comparison: NNM (Allouah et al., 2023a) vs. ARC
+(a) Varying Model Size $d$
+(b) Varying System Size $n$
+Figure 26: Computational performance of ARC vs NNM in two different settings.
+In the left plot, we fix the total number of workers $n=30$, the number of adversarial workers $f=3$, and vary the model size $d$.
+In the right plot, we fix the model size $d=1310922$ (the size of the model we used on CIFAR-10) and $f=3$, and vary the number of workers $n$ in the distributed system.